List of References on Evolutionary Multiobjective Optimization

Carlos A. Coello Coello

         ccoello@cs.cinvestav.mx

CINVESTAV-IPN Sección de Computación Departamento de Ingeniería Eléctrica Av. IPN # 2508 Col. San Pedro Zacatenco México, D.F. 07300 MEXICO

October 23, 2002

[1] Hussein A. Abbass. The Self-Adaptive Pareto Differential Evolution Algorithm. In Congress on Evolutionary Computation (CEC’2002), volume 1, pages 831–836, Piscataway, New Jersey, May 2002. IEEE Service Center.

[2] Hussein A. Abbass, Ruhul Sarker, and Charles Newton. PDE: A Paretofrontier Differential Evolution Approach for Multi-objective Optimization Problems. In Proceedings of the Congress on Evolutionary Computation 2001 (CEC’2001), volume 2, pages 971–978, Piscataway, New Jersey, May 2001. IEEE Service Center.

[3] M. F. Abbod, D. A. Linkens, and M. Mahfouf. Multi-Objective Genetic Optimization for Self-Organizing Fuzzy Logic Control. In Proceedings of UKACC Control’98, pages 1575–1580, University of Wales Swansea, UK, september 1998. IEE.

[4] M.A. Abido. A new multiobjective evolutionary algorithm for environmental/economic power dispatch. In Power Engineering Society Summer Meeting, volume 2, pages 1263–1268. IEEE, 2001.

[5] Jose Aguilar and Pablo Miranda. Approaches Based on Genetic Algorithms for Multiobjective Optimization Problems. In Wolfgang Banzhaf, Jason Daida, Agoston E. Eiben, Max H. Garzon, Vasant Honavar, Mark Jakiela, and Robert E. Smith, editors, GECCO–99: Proceedings of the Genetic and Evolutionary Computation Conference, volume 1, pages 3–10, Orlando, Florida, USA, 1999. Morgan Kaufmann Publishers.

[6] Hernán E. Aguirre, Kiyoshi Tanaka, Tatsuo Sugimura, and Shinjiro Oshita. Halftone Image Generation with Improved Multiobjective Genetic Algorithm. In Eckart Zitzler, Kalyanmoy Deb, Lothar Thiele, Carlos A. Coello Coello, and David Corne, editors, First International Conference on Evolutionary Multi-Criterion Optimization, pages 501–515. Springer-Verlag. Lecture Notes in Computer Science No. 1993, 2001.

[7] José Luis Bernal Agustín. Aplicación de Algoritmos Genéticos al Diseño Óptimo de Sistemas de Distribución de Energía Eléctrica. PhD thesis, Department of Electrical Engineering, University of Zaragoza, Spain, January 1998. (In Spanish).

[8] Frank J. Aherne, Neil A. Thacker, and Peter I. Rockett. Automatic Parameter Selection for Object Recognition using a Parallel Multiobjective Genetic Algorithm. In Proceedings of the 7th International Conference on Computer Analysis of Images and Patterns (CAIP’97), Lecture Notes in Computer Science 1296, pages 559–566, Kiel, Germany, September 1997. Springer Verlag.

[9] Frank J. Aherne, Neil A. Thacker, and Peter I. Rockett. Optimal Pairwise Geometric Histograms. In Adrian F. Clark, editor, Electronic Proceedings of the Eighth British Machine Vision Conference, BMVC97, University of Essex, United Kingdom, September 1997.

[10] Frank J. Aherne, Neil A. Thacker, and Peter I. Rockett. Optimising Object Recognition Parameters using a Parallel Multiobjective Genetic Algorithm. In Proceedings of the 2nd IEE/IEEE International Conference on Genetic Algorithms in Engineering Systems: Innovations and Applications (GALESIA’97), pages 1–6, Glasgow, Scotland, September 1997. IEE.

[11] Ahmad Al-Yamani, Sadiq M. Sait, and Hassan R. Barada. HPTS: Heterogeneous Parallel Tabu Search for VLSI Placement. In Congress on Evolutionary Computation (CEC’2002), volume 1, pages 351–355, Piscataway, New Jersey, May 2002. IEEE Service Center.

[12] M. E. Alemany, C. Andrés, and E. Vicens. A new genetic algorithm for the machine grouping problem with multiple objectives. In Ashayeri, Sullivan, and Ahmad, editors, Proceedings of Flexible Automation and Intelligent Manufacturing (FAIM 1999), pages 461–473, 1999.

[13] Robin Allenson. Genetic Algorithms with Gender for Multi-function Optimisation. Technical Report EPCC-SS92-01, Edinburgh Parallel Computing Centre, Edinburgh, Scotland, 1992.

[14] P. Alotto, A. V. Kuntsevitch, Ch. Magele, G. Molinari, C. Paul, K. Preis, M. Repetto, and K. R. Richter. Multiobjective Optimization in Magnetostatics:

A Proposal for Benchmark Problems. Technical report, Institut für Grundlagen und Theorie Electrotechnik, Technische Universität Graz, Graz, Austria, 1996. http://www-igte.tu-graz.ac.at/team/berl01.htm.

[15] Kevin P. Anchor, Jesse B. Zydallis, Gregg H. Gunsch, and Gary B. Lamont. Extending the Computer Defense Immune System: Network Intrusion Detection with a Multiobjective Evolutionary Programming Approach. In Jonathan Timmis and Peter J. Bentley, editors, First International Conference on Artificial Immune Systems (ICARIS’2002), pages 12–21. University of Kent at Canterbury, UK, September 2002. ISBN 1-902671-32-5.

[16] Jessica M. Anderson, Tessa M. Sayers, and M. G. H. Bell. Optimization of a Fuzzy Logic Traffic Signal Controller by a Multiobjective Genetic Algorithm. In Proceedings of the Ninth International Conference on Road Transport Information and Control, pages 186–190, London, April 1998. IEE.

[17] M. B. Anderson and W. R. Lawrence. Launch Conditions and Aerodynamic Data Extraction By An Elitist Pareto Genetic Algorithm. In AIAA Atmospheric Flight Mechanics Conference, San Diego, California, July 1996. AIAA Paper 96-3361.

[18] M. B. Anderson, W. R. Lawrence, and G. A. Gebert. Using an Elitist Pareto Genetic Algorithm for Aerodynamic Data Extraction. In 4th Aerospace Sciences Meeting and Exhibit, Reno, Nevada, January 1996. AIAA Paper 96-0514.

[19] M.B. Anderson, J.E. Burkhalter, and R.M. Jenkins. Missile aerodynamic shape optimization using genetic algorithms. Journal of Spacecraft and Rockets, 37(5):663–669, September-October 2000.

[20] Murray B. Anderson. Using Pareto Genetic Algorithms for Preliminary Subsonic Wing Design. In 6th AIAA/NASA/USAF Multidisciplinary Analysis and Optimization Symposium, Bellevue, Washington, September 1996. AIAA Paper 96-4023.

[21] Murray B. Anderson, J. E. Burkhalter, and R. M. Jenkins. Missile Aerodynamic Shape Optimization Using Genetic Algorithms. In AIAA Aerospace Sciences Meeting and Exhibit, Reno, Nevada, January 1999. AIAA Paper 99-0261.

[22] Murray B. Anderson and Glenn A. Gerbert. Using Pareto Genetic Algorithms for Preliminary Subsonic Wing Design. Technical Report AIAA-96-4023-CP, AIAA, Washington, D.C., 1996.

[23] Johan Andersson. A Survey of Multiobjective Optimization in Engineering Design. Technical Report LiTH-IKP-R-1097, Department of Mechanical Engineering, Linköping University, Linköping, Sweden, 2000.

[24] Johan Andersson. Multiobjective Optimization in Engineering Design— Applications to Fluid Power Systems. PhD thesis, Division of Fluid and Mechanical Engineering Systems. Department of Mechanical Engineering. Linköping Universitet, Linköping, Sweden, 2001.

[25] Johan Andersson and Petter Krus. Metamodel representations for robustness assessment in multiobjective optimization. In Proceedings of the 13th International Conference on Engineering Design (ICED 01), Glasgow, UK, August 2001.

[26] Johan Andersson, Petter Krus, and David Wallace. Multi-Objective Optimization of Hydraulic Actuation Systems. In Proceedings of the 2000 ASME Design Automation Conference, Baltimore, Maryland, September 2000. ASME Press.

[27] Johan Andersson, Jochen Pohl, and Petter Krus. Design of Objective Functions for Optimization of Multi-domain Systems. In Proceedings of the 1998 ASME Winter Meeting, Anaheim, California, November 1998. ASME Press.

[28] Kiam Heong Ang, Gregory Chong, and Yun Li. Preliminary Statement on the Current Progress of Multi-Objective Evolutionary Algorithm Performance Measurement. In Congress on Evolutionary Computation (CEC’2002), volume 2, pages 1139–1144, Piscataway, New Jersey, May 2002. IEEE Service Center.

[29] Kiam Heong Ang and Yun Li. Multi-Objective Benchmark Studies for Evolutionary Computation. In 2001 Genetic and Evolutionary Computation Conference. Workshop Program, pages 393–396, San Francisco, California, July 2001.

[30] Kiam Heong Ang, Yun Li, and Kay Chen Tan. Multi-Objective Benchmark Functions and Benchmark Studies for Evolutionary Computation. In Proceedings of the International Conference on Computational Intelligence for Modelling and Automation (CIMCA’2001), pages 132–139, Las Vegas, Nevada, July 2001.

[31] M. Jo ao Alves and Jo ao Clímaco. An Interactive Method for 0-1 Multiobjective Problems Using Simulated Annealing and Tabu Search. Journal of Heuristics, 6(3):385–403, August 2000.

[32] B. Arkov, D.C. Evans, P.J. Fleming, D.C. Hill, J.P. Norton, I. Pratt, D. Rees, and K. Rodríguez Vázquez. System Identification Strategies Applied to Aircraft Gas-Turbine Engines. In 14th IFAC World Congress, volume 1, pages 145– 152, Beijing, China, July 1999.

[33] T. Arslan, D. H. Horrocks, and E. Ozdemir. Structural Synthesis of Cell-based VLSI Circuits using a Multi-Objective Genetic Algorithm. IEE Electronic Letters, 32(7):651–652, March 1996.

[34] T. Arslan, E. Ozdemir, M. S. Bright, and D. H. Horrocks. Genetic Synthesis Techniques for Low-Power Digital Signal Processing Circuits. In Proceedings Of The IEE Colloquium On Digital Synthesis, pages 7/1–7/5, London, UK, February 1996. IEE.

[35] A. Augugliaro, L. Dusonchet, and E. Riva Sanseverino. Multiobjective service restoration in distribution networks using an evolutionary approach and fuzzy sets. Electrical Power and Energy Systems, 22(2):103–110, February 2000.

[36] Antonino Augugliaro, Luigi Dusonchet, and Eleonora Riva Sanseverino. Evolving non-dominated solutions in multiobjective service restoration for automated distribution networks. Electric Power Systems Research, 59(3):185– 195, October 2001.

[37] B. Awadh, N. Sepehri, and O. Hawaleshka. A Computer-Aided Process Planning Model Based on Genetic Algorithms. Computers in Operations Research, 22(8):651–652, 1995.

[38] Shapour Azarm, Brian J. Reynolds, and Sanjay Narayanan. Comparison of Two Multiobjective Optimization Techniques With and Within Genetic Algorithms. In CD-ROM Proceedings of the 25th ASME Design Automation Conference, volume Paper No. DETC99/DAC-8584, Las Vegas, Nevada, September 1999.

[39] Thomas Bäck, David B. Fogel, and Zbigniew Michalewicz, editors. Handbook of Evolutionary Computation. Institute of Physics Publishing and Oxford University Press, 1997.

[40] Tapan P. Bagchi. Multiobjective Scheduling by Genetic Algorithms. Kluwer Academic Publishers, Boston, 1999.

[41] Tapan P. Bagchi. Pareto-Optimal Solutions for Multi-objective Production Scheduling Problems. In Eckart Zitzler, Kalyanmoy Deb, Lothar Thiele, Carlos A. Coello Coello, and David Corne, editors, First International Conference on Evolutionary Multi-Criterion Optimization, pages 458–471. Springer-Verlag. Lecture Notes in Computer Science No. 1993, 2001.

[42] Flavio Baita, Francesco Mason, Carlo Poloni, and Walter Ukovich. Genetic algorithm with redundancies for the vehicle scheduling problem. In J. Biethahn and Volker Nissen, editors, Evolutionary Algorithms in Management Applications, pages 341–353. Springer-Verlag, Berlin, 1995.

[43] A. Bajwa, T. Williams, and M.A. Stuchly. Design of broadband radar absorbers with genetic algorithm. In IEEE International Symposium of the Antennas and Propagation Society, volume 4, pages 672–675. IEEE, 2001.

[44] Jerzy Balicki. Evolutionary algorithms for multicriteria optimization of program module allocations. In M. Koksalan and S. Zionts, editors, 15th International Conference on Multiple Criteria Decision Making (MCDM), pages 273–281. Springer-Verlag. Lecture Notes in Economics and Mathematical Sciences. Volume 507, 2001.

[45] Jerzy Balicki and Zybmunt Kitowski. Multicriteria evolutionary algorithm with tabu search for task assignment. In Eckart Zitzler, Kalyanmoy Deb, Lothar Thiele, Carlos A. Coello Coello, and David Corne, editors, First International Conference on Evolutionary Multi-Criterion Optimization, pages 373– 384. Springer-Verlag. Lecture Notes in Computer Science No. 1993, 2001.

[46] Jerzy Balicki and Zygmunt Kitowski. Evolutionary Algorithms for Navigation of Underwater Vehicle. In M. Galicki and K. Tchon, editors, Proceedings of the Second International Workshop on Robot Motion and Control, pages 103– 108, 2001.

[47] R. J. Balling, J. T. Taber, K. Day, and S. Wilson. City Planning with a Multiobjective Genetic Algorithm and a Pareto Set Scanner. In Ian C. Parmee, editor, Proceedings of the Fourth International Conference on Adaptive Computing in Design and Manufacture (ACDM’2000), pages 237–247. PEDC, University of Plymouth, UK, Springer London, 2000.

[48] Richard Balling. Pareto sets in decision-based design. Journal of Engineering Valuation and Cost Analysis, 3:189–198, 2000.

[49] Richard Balling and Scott Wilson. The Maximim Fitness Function for Multiobjective Evolutionary Computation: Application to City Planning. In Lee Spector, Erik D. Goodman, Annie Wu, W.B. Langdon, Hans-Michael Voigt, Mitsuo Gen, Sandip Sen, Marco Dorigo, Shahram Pezeshk, Max H. Garzon, and Edmund Burke, editors, Proceedings of the Genetic and Evolutionary Computation Conference (GECCO’2001), pages 1079–1084, San Francisco, California, 2001. Morgan Kaufmann Publishers.

[50] Richard J. Balling, John T. Taber, Michael R. Brown, and Kirsten Day. Multiobjective Urban Planning Using a Genetic Algorithm. ASCE Journal of Urban Planning and Development, 125(2):86–99, June 1999.

[51] P. Di Barba, M. Farina, and A. Savini. A Meta Optimization Approach to the Pareto Optimal Design of Multiple Wingdings Solenoids. In Kaisa Miettinen, Marko M. Mäkelä, Pekka Neittaanmäki, and Jacques Periaux, editors, Proceedings of EUROGEN’99, Jyväskyl, Finland, 1999. University of Jyváskylä.

[52] P. Di Barba, M. Farina, and A. Savini. Multicriteria Optimization of Air-cored Solenoids with Multiple Windings. In Proceedings of the International Symposium on Non-linear Electromagnetic Systems, ISEM’99, Pavia, Italy, May 1999. IOS Press.

[53] P. Di Barba, M. Farina, and A. Savini. Vector Shape Optimization of an Electrostatic Micromotor using a Genetic Algorithm. In Proceedings of the International Symposium on Electromagnetic Fields in Electrical Engineering, ISEF’99, Pavia, Italy, September 1999.

[54] P. Di Barba, M. Farina, and A. Savini. Multiobjective Design Optimization of Real-Life Devices in Electrical Engineering: A Cost-Effective Evolutionary Approach. In Eckart Zitzler, Kalyanmoy Deb, Lothar Thiele, Carlos A. Coello Coello, and David Corne, editors, First International Conference on Evolutionary Multi-Criterion Optimization, pages 560–573. Springer-Verlag. Lecture Notes in Computer Science No. 1993, 2001.

[55] Helio J. C. Barbosa. A coevolutionary genetic algorithm for a game approach to structural optimization. In Thomas Bäck, editor, Proceedings of the Seventh International Conference on Genetic Algorithms, pages 545–552, San Mateo, California, July 1997. Michigan State University, Morgan Kaufmann Publishers.

[56] Helio J.C. Barbosa and André M.S. Barreto. An interactive genetic algorithm with co-evolution of weights for multiobjective problems. In Lee Spector, Erik D. Goodman, Annie Wu, W.B. Langdon, Hans-Michael Voigt, Mitsuo Gen, Sandip Sen, Marco Dorigo, Shahram Pezeshk, Max H. Garzon, and Edmund Burke, editors, Proceedings of the Genetic and Evolutionary Computation Conference (GECCO’2001), pages 203–210, San Francisco, California, 2001. Morgan Kaufmann Publishers.

[57] L. Barone, L. While, and P. Hingston. Designing Crushers with a Multi-Objective Evolutionary Algorithm. In W.B. Langdon, E. Cantú-Paz, K. Mathias, R. Roy, D. Davis, R. Poli, K. Balakrishnan, V. Honavar, G. Rudolph, J. Wegener, L. Bull, M.A. Potter, A.C. Schultz, J.F. Miller, E. Burke, and N. Jonoska, editors, Proceedings of the Genetic and Evolutionary Computation Conference (GECCO’2002), pages 995–1002, San Francisco, California, July 2002. Morgan Kaufmann Publishers.

[58] Matthieu Basseur, Franck Seynhaeve, and El ghazali Talbi. Design of multiobjective evolutionary algorithms: Application to the flow-shop. In Congress on Evolutionary Computation (CEC’2002), volume 2, pages 1151–1156, Piscataway, New Jersey, May 2002. IEEE Service Center.

[59] L. Bastidas, H. V. Gupta, S. Sorooshian, W. J. Shuttleworth, and Z. L. Yang. Parameter Estimation of a Land Surface Scheme using Multi-Criteria Methods. Journal of Geophysical Research—Atmospheres, 104(D16):19491, 1999.

[60] R. A. Bates, R. Fontana, L. Pronzato, and H. P. Wynn. Multi-Domain Optimisation Using Computer Experiments for Concurrent Engineering. In Ian Parmee, editor, The Integration of Evolutionary and Adaptive Computing Technologies with Product/System Design and Realisation, pages 355–364, Plymouth, United Kingdom, April 1998. Plymouth Engineering Design Centre, Springer-Verlag.

[61] R.A. Bates and H.P. Wynn. The Optimisation of Multivariate Robust Design Criteria. In I.C. Parmee, editor, Proceedings of the Fifth International Conference on Adaptive Computing Design and Manufacture (ACDM 2002), volume 5, pages 285–294, University of Exeter, Devon, UK, April 2002. Springer-Verlag.

[62] A. Baykasoglu. Goal Programming using Multiple Objective Tabu Search. Journal of the Operational Research Society, 52(12):1359–1369, December 2001.

[63] Ricardo P. Beausoleil Delgado. Multiple Criteria Scatter Search. In Jorge Pinho de Sousa, editor, Proceedings of the 4th Metaheuristics International Conference (MIC’2001), pages 539–543. Program Operational Ciencia, Tecnologia, Inovaçao do Quadro Comunitário de Apoio III de Fundaçao para a Ciencia e Tecnologia, Porto, Portugal, July 16–20 2001.

[64] A. D. Belegundu, D. V. Murthy, R. R. Salagame, and E. W. Constants. Multiobjective Optimization of Laminated Ceramic Composites Using Genetic Algorithms. In Fifth AIAA/USAF/NASA Symposium on Multidisciplinary Analysis and Optimization, pages 1015–1022, Panama City, Florida, 1994. AIAA. Paper 84-4363-CP.

[65] A. D. Belegundu and P. L. N. Murthy. A New Genetic Algorithm for Multiobjective Optimization. Technical Report AIAA-96-4180-CP, AIAA, Washington, D.C., 1996.

[66] Ernesto Benini and Andrea Toffolo. Development of High-Performance Airfoils for Axial Flow Compressors Using Evolutionary Computation. Journal of Propulsion and Power, 18(3):544–554, May-June 2002.

[67] P. J. Bentley, editor. Evolutionary Design by Computers. Academic Press Ltd., London, 1999.

[68] P. J. Bentley and J. P. Wakefield. An Analysis of Multiobjective optimization within Genetic Algorithms. Technical Report ENGPJB96, University of Huddersfield, UK, 1996.

[69] P. J. Bentley and J. P. Wakefield. Overview of Generic Evolutionary Design Systems. In Proceedings of the 2nd On-Line World Conference on Evolutionary Computation (WEC2), pages 53–56, 1996.

[70] P. J. Bentley and J. P. Wakefield. Finding Acceptable Solutions in the Pareto-Optimal Range using Multiobjective Genetic Algorithms. In P. K. Chawdhry, R. Roy, and R. K. Pant, editors, Soft Computing in Engineering Design and Manufacturing, Part 5, pages 231–240, London, June 1997. Springer Verlag London Limited. (Presented at the 2nd On-line World Conference on Soft Computing in Design and Manufacturing (WSC2)).

[71] Ester Bernardó i Mansilla and Josep Maria Garrell i Guiu. MOLeCS: A Multiobjective Learning Classifier System. In Darrell Whitley, David Goldberg, Erick Cantú-Paz, Lee Spector, Ian Parmee, and Hans-Georg Beyer, editors, Proceedings of the Genetic and Evolutionary Computation Conference (GECCO’2000), page 390, San Francisco, California, 2000. Morgan Kaufmann.

[72] Alain Berro and Yves Duthen. Search for optimum in dynamic environment: a efficient agent-based method. In 2001 Genetic and Evolutionary Computation Conference. Workshop Program, pages 51–54, San Francisco, California, July 2001.

[73] Zafer Bingul, Ali Sekmen, and Saleh Zein-Sabatto. Adaptive Genetic Algorithms Applied to Dynamic Multi-Objective Problems. In Cihan H. Dagli, Anna L. Buczak, Joydeep Ghosh, Mark Embrechts, Okan Ersoy, and Stephen Kercel, editors, Proceedings of the Artificial Neural Networks in Engineering Conference (ANNIE’2000), pages 273–278, New York, 2000. ASME Press.

[74] Zafer Bingul, Ali Sekmen, and Saleh Zein-Sabatto. Genetic Algorithms Applied to Real Time Multi-Objective Optimization Problems. In Proceedings of the 2000 IEEE SouteastCon Conference (SoutheastCON’2000), pages 95–103, Nashville, Tennessee, April 2000. IEEE.

[75] To Thanh Binh. Eine Entwurfsstrategie für Mehrgrößensysteme zur Polgebietsvorgabe. PhD thesis, Institute of Automation, University of Magdeburg, Germany, 1994.

[76] To Thanh Binh. A Multiobjective Evolutionary Algorithm: The Study Cases. Technical report, Institute for Automation and Communication, Barleben, Germany, January 1999.

[77] To Thanh Binh. A Multiobjective Evolutionary Algorithm: The Study Cases. In Annie S. Wu, editor, Proceedings of the 1999 Genetic and Evolutionary Computation Conference. Workshop Program, pages 127–128, Orlando, Florida, July 1999.

[78] To Thanh Binh and Urlich Korn. An evolution strategy for the multiobjective optimization. In The Second International Conference on Genetic Algorithms (Mendel 96), pages 23–28, Brno, Czech Republic, 1996.

[79] To Thanh Binh and Urlich Korn. Ein multikriterielles Design-Tool für Mehrgrößensysteme mittels Evolutionsstrategien. Fachtagung 1997: Moderne Methoden des Regelungs- und ungsentwurfes, pages 35–42, 1997.

[80] To Thanh Binh and Urlich Korn. MOBES: A multiobjective evolution strategy for constrained optimization problems. In The Third International Conference on Genetic Algorithms (Mendel 97), pages 176–182, Brno, Czech Republic, 1997.

[81] To Thanh Binh and Urlich Korn. Multicriteria control system design using an intelligent evolution strategy with dynamical constraints boundaries. In Proceedings of the Conference for Control of Industrial Systems (CIS’97), volume 2, pages 242–247, Belfort, France, 1997.

[82] To Thanh Binh and Urlich Korn. Multiobjective Evolution Strategy for Constrained Optimization Problems. In Proc. of the 15^{th} IMACS World Congress on Scientific Computation, Modelling and Applied Mathematics, pages 357– 362, Berlin, Germany, 1997.

[83] To Thanh Binh and Urlich Korn. Scalar Optimization With Linear and Nonlinear Constraints Using Evolution Strategies. In Bernd Reusch, editor, Lecture Notes in Computer Science, pages 381–392. Springer-Verlag, April 1997.

[84] To Thanh Binh and Urlich Korn. A parallel multiobjective evolutionary algorithm. Submission paper for Evolutionary Computing 1999, 1999.

[85] To Thanh Binh, Urlich Korn, and J. Kliche. Evolution Strategy Toolbox for use with MATLAB. Technical report, Institute of Automation, University of Magdeburg, Germany, March 1996.

[86] Stefan Bleuler, Martin Brack, Lothar Thiele, and Eckart Zitzler. Multiobjective Genetic Programming: Reducing Bloat Using SPEA2. In Proceedings of the Congress on Evolutionary Computation 2001 (CEC’2001), volume 1, pages 536–543, Piscataway, New Jersey, May 2001. IEEE Service Center.

[87] Tobias Blickle, Jürgen Teich, and Lothar Thiele. System-level synthesis using evolutionary algorithms. Technical Report TIK Report-Nr. 16, Computer Engineering and Communication Networks Lab (TIK), Swiss Federal Institute of Technology (ETH), Gloriastrasse 35, 8092 Zurich, April 1996.

[88] Tobias Blickle, Jürgen Teich, and Lothar Thiele. An evolutionary approach to system-level synthesis. In Proc. 5th International Workshop on Hardware/Software Codesign, pages 167–172. IEEE Computer Society Press, 1997.

[89] Anna Blumel and Brian White. Multiobjective Optimization of Fuzzy Logic Scheduled Controllers for Missile Autopilot Design. In Joint 9th IFSA World Congress and 20th NAFIPS International Conference, volume 3, pages 1758– 1763. IEEE, 2001.

[90] Anna L. Blumel, Evan J. Hughes, and Brian A. White. Fuzzy Autopilot Design using a Multiobjective Evolutionary Algorithm. In 2000 Congress on Evolutionary Computation, volume 1, pages 54–61, Piscataway, New Jersey, July 2000. IEEE Service Center.

[91] Anna L. Blumel, Evan J. Hughes, and Brian A. White. Multi-objective Evolutionary Design of Fuzzy Autopilot Controller. In Eckart Zitzler, Kalyanmoy Deb, Lothar Thiele, Carlos A. Coello Coello, and David Corne, editors, First International Conference on Evolutionary Multi-Criterion Optimization, pages 668–680. Springer-Verlag. Lecture Notes in Computer Science No. 1993, 2001.

[92] Carlos C.H. Borges and Helio J.C. Barbosa. A Non-generational Genetic Algorithm for Multiobjective Optimization. In 2000 Congress on Evolutionary Computation, volume 1, pages 172–179, San Diego, California, July 2000. IEEE Service Center.

[93] Pedro Castro Borges. CHESS-Changing Horizon Efficient Set Search: A Simple Principle for Multiobjective Optimization. Journal of Heuristics, 6(3):405– 418, August 2000.

[94] Pedro Castro Borges and Michael Pilegaard Hansen. A basis for future successes in multiobjective combinatorial optimization. Technical Report IMM-REP-1998-8, Institute of Mathematical Modelling, Technical University of Denmark, March 1998.

[95] C. A. Borghi, D. Casadei, M. Fabbri, and G. Serra. Reduction of the torque ripple in permanent magnet actuators by a multiobjective minimization technique. IEEE Transactions on Magnetics, 34(5):2869–2872, September 1998.

[96] F. Boschetti, M. Dentith, and R. List. Genetic Algorithms Incorporating a Pseudo Subspace Method. In Proceedings of the Second International Conference on Evolutionary Computation, pages 557–560, November 1995.

[97] Martijn C.J. Bot. Improving Induction of Linear Classification Trees with Genetic Programming. In Darrell Whitley, David Goldberg, Erick Cantú-Paz, Lee Spector, Ian Parmee, and Hans-Georg Beyer, editors, Proceedings of the Genetic and Evolutionary Computation Conference (GECCO’2000), pages 403– 410, San Francisco, California, 2000. Morgan Kaufmann.

[98] Jürgen Branke, Thomas Kaußler, and Hartmut Schmeck. Guiding Multi Objective Evolutionary Algorithms Towards Interesting Regions. Technical Report 398, Institute für Angewandte Informatik und Formale Beschreibungsverfahren, Universität Karlsruhe, Karlsruhe, Germany, February 2000.

[99] Jürgen Branke, Thomas Kaußler, and Hartmut Schmeck. Guiding Multi-Objective Evolutionary Algorithms Towards Interesting Regions. In Ian C. Parmee, editor, Fourth International Conference on Adaptive Computing in Design and Manufacture (ACDM 2000), Poster Proceedings, pages 1–4. Plymouth Engineering Design Centre, University of Plymouth, April 2000.

[100] Jürgen Branke, Thomas Kaußler, and Harmut Schmeck. Guidance in Evolutionary Multi-Objective Optimization. Advances in Engineering Software, 32:499–507, 2001.

[101] M. S. Bright and T. Arslan. Effective Design Exploration through Multi-Objective High-Level DSP Synthesis. In IEEE International ASIC/SOC Conference, pages 115–118, Florida, USA, September 1999.

[102] M. S. Bright and T. Arslan. Multi-Objective Design Strategy for High-Level Low-Power Design of DSP Systems. In IEEE International Symposium on Circuits and Systems, ISCAS 99, volume 1, pages 80–83, Florida, USA, May– June 1999.

[103] M. S. Bright and T. Arslan. Optimal Supply Voltage Selection through a Multiobjective Design Strategy. In IEEE 33rd Asilomar Conference on Signals, Systems, and Computers, pages 374–377, Pacific Grove, California, October 1999.

[104] Marc Stephen Bright. Evolutionary Strategies for the High-Level Synthesis of VLSI-Based DSP Systems for Low Power. PhD thesis, University Of Wales Cardiff, School Of Engineering, Circuits And Systems Research Group, Cardiff, Wales, UK, October 1998.

[105] M.-O. Bristeau, R. Glowinski, B. Mantel, J. Periaux, and M. Sefrioui. Genetic algorithms for RCS minimization of aerodynamic shapes by active reflectors. In Y. Rahmat-Samii and E. Michielssen, editors, System Design Using Evolutionary Optimization: Genetic Algorithms. John Wiley, 1998.

[106] Marie-Odile Bristeau, Roland Glowinski, Bertrand Mantel, Jacques Périaux, and Mourad Sefrioui. Genetic Algorithms for Electromagnetic Backscattering: Multiobjective Optimization. In Yahya Rahmat-Samii and Eric Michielssen, editors, Electromagnetic Optimization by Genetic Algorithms, pages 399–434. John Wiley and Sons, Inc., New York, 1999.

[107] C. Brizuela, N. Sannomiya, and Y. Zhao. Multi-Objective Flow-Shop: Preliminary Results. In Eckart Zitzler, Kalyanmoy Deb, Lothar Thiele, Carlos A. Coello Coello, and David Corne, editors, First International Conference on Evolutionary Multi-Criterion Optimization, pages 443–457. Springer-Verlag. Lecture Notes in Computer Science No. 1993, 2001.

[108] Rob A. C. M. Broekmeulen. Facility Management of Distribution Centers for Vegetables and Fruits. In J. Biethahn and Volker Nissen, editors, Evolutionary Algorithms in Management Applications, pages 199–210. Springer-Verlag, Berlin, 1995.

[109] A. J. Brown and M. Thomas. Reengineering the Naval Ship Design Process. In Proceedings of From Research to Reality in Ship Systems Engineering Symposium, page 277, University of Essex, United Kingdom, 1998. ASNE.

[110] Anthony Bucci and Jordan Pollack. Order-theoretic analysis of coevolution problems: Coevolutionary statics. In Alwyn M. Barry, editor, GECCO 2002: Proceedings of the Bird of a Feather Workshops, Genetic and Evolutionary Computation Conference, pages 229–235, New York, July 2002. AAAI.

[111] Dirk Büche and Rolf Dornberger. New Evolutionary Algorithm for Multiobjective Optimization and the Application to Engineering Design Problems. In Proceedings of the Fourth World Congress of Structural and Multidisciplinary Optimization, Dalian, China, 2001.

[112] Dirk Büche, Gianfranco Guidati, Peter Stoll, and Petros Kourmoursakos. Self-Organizing Maps for Pareto Optimization of Airfoils. In Juan Julián Merelo Guervós, Panagiotis Adamidis, Hans-Georg Beyer, José-Luis Fernández-Villaca nas, and Hans-Paul Schwefel, editors, Parallel Problem Solving from Nature—PPSN VII, pages 122–131, Granada, Spain, September 2002. Springer-Verlag. Lecture Notes in Computer Science No. 2439.

[113] Dirk Büche, Michele Milano, and Petros Koumoutsakos. Self-Organizing Maps for Multi-Objective Optimization. In Alwyn M. Barry, editor, GECCO 2002: Proceedings of the Bird of a Feather Workshops, Genetic and Evolutionary Computation Conference, pages 152–155, New York, July 2002. AAAI.

[114] Anna L. Buczak and V. R. Jamalabad. Self-organization of a Heterogeneous Sensor Network by Genetic Algorithms. In C. H. Dagli, M. Akay, A. L. Buczak, O. Ersoy, and B.R. Fernández, editors, Intelligent Engineering Systems Through Artificial Neural Networks, volume 8, pages 259–264, New York, 1998. ASME Press.

[115] Edmund K. Burke, Patrick De Causmaecker, Sanja Petrovic, and Greet Vanden Berghe. A Multi Criteria Meta-heuristic Approach to Nurse Rostering. In Congress on Evolutionary Computation (CEC’2002), volume 2, pages 1197– 1202, Piscataway, New Jersey, May 2002. IEEE Service Center.

[116] E.K. Burke, P. Cowling, J.D. Landa Silva, and S. Petrovic. Combining Hybrid Metaheuristics and Populations for the Multiobjective Optimisation of Space Allocation Problems. In Lee Spector, Erik D. Goodman, Annie Wu, W.B. Langdon, Hans-Michael Voigt, Mitsuo Gen, Sandip Sen, Marco Dorigo, Shahram Pezeshk, Max H. Garzon, and Edmund Burke, editors, Proceedings of the Genetic and Evolutionary Computation Conference (GECCO’2001), pages 1252–1259, San Francisco, California, 2001. Morgan Kaufmann Publishers.

[117] E.K. Burke, P. Cowling, J.D. Landa Silva, and S. Petrovic. On the Performance of Population-Based Metaheuristics for the Space Allocation Problem: An Extended Abstract. In Proceedings of the Metaheuristics International Conference (MIC’2001), pages 579–583, Porto, Portugal, July 2001.

[118] J. M. Cadenas and F. Jiménez. A genetic algorithm for the multiobjective solid transportation problem: a fuzzy approach. In International Symposium on Automotive Technology and Automation, Proceedings for the dedicated conferences on Mechatronics and Supercomputing Applications in the Transportation Industries, pages 327–334, Aachen, Germany, 1994.

[119] Eduardo Camponogara and Sarosh N. Talukdar. A Genetic Algorithm for Constrained and Multiobjective Optimization. In Jarmo T. Alander, editor, 3rd Nordic Workshop on Genetic Algorithms and Their Applications (3NWGA), pages 49–62, Vaasa, Finland, August 1997. University of Vaasa.

[120] Alain Cardon, Theirry Galinho, and Jean-Philippe Vacher. An Agent Based Architecture for Job-Shop Scheduling Problem Using the Spirit of Genetic Algorithm. In Proceedings of Evolutionary Algorithms in Engineering and Computer Science, EUROGEN’99, pages 12–19, Jyväskylä, Finland, May 1999.

[121] Alain Cardon, Theirry Galinho, and Jean-Philippe Vacher. A Multi-Objective Genetic Algorithm in Job Shop Scheduling Problem to Refine an Agents’ Architecture. In Kaisa Miettinen, Marko M. Mäkelä, Pekka Neittaanmäki, and Jacques Periaux, editors, Proceedings of EUROGEN’99, Jyväskyl, Finland, May 1999. University of Jyváskylä.

[122] Alain Cardon, Thierry Galinho, and Jean-Philippe Vacher. Using Genetic Algorithm in Job-Shop Scheduling Problem to Constraints Negociators’ Agents. In Proceedings of Evolutionary Algorithms in Engineering and Computer Science, EUROGEN’99, pages 20–27, Jyväskylä, Finland, May 1999.

[123] Alain Cardon, Thierry Galinho, and Jean-Philippe Vacher. Genetic Algorithms using Multi-Objectives in a Multi-Agent System. Robotics and Autonomous Systems, 33(2–3):179–190, November 2000.

[124] Alain Cardon and Jean-Philippe Vacher. Algorithmes Génétiques dans un Système Multi-Agents pour l’Ordonnancement. Technical report, Crihan, 1998. (In French).

[125] Alain Cardon and Jean-Philippe Vacher. Rapport Technique pour Ouverture de Compte au Crihan sur Machine Parallèle Illiac8. Technical report, Crihan, 1998. http://www.crihan.fr (In French).

[126] W. Matthew Carlyle, Bosun Kim, John W. Fowler, and Esma S. Gel. Comparison of Multiple Objective Genetic Algorithms for Parallel Machine Scheduling Problems. In Eckart Zitzler, Kalyanmoy Deb, Lothar Thiele, Carlos A. Coello Coello, and David Corne, editors, First International Conference on Evolutionary Multi-Criterion Optimization, pages 472–485. Springer-Verlag. Lecture Notes in Computer Science No. 1993, 2001.

[127] David J. Caswell and Gary B. Lamont. Wire-Antenna Geometry Design with Multiobjective Genetic Algorithms. In Congress on Evolutionary Computation (CEC’2002), volume 1, pages 103–108, Piscataway, New Jersey, May 2002. IEEE Service Center.

[128] Sergio Cavalieri and Paolo Gaiardelli. Hybrid genetic algorithms for a multiple-objective scheduling problem. Journal of Intelligent Manufacturing, 9:361–367, 1998.

[129] N. Chakraborti, R. Kumar, and D. Jain. A study of the continuous casting mold using a pareto-converging genetic algorithm. Applied Mathematical Modelling, 25:287–297, 2001.

[130] C. S. Chang, W. Wang, A. C. Liew, and F. S. Wen. Bicriterion optimisation for traction substations in DC railway system using genetic algorithm. IEE Proceedings. B, Electric Power Applications, 145(1):49–56, January 1998.

[131] C. S. Chang, W. Wang, A. C. Liew, F. S. Wen, and D. Srinivasan. Genetic Algorithm Based Bicriterion Optimization for Traction Sustations in DC Railway System. In Proceedings of the Second IEEE International Conference on Evolutionary Computation, pages 11–16, Piscataway, New Jersey, 1995. IEEE Press.

[132] Ni-Bin Chang and H. W. Chen. The Use of Fuzzy Interval Genetic Algorithm for Solving Multiobjective Nonlinear Mixed Integer Programming Model. In Proceedings of the First International Conference on Operations and Quantitative Management, pages 76–82, 1997.

[133] T. J. Chang, N. Meade, and J. E. Beasley. Heuristics for Cardinality Constrained Portfolio Optimization. Technical report, The Management School, Imperial College, London SW7 2AZ, England, May 1998.

[134] A. Chattopadhyay and C.E. Seeley. A simulated annealing technique for multiobjective optimization of intelligent structures. Smart Materials and Structures, 3(2):98–106, June 1994.

[135] H. W. Chen and Ni-Bin Chang. Water pollution control in the river basin by fuzzy genetic algorithm-based multiobjective programming modeling. Water Science and Technology, 37(8):55–63, 1998.

[136] Jian-Jung Chen, David E. Goldberg, Shinn-Ying Ho, and Kumara Sastry. Fitness Inheritance in Multi-Objective Optimization. In W.B. Langdon, E. Cantú-Paz, K. Mathias, R. Roy, D. Davis, R. Poli, K. Balakrishnan, V. Honavar, G. Rudolph, J. Wegener, L. Bull, M.A. Potter, A.C. Schultz, J.F. Miller, E. Burke, and N. Jonoska, editors, Proceedings of the Genetic and Evolutionary Computation Conference (GECCO’2002), pages 319–326, San Francisco, California, July 2002. Morgan Kaufmann Publishers.

[137] Jiang-Hung Chen and Shinn-Ying Ho. Multi-Objective Optimization of Flexible Manufacturing Systems. In Lee Spector, Erik D. Goodman, Annie Wu, W.B. Langdon, Hans-Michael Voigt, Mitsuo Gen, Sandip Sen, Marco Dorigo, Shahram Pezeshk, Max H. Garzon, and Edmund Burke, editors, Proceedings of the Genetic and Evolutionary Computation Conference (GECCO’2001), pages 1260–1267, San Francisco, California, 2001. Morgan Kaufmann Publishers.

[138] Xianming Chen. Pareto Tree Searching Genetic Algorithm: Approaching Pareto Optimal Front by Searching Pareto Optimal Tree. Technical Report NK-CS-2001-002, Department of Computer Science, Nankai University, Tianjin, China, 2001.

[139] Yan Chen, Masakuni Narita, Masashi Tsuji, and Sangduk Sa. A Genetic Algorithm Approach to Optimization for the Radiological Worker Allocation Problem. Health Physics, 70(2):180–186, February 1996.

[140] Franklin Y. Cheng and Dan Li. Multiobjective Optimization Design with Pareto Genetic Algorithm. Journal of Structural Engineering, 123(9):1252–1261, September 1997.

[141] Runwei Cheng, Mitsuo Gen, and Shmuel S. Oren. An Adaptive Hyperplane Approach for Multiple Objective Optimization Problems with Complex Constraints. In Darrell Whitley, David Goldberg, Erick Cantú-Paz, Lee Spector, Ian Parmee, and Hans-Georg Beyer, editors, Proceedings of the Genetic and Evolutionary Computation Conference (GECCO’2000), pages 299–306, San Francisco, California, 2000. Morgan Kaufmann.

[142] A. J. Chipperfield, N. V. Dakev, P. J. Fleming, and J. F. Whidborne. Multiobjective robust control using evolutionary algorithms. In IEEE International Conference on Industrial Technology, Shanghai, China, December 1996.

[143] A. J. Chipperfield and C. M. Fonseca P. J. Fleming. Development of Genetic Optimization Tools for Multi-Objective Optimization Problems in CACSD. In IEE Colloquium, Genetic Algorithms for Control and Systems Engineering, Halifax Hall, University of Sheffield, UK, May 1995. Digest No. 1992/106.

[144] A. J. Chipperfield and P. J. Fleming. Gas Turbine Engine Controller Design using Multiobjective Genetic Algorithms. In A. M. S. Zalzala, editor, Proceedings of the First IEE/IEEE International Conference on Genetic Algorithms in Engineering Systems : Innovations and Applications, GALESIA’95, pages 214–219, Halifax Hall, University of Sheffield, UK, September 1995. IEEE.

[145] A. J. Chipperfield and P. J. Fleming. Multiobjective Gas Turbine Engine Controller Design Using Genetic Algorithms. IEEE Transactions on Industrial Electronics, 43(5), October 1996.

[146] A. J. Chipperfield and P. J. Fleming. Evolutionary Design of Gas Turbine Aero-Engine Controllers. In Frank DiCesare and Mohsen Jafari, editors, Proceedings of the 1998 IEEE International Conference on Systems, Man, and Cybernetics, volume 3, pages 2401–2406. IEEE, 1998.

[147] A.J. Chipperfield, J.F. Whidborne, and P.J. Fleming. Evolutionary Algorithms and Simulated Annealing for MCDM. In T. Gal, T.J. Stewart, and T. Hanne, editors, Multicriteria Decicion Making—Advances in MCDM Models, Algorithms, Theory, and Applications, pages 16.1–16.32. Kluwer Academic Publishing, Boston, Massachusetts, 1999.

[148] Seunghoon Choi and Chisu Wu. Partitioning and Allocation of Objects in Heterogeneous Distributed Environments Using the Niched Pareto Genetic-Algorithm. In Proceedings of 1998 Asia Pacific Software Engineering Conference (APSEC 98), pages 322–329, Taipei, Taiwan, December 1998.

[149] C. Rick Chow. An Evolutionary Approach to Search for NCR-Boards. In David B. Fogel, editor, Proceedings of the 1998 International Conference on Evolutionary Computation, pages 295–300, Piscataway, New Jersey, 1998. IEEE.

[150] Scott E. Cieniawski, J. W. Eheart, and S. Ranjithan. Using Genetic Algorithms to Solve a Multiobjective Groundwater Monitoring Problem. Water Resources Research, 31(2):399–409, February 1995.

[151] T. Clarke and R. Davies. Robust eigenstructure assignment using the genetic algorithm and constrained state feedback. Journal of Systems and Control Engineering, 211(1):53–61, February 1997.

[152] M. Clergue and P. Collard. Dual Genetic Algorithms and Pareto Optimization. In George D. Smith, Nigel C. Steele, and Rudolf F. Albrecht, editors, Artificial Neural Nets and Genetic Algorithms, pages 188–197, Norwich, UK, April 1997. Springer-Verlag.

[153] M. Clergue, P. Collard, and A. Gaspar. DGA and Pareto Elitism : Improving Pareto Optimization. In Second International ICSC Symposium on Soft Computing (SOCO’97) at the Ecole pour les Etudes et la Recherche en Informatique et Electronique (EERIE), Nîmes, France, September 1997.

[154] Jeffery K. Cochran, Shwu-Min Horng, and John W. Fowler. A Multi-Population Genetic Algorithm to Solve Multi-Objective Scheduling Problems for Parallel Machines. Working Paper, 2000. (Submitted to Computers and Operations Research).

[155] Carlos A. Coello Coello. An Updated Survey of GA-Based Multiobjective Optimization Techniques. Technical Report Lania-RD-98-08, Laboratorio Nacional de Informática Avanzada (LANIA), Xalapa, Veracruz, México, December 1998.

[156] Carlos A. Coello Coello. Using a Min-Max Method to solve Multiobjective Optimization Problems with Genetic Algorithms. In IBERAMIA’98. Lecture Notes in Computer Science, pages 303–314, Lisbon, Portugal, October 1998. Springer-Verlag.

[157] Carlos A. Coello Coello. A Comprehensive Survey of Evolutionary-Based Multiobjective Optimization Techniques. Knowledge and Information Systems. An International Journal, 1(3):269–308, August 1999.

[158] Carlos A. Coello Coello. Constraint handling through a multiobjective optimization technique. In Annie S. Wu, editor, Proceedings of the 1999 Genetic and Evolutionary Computation Conference. Workshop Program, pages 117– 118, Orlando, Florida, July 1999.

[159] Carlos A. Coello Coello. An Updated Survey of Evolutionary Multiobjective Optimization Techniques : State of the Art and Future Trends. In 1999 Congress on Evolutionary Computation, pages 3–13, Washington, D.C., July 1999. IEEE Service Center.

[160] Carlos A. Coello Coello. Constraint-handling using an evolutionary multiobjective optimization technique. Civil Engineering and Environmental Systems, 17:319–346, 2000.

[161] Carlos A. Coello Coello. Handling Preferences in Evolutionary Multiobjective Optimization: A Survey. In 2000 Congress on Evolutionary Computation, volume 1, pages 30–37, Piscataway, New Jersey, July 2000. IEEE Service Center.

[162] Carlos A. Coello Coello. Treating Constraints as Objectives for Single-Objective Evolutionary Optimization. Engineering Optimization, 32(3):275– 308, 2000.

[163] Carlos A. Coello Coello. An Updated Survey of GA-Based Multiobjective Optimization Techniques. ACM Computing Surveys, 32(2):109–143, June 2000.

[164] Carlos A. Coello Coello. A Short Tutorial on Evolutionary Multiobjective Optimization. In Eckart Zitzler, Kalyanmoy Deb, Lothar Thiele, Carlos A. Coello Coello, and David Corne, editors, First International Conference on Evolutionary Multi-Criterion Optimization, pages 21–40. Springer-Verlag. Lecture Notes in Computer Science No. 1993, 2001.

[165] Carlos A. Coello Coello. Evolutionary Multi-Objective Optimization: A Critical Review. In Ruhul Sarker, Masoud Mohammadian, and Xin Yao, editors, Evolutionary Optimization, pages 117–146. Kluwer Academic Publishers, New York, February 2002. ISBN 0-7923-7654-4.

[166] Carlos A. Coello Coello and Alan D. Christiansen. An Approach to Multiobjective Optimization Using Genetic Algorithms. In Cihan H. Dagli, Metin Akay, C. L. Philip Chen, Benito R. Fernández, and Joydeep Ghosh, editors, Intelligent Engineering Systems Through Artificial Neural Networks. Volume 5. Fuzzy Logic and Evolutionary Programming, pages 411–416, St. Louis, Missouri, USA, November 1995. ASME Press.

[167] Carlos A. Coello Coello and Alan D. Christiansen. Two New GA-based methods for multiobjective optimization. Civil Engineering Systems, 15(3):207– 243, 1998.

[168] Carlos A. Coello Coello and Alan D. Christiansen. MOSES : A Multiobjective Optimization Tool for Engineering Design. Engineering Optimization, 31(3):337–368, 1999.

[169] Carlos A. Coello Coello and Alan D. Christiansen. Multiobjective optimization of trusses using genetic algorithms. Computers and Structures, 75(6):647–660, May 2000.

[170] Carlos A. Coello Coello, Alan D. Christiansen, and A. Hernández Aguirre. Multiobjective Design Optimization of Counterweight Balancing of a Robot Arm using Genetic Algorithms. In Proceedings of the Seventh International Conference on Tools with Artificial Intelligence, pages 20–23, Herndon, Virginia, U.S.A., November 1995. IEEE Computer Society Press.

[171] Carlos A. Coello Coello, Alan D. Christiansen, and Arturo Hernández Aguirre. Use of Genetic Algorithms for Multiobjective Optimization of Counterweight Balancing of Robot Arms. In Jacob J. G. Chen, editor, EXPERSYS-95 Expert Systems Applications and Artificial Intelligence, pages 243–248, San Francisco, California, November 1995. I. I. T. T. International, Technology Transfer Series.

[172] Carlos A. Coello Coello, Alan D. Christiansen, and Arturo Hernández Aguirre. Using a New GA-Based Multiobjective Optimization Technique for the Design of Robot Arms. Robotica, 16(4):401–414, July–August 1998.

[173] Carlos A. Coello Coello and Nareli Cruz Cortés. An Approach to Solve Multiobjective Optimization Problems Based on an Artificial Immune System. In Jonathan Timmis and Peter J. Bentley, editors, First International Conference on Artificial Immune Systems (ICARIS’2002), pages 212–221. University of Kent at Canterbury, UK, September 2002. ISBN 1-902671-32-5.

[174] Carlos A. Coello Coello, Arturo Hernández Aguirre, and Bill P. Buckles. Evolutionary Multiobjective Design of Combinational Logic Circuits. In Jason Lohn, Adrian Stoica, Didier Keymeulen, and Silvano Colombano, editors, Proceedings of the Second NASA/DoD Workshop on Evolvable Hardware, pages 161–170, Los Alamitos, California, July 2000. IEEE Computer Society.

[175] Carlos A. Coello Coello and Carlos E. Mariano Romero. Evolutionary Algorithms and Multiple Objective Optimization. In Matthias Ehrgott and Xavier Gandibleux, editors, Multiple Criteria Optimization: State of the Art Annotated Bibliographic Surveys, pages 277–331. Kluwer Academic Publishers, Boston, 2002.

[176] Carlos A. Coello Coello and Maximino Salazar Lechuga. MOPSO: A Proposal for Multiple Objective Particle Swarm Optimization. In Congress on Evolutionary Computation (CEC’2002), volume 2, pages 1051–1056, Piscataway, New Jersey, May 2002. IEEE Service Center.

[177] Carlos A. Coello Coello and Gregorio Toscano Pulido. A Micro-Genetic Algorithm for Multiobjective Optimization. In Eckart Zitzler, Kalyanmoy Deb, Lothar Thiele, Carlos A. Coello Coello, and David Corne, editors, First International Conference on Evolutionary Multi-Criterion Optimization, pages 126–140. Springer-Verlag. Lecture Notes in Computer Science No. 1993, 2001.

[178] Carlos A. Coello Coello and Gregorio Toscano Pulido. Multiobjective Optimization using a Micro-Genetic Algorithm. In Lee Spector, Erik D. Goodman, Annie Wu, W.B. Langdon, Hans-Michael Voigt, Mitsuo Gen, Sandip Sen, Marco Dorigo, Shahram Pezeshk, Max H. Garzon, and Edmund Burke, editors, Proceedings of the Genetic and Evolutionary Computation Conference (GECCO’2001), pages 274–282, San Francisco, California, 2001. Morgan Kaufmann Publishers.

[179] Carlos A. Coello Coello, David A. Van Veldhuizen, and Gary B. Lamont. Evolutionary Algorithms for Solving Multi-Objective Problems. Kluwer Academic Publishers, New York, May 2002. ISBN 0-3064-6762-3.

[180] Carlos Artemio Coello Coello. An Empirical Study of Evolutionary Techniques for Multiobjective Optimization in Engineering Design. PhD thesis, Department of Computer Science, Tulane University, New Orleans, LA, April 1996.

[181] David A. Coley. Evolving Green Buildings. In Erick Cantú-Paz, editor, 2002 Genetic and Evolutionary Computation Conference. Late-Breaking Papers, pages 62–68, New York, July 2002.

[182] Y. Collette, P. Siarry, and H.-I. Wong. Multidimensional Data Representation Aimed at the Interpretation of Results from Multiobjective Optimization. In PPSN/SAB Workshop on Multiobjective Problem Solving from Nature (MPSN), Paris, France, September 2000.

[183] David W. Corne, Nick R. Jerram, Joshua D. Knowles, and Martin J. Oates. PESA-II: Region-based Selection in Evolutionary Multiobjective Optimization. In Lee Spector, Erik D. Goodman, Annie Wu, W.B. Langdon, Hans-Michael Voigt, Mitsuo Gen, Sandip Sen, Marco Dorigo, Shahram Pezeshk, Max H. Garzon, and Edmund Burke, editors, Proceedings of the Genetic and Evolutionary Computation Conference (GECCO’2001), pages 283–290, San Francisco, California, 2001. Morgan Kaufmann Publishers.

[184] David W. Corne, Joshua D. Knowles, and Martin J. Oates. The Pareto Envelope-based Selection Algorithm for Multiobjective Optimization. In Marc Schoenauer, Kalyanmoy Deb, Günter Rudolph, Xin Yao, Evelyne Lutton, Juan Julian Merelo, and Hans-Paul Schwefel, editors, Proceedings of the Parallel Problem Solving from Nature VI Conference, pages 839–848, Paris, France, 2000. Springer. Lecture Notes in Computer Science No. 1917.

[185] Lino Costa and Pedro Oliveira. An Elitist Genetic Algorithm for Multiobjective Optimization. In Jorge Pinho de Sousa, editor, Proceedings of the 4th Metaheuristics International Conference—MIC 2001, pages 205–209, Porto, Portugal, July 16–20 2001. Program Operational Ciencia, Tecnologia, Inovaçao do Quadro Comunitário de Apoio III de Fundaçao para a Ciencia e Tecnologia.

[186] Lino Costa and Pedro Oliveira. Evolutionary Algorithms Approach to the Solution of Mixed Integer Non-Linear Programming Problems. Computers and Chemical Engineering, 25:257–266, 2001.

[187] Lino Costa and Pedro Oliveira. An Evolution Strategy for Multiobjective Optimization. In Congress on Evolutionary Computation (CEC’2002), volume 1, pages 97–102, Piscataway, New Jersey, May 2002. IEEE Service Center.

[188] Lino A. Costa, Pedro Oliveira, Isabel N. Figueiredo, and Rogério P. Leal. Compliance Minimization of a Composite Laminated Plate by Genetic Algorithms. In Proceedings of the European Conference on Computational Mechanics— ECCM’99, pages 740–741, Munich, Germany, 1999.

[189] Lino A. Costa, Pedro Oliveira, Isabel N. Figueiredo, Luís F. Roseiro, and Rogério P. Leal. Structural Optimization of Laminated Plates with Genetic Algorithms. In Darrell Whitley, David Goldberg, Erick Cantú-Paz, Lee Spector, Ian Parmee, and Hans-Georg Beyer, editors, Proceedings of the Genetic and Evolutionary Computation Conference (GECCO’2000), pages 621–627, San Francisco, California, 2000. Morgan Kaufmann.

[190] Dan Costelloe, Peter Mooney, and Adam Winstanley. Multi-Objective Optimisation and Dynamic Routing Algorithms in Transportation Networks. In First International Conference on Geographic Information Science, Savannah, Georgia, October 2000.

[191] Dan Costelloe, Peter Mooney, and Adam Winstanley. Multi-Objective Optimisation on Transportation Networks. In Proceedings of the 4th AGILE Conference, Brno, Czech Republic, April 2001.

[192] V. Coverstone-Caroll, J.W. Hartmann, and W.M. Mason. Optimal Multi-Objective Low-Thrust Spacecraft Trajectories. Computer Methods in Applied Mechanics and Engineering, 186:387–402, 2000.

[193] William A. Crossley. Genetic Algorithm Approaches for Multiobjective Design of Rotor Systems. In Proceedings of the 6th AIAA/NASA/ISSMO Symposium on Multidisciplinary Analysis and Optimization, pages 384–394, Bellevue, Washington, September 1996. AIAA Paper 96-4025.

[194] William A. Crossley. Genetic Algorithm with the Kreisselmeier-Steinhauser Function for Multiobjective Constrained Optimization of Rotor Systems. In AIAA 35th Aerospace Sciences Meeting and Exhibit, Reno, Nevada, January 1997. AIAA Paper 97-0080.

[195] William A. Crossley, A. M. Cook, D. W. Fanjoy, and V. B. Venkayya. Using the Two-Branch Tournament Genetic Algorithm for Multiobjective Design. In AIAA/ASME/ASCE/AHS/ASC Structures, Structural Dynamics, and Materials Conference, Long Beach, California, April 1998. AIAA Paper 98-1914.

[196] Xunxue Cui, Miao Li, and Tingjian Fang. Study of Population Diversity of Multiobjective Evolutionary Algorithm Based on Immune and Entropy Principles. In Proceedings of the Congress on Evolutionary Computation 2001 (CEC’2001), volume 2, pages 1316–1321, Piscataway, New Jersey, May 2001. IEEE Service Center.

[197] A. Gaspar Cunha, Pedro Oliveira, and José A. Covas. Use of Genetic Algorithms in Multicriteria Optimization to Solve Industrial Problems. In Thomas Bäck, editor, Proceedings of the Seventh International Conference on Genetic Algorithms, pages 682–688, San Mateo, California, July 1997. Michigan State University, Morgan Kaufmann Publishers.

[198] A. Gaspar Cunha, Pedro Oliveira, and José A. Covas. Genetic Algorithms in Multiobjective Optimization Problems: An Application to Polymer Extrusion. In Annie S. Wu, editor, Proceedings of the 1999 Genetic and Evolutionary Computation Conference. Workshop Program, pages 129–130, Orlando, Florida, July 1999.

[199] Dragan Cvetković. Evolutionary Multi–Objective Decision Support Systems for Conceptual Design. PhD thesis, School of Computing, University of Plymouth, Plymouth, UK, November 2000.

[200] Dragan Cvetković, Ian Parmee, and Eric Webb. Multi-objective Optimisation and Preliminary Airframe Design. In Ian Parmee, editor, The Integration of Evolutionary and Adaptive Computing Technologies with Product/System Design and Realisation, pages 255–267, Plymouth, United Kingdom, April 1998. Plymouth Engineering Design Centre, Springer-Verlag.

[201] Dragan Cvetković and Ian C. Parmee. Evolutionary design and multi–objective optimisation. In 6th European Congress on Intelligent Techniques and Soft Computing EUFIT’98, pages 397–401, Aachen, Germany, September 1998.

[202] Dragan Cvetković and Ian C. Parmee. Genetic Algorithm–based Multi– objective Optimisation and Conceptual Engineering Design. In Congress on Evolutionary Computation – CEC99, volume 1, pages 29–36, Washington D.C., USA, 1999. IEEE.

[203] Dragan Cvetković and Ian C. Parmee. Genetic Algorithms Based Systems for Conceptual Engineering Design. In U. Lindemann, H. Birkhofer, H. Meerkamm, and S. Vajna, editors, Proceedings of the 12th International Conference on Engineering Design ICED’99, volume 2, pages 1035–1038, München, Germany, August 1999. TU München.

[204] Dragan Cvetković and Ian C. Parmee. Use of Preferences for GA–based Multi– objective Optimisation. In Wolfgang Banzhaf, Jason Daida, Agoston E. Eiben, Max H. Garzon, Vasant Honavar, Mark Jakiela, and Robert E. Smith, editors, GECCO–99: Proceedings of the Genetic and Evolutionary Computation Conference, volume 2, pages 1504–1509, Orlando, Florida, USA, 1999. Morgan Kaufmann Publishers.

[205] Dragan Cvetković and Ian C. Parmee. Designer’s preferences and multi– objective preliminary design processes. In Ian C. Parmee, editor, Proceedings of the Fourth International Conference on Adaptive Computing in Design and Manufacture (ACDM’2000), pages 249–260. PEDC, University of Plymouth, UK, Springer London, 2000.

[206] Dragan Cvetković and Ian C. Parmee. Agent-based Support within an Interactive Evolutionary Design System. In I.C. Parmee, editor, Proceedings of the Fifth International Conference on Adaptive Computing Design and Manufacture (ACDM 2002), volume 5, pages 355–367, University of Exeter, Devon, UK, April 2002. Springer-Verlag.

[207] Dragan Cvetković and Ian C. Parmee. Preferences and their Application in Evolutionary Multiobjective Optimisation. IEEE Transactions on Evolutionary Computation, 6(1):42–57, February 2002.

[208] P. Czyzak and A. Jaszkiewicz. A multiobjective metaheuristic approach to the localization of a chain of petrol stations by the capital budgeting model. Control and Cybernetics, 25(1):177–187, 1996.

[209] P. Czyzak and A. Jaszkiewicz. The Multiobjective Metaheuristic Approach for Optimization of Complex Manufacturing Systems. In G. Fandel and T. Gal, editors, Multiple Criteria Decision Making. Proceedings of the XIIth International Conference, pages 591–592, Hagen, Germany, 1997. Springer-Verlag.

[210] P. Czyzak and A. Jaszkiewicz. Pareto Simulated Annealing. In G. Fandel and T. Gal, editors, Multiple Criteria Decision Making. Proceedings of the XIIth International Conference, pages 297–307, Hagen, Germany, 1997. Springer-Verlag.

[211] P. Czyzak and A. Jaszkiewicz. Pareto simulated annealing—a metaheuristic technique for multiple-objective combinatorial optimization. Journal of Multi-Criteria Decision Analysis, 7:34–47, 1998.

[212] Viviane Grunert da Fonseca, Carlos M. Fonseca, and Andreia O. Hall. Inferential Performance Assessment of Stochastic Optimisers and the Attainment Function. In Eckart Zitzler, Kalyanmoy Deb, Lothar Thiele, Carlos A. Coello Coello, and David Corne, editors, First International Conference on Evolutionary Multi-Criterion Optimization, pages 213–225. Springer-Verlag. Lecture Notes in Computer Science No. 1993, 2001.

[213] N. V. Dakev, A. J. Chipperfield, J. F. Whidborne, and P. J. Fleming. An evolutionary algorithm approach for solving optimal control problems. In Proceedings of the 13th International Federation of Automatic Control (IFAC) World Congress, San Francisco, California, 1996.

[214] N. V. Dakev, J. F. Whidborne, and A. J. Chipperfield. H_{∞} design of an EMS control system for a maglev vehicle using evolutionary algorithms. In Proceedings of the First IEE/IEEE International Conference on Genetic Algorithms in Engineering Systems : Innovations and Applications, GALESIA’95, pages 226–231, Sheffield, U.K., 1995.

[215] N. V. Dakev, J. F. Whidborne, A. J. Chipperfield, and P. J. Fleming. H_{∞} design of an EMS control system for a maglev vehicle using evolutionary algorithms. Proceedings of IMechE-I Part I: Journal of Systems and Control Engineering, 311(4):345–355, 1997.

[216] N.V. Dakev, A.J. Chipperfield, and P.J. Fleming. A General Approach for Solving Optimal Control Problems using Optimization Techniques. In IEEE International Conference on Systems, Man, and Cybernetics, volume 5, pages 4503–4508. IEEE, 1995.

[217] P. J. Darwen and X. Yao. A Dilemma for Fitness Sharing with a Scaling Function. In Proceedings of the Second IEEE International Conference on Evolutionary Computation, Piscataway, New Jersey, 1995. IEEE Press.

[218] Dipankar Dasgupta and Fabio A. González. Evolving Complex Fuzzy Classifier Rules Using a Linear Tree Genetic Representation. In Lee Spector, Erik D. Goodman, Annie Wu, W.B. Langdon, Hans-Michael Voigt, Mitsuo Gen, Sandip Sen, Marco Dorigo, Shahram Pezeshk, Max H. Garzon, and Edmund Burke, editors, Proceedings of the Genetic and Evolutionary Computation Conference (GECCO’2001), pages 299–305, San Francisco, California, 2001. Morgan Kaufmann Publishers.

[219] Carlos Manuel Mira de Fonseca. Multiobjective Genetic Algorithms with Applications to Control Engineering Problems. PhD thesis, Department of Automatic Control and Systems Engineering, University of Sheffield, Sheffield, UK, September 1995.

[220] Edwin D. de Jong, Richard A. Watson, and Jordan B. Pollack. Reducing Bloat and Promoting Diversity using Multi-Objective Methods. In Lee Spector, Erik D. Goodman, Annie Wu, W.B. Langdon, Hans-Michael Voigt, Mitsuo Gen, Sandip Sen, Marco Dorigo, Shahram Pezeshk, Max H. Garzon, and Edmund Burke, editors, Proceedings of the Genetic and Evolutionary Computation Conference (GECCO’2001), pages 11–18, San Francisco, California, 2001. Morgan Kaufmann Publishers.

[221] Pierre De Lit, Patrice Latinne, Brahim Rekiek, and Alain Delchambre. Assembly planning with an ordering genetic algorithm. International Journal of Production Research, 39(16):3623–3640, November 2001.

[222] Francisco de Toro, Julio Ortega, Javier Fernández, and A. Díaz. PSFGA: A Parallel Genetic Algorithm for Multiobjective Optimization. In F. Vajda and N. Podhorszki, editors, 10th Euromicro Workshop on Parallel, Distributed and Network-Based Processing, pages 384–391. IEEE, 2002.

[223] Kalyanmoy Deb. Multi-Objective Genetic Algorithms: Problem Difficulties and Construction of Test Problems. Technical Report CI-49/98, Dortmund: Department of Computer Science/LS11, University of Dortmund, Germany, 1998.

[224] Kalyanmoy Deb. Evolutionary Algorithms for Multi-Criterion Optimization in Engineering Design. In Kaisa Miettinen, Marko M. Mäkelä, Pekka Neittaanmäki, and Jacques Periaux, editors, Evolutionary Algorithms in Engineering and Computer Science, chapter 8, pages 135–161. John Wiley & Sons, Ltd, Chichester, UK, 1999.

[225] Kalyanmoy Deb. Multi–Objective Evolutionary Algorithms: Introducing Bias Among Pareto–Optimal Solutions. KanGAL report 99002, Indian Institute of Technology, Kanpur, India, 1999.

[226] Kalyanmoy Deb. Multi-Objective Genetic Algorithms: Problem Difficulties and Construction of Test Problems. Evolutionary Computation, 7(3):205–230, Fall 1999.

[227] Kalyanmoy Deb. Non-Linear Goal Programming using Multi-Objective Genetic Algorithms. Technical Report CI-60/98, Dortmund: Department of Computer Science/LS11, University of Dortmund, Germany, 1999.

[228] Kalyanmoy Deb. Solving Goal Programming Problems Using Multi-Objective Genetic Algorithms. In 1999 Congress on Evolutionary Computation, pages 77–84, Washington, D.C., July 1999. IEEE Service Center.

[229] Kalyanmoy Deb. Multi-objective Evolutionary Optimization: Past, Present and Future. In Ian C. Parmee, editor, Proceedings of the Fourth International Conference on Adaptive Computing in Design and Manufacture (ACDM’2000), pages 225–236. PEDC, University of Plymouth, UK, Springer London, 2000.

[230] Kalyanmoy Deb. Multi-Objective Optimization using Evolutionary Algorithms. John Wiley & Sons, Chichester, UK, 2001. ISBN 0-471-87339-X.

[231] Kalyanmoy Deb. Nonlinear goal programming using multi-objective genetic algorithms. Journal of the Operational Research Society, 52(3):291–302, 2001.

[232] Kalyanmoy Deb, Samir Agrawal, Amrit Pratab, and T. Meyarivan. A Fast Elitist Non-Dominated Sorting Genetic Algorithm for Multi-Objective Optimization: NSGA-II. KanGAL report 200001, Indian Institute of Technology, Kanpur, India, 2000.

[233] Kalyanmoy Deb, Samir Agrawal, Amrit Pratab, and T. Meyarivan. A Fast Elitist Non-Dominated Sorting Genetic Algorithm for Multi-Objective Optimization: NSGA-II. In Marc Schoenauer, Kalyanmoy Deb, Günter Rudolph, Xin Yao, Evelyne Lutton, Juan Julian Merelo, and Hans-Paul Schwefel, editors, Proceedings of the Parallel Problem Solving from Nature VI Conference, pages 849–858, Paris, France, 2000. Springer. Lecture Notes in Computer Science No. 1917.

[234] Kalyanmoy Deb and Tushar Goel. Controlled Elitist Non-dominated Sorting Genetic Algorithms for Better Convergence. In Eckart Zitzler, Kalyanmoy Deb, Lothar Thiele, Carlos A. Coello Coello, and David Corne, editors, First International Conference on Evolutionary Multi-Criterion Optimization, pages 67–81. Springer-Verlag. Lecture Notes in Computer Science No. 1993, 2001.

[235] Kalyanmoy Deb and Tushar Goel. A Hybrid Multi-Objective Evolutionary Approach to Engineering Shape Design. In Eckart Zitzler, Kalyanmoy Deb, Lothar Thiele, Carlos A. Coello Coello, and David Corne, editors, First International Conference on Evolutionary Multi-Criterion Optimization, pages 385–399. Springer-Verlag. Lecture Notes in Computer Science No. 1993, 2001.

[236] Kalyanmoy Deb and Tushar Goyal. Controlled Elitist Non-dominated Sorting Genetic Algorithms for Better Convergence. KanGAL report 200004, Indian Institute of Technology, Kanpur, India, 2000.

[237] Kalyanmoy Deb and Tushar Goyal. Multi-Objective Evolutionary Algorithms for Engineering Shape Design. KanGAL report 200003, Indian Institute of Technology, Kanpur, India, 2000.

[238] Kalyanmoy Deb and Amarendra Kumar. Real-coded Genetic Algorithms with Simulated Binary Crossover: Studies on Multimodal and Multiobjective Problems. Complex Systems, 9:431–454, 1995.

[239] Kalyanmoy Deb and T. Meyarivan. Constrained Test Problems for Multi-Objective Evolutionary Optimization. KanGAL report 200005, Indian Institute of Technology, Kanpur, India, 2000.

[240] Kalyanmoy Deb, A. Patrap, and S. Moitra. Mechanical Component Design for multi-objective using Elitist non-dominated sorting GA. KanGAL report 200002, Indian Institute of Technology, Kanpur, India, 2000.

[241] Kalyanmoy Deb, Amrit Pratab, and Subrajyoti Moitra. Mechanical Component Design for Multiple Objectives Using Elitist Non-dominated Sorting GA. In Marc Schoenauer, Kalyanmoy Deb, Günter Rudolph, Xin Yao, Evelyne Lutton, Juan Julian Merelo, and Hans-Paul Schwefel, editors, Proceedings of the Parallel Problem Solving from Nature VI Conference, pages 859–868, Paris, France, 2000. Springer. Lecture Notes in Computer Science No. 1917.

[242] Kalyanmoy Deb, Amrit Pratap, Sameer Agarwal, and T. Meyarivan. A Fast and Elitist Multiobjective Genetic Algorithm: NSGA–II. IEEE Transactions on Evolutionary Computation, 6(2):182–197, April 2002.

[243] Kalyanmoy Deb, Amrit Pratap, and T. Meyarivan. Constrained Test Problems for Multi-objective Evolutionary Optimization. In Eckart Zitzler, Kalyanmoy Deb, Lothar Thiele, Carlos A. Coello Coello, and David Corne, editors, First International Conference on Evolutionary Multi-Criterion Optimization, pages 284–298. Springer-Verlag. Lecture Notes in Computer Science No. 1993, 2001.

[244] Kalyanmoy Deb, Lothar Thiele, Marco Laumanns, and Eckart Zitzler. Scalable Test Problems for Evolutionary Multi-Objective Optimization. Technical Report 112, Computer Engineering and Networks Laboratory (TIK), Swiss Federal Institute of Technology (ETH), Zurich, Switzerland, 2001.

[245] Kalyanmoy Deb, Lothar Thiele, Marco Laumanns, and Eckart Zitzler. Scalable Multi-Objective Optimization Test Problems. In Congress on Evolutionary Computation (CEC’2002), volume 1, pages 825–830, Piscataway, New Jersey, May 2002. IEEE Service Center.

[246] A. K. Dhingra and B. H. Lee. A Genetic Algorithm Approach to Single and Multiobjective Structural Optimization with Discrete-Continuous Variables. International Journal for Numerical Methods in Engineering, 37:4059–4080, 1994.

[247] P. Di Barba, M. Farina, and A. Savini. Vector Shape Optimisation of an Electrostatic Micromotor using a Genetic Algorithm. COMPEL International Journal for Computation and Mathematics in Electrical and Electronic Engineering, 19(2):576–581, 2000.

[248] P. Di Barba, M. Farina, and A. Savini. An improved technique for enhancing diversity in pareto evolutionary optimization of electromagnetic devices. COMPEL International Journal for Computation and Mathematics in Electrical and Electronic Engineering, 20(2):482–496, 2001.

[249] Alexandre H.F. Dias and Jõao A. de Vasconcelos. Multiobjective genetic algorithms applied to solve optimization problems. IEEE Transactions on Magnetics, 38(2):1133–1136, March 2002.

[250] Renio Díaz and Alejandro Rosete Suárez. A Study of the Capacity of the Stochastic Hill Climbing to Solve Multi-Objective Problems. In Proceedings of the Third International Symposium on Adaptive Systems—Evolutionary Computation and Probabilistic Graphical Models, pages 37–40, Havana, Cuba, March 19–23 2001. Institute of Cybernetics, Mathematics and Physics.

[251] Robert P. Dick and Niraj K. Jha. MOGAC: A Multiobjective Genetic Algorithm for the Co-Synthesis of Hardware-Software Embedded Systems. In IEEE/ACM Conference on Computer Aided Design, pages 522–529, Los Alamitos, California, 1997. IEEE Computer Society Press.

[252] Robert P. Dick and Niraj K. Jha. CORDS: Hardware-Software Co-Synthesis of Reconfigurable Real-Time Distributed Embedded Systems. In Proceedings of the International Conference on Computer-Aided Design, pages 62–68, November 1998.

[253] Robert P. Dick and Niraj K. Jha. MOGAC: A Multiobjective Genetic Algorithm for Hardware-Software Co-synthesis of Hierarchical Heterogeneous Distributed Embedded Systems. IEEE Transactions on Computer-Aided Design of Integrated Circuits and Systems, 17(10):920–935, October 1998.

[254] Robert P. Dick and Niraj K. Jha. MOCSYN: Multiobjective Core-Based Single-Chip System Synthesis. In Proc. Design, Automation and Test in Europe, pages 263–270, March 1999.

[255] C. Dimopoulos and A. M. S. Zalzala. Evolutionary Computation Approaches to Cell Optimisation. In Ian Parmee, editor, The Integration of Evolutionary and Adaptive Computing Technologies with Product/System Design and Realisation, pages 69–83, Plymouth, United Kingdom, April 1998. Plymouth Engineering Design Centre, Springer-Verlag.

[256] C. Dimopoulos and A. M. S. Zalzala. Optimization of Cell Configuration and Comparisons using Evolutionary Computation Approaches. In David B. Fogel, editor, Proceedings of the 1998 International Conference on Evolutionary Computation, pages 148–153, Piscataway, New Jersey, 1998. IEEE.

[257] D. C. Donha, D. S. Desanj, and M. R. Katebi. Genetic Algorithm for Weight Selection in h_{∞} Control Design. In Thomas Bäck, editor, Proceedings of the Seventh International Conference on Genetic Algorithms, pages 599–606, San Mateo, California, July 1997. Michigan State University, Morgan Kaufmann Publishers.

[258] Gerry V. Dozier, Shaun McCullough, Abdollah Homaifar, and Loretta Moore. Multiobjective Evolutionary Path Planning via Fuzzy Tournament Selection. In IEEE International Conference on Evolutionary Computation (ICEC’98), pages 684–689, Piscataway, New Jersey, May 1998. IEEE Press.

[259] Nicole Drechsler, Rolf Drechsler, and Bernd Becker. Multi-Objective Optimization in Evolutionary Algorithms Using Satisfyability Classes. In International Conference on Computational Intelligence (6th Fuzzy Days 1999), volume LNCS 1625, Dortmund, Germany, 1999. Springer-Verlag.

[260] Nicole Drechsler, Rolf Drechsler, and Bernd Becker. Multi-objective Optimisation Based on Relation favour. In Eckart Zitzler, Kalyanmoy Deb, Lothar Thiele, Carlos A. Coello Coello, and David Corne, editors, First International Conference on Evolutionary Multi-Criterion Optimization, pages 154– 166. Springer-Verlag. Lecture Notes in Computer Science No. 1993, 2001.

[261] Rolf Drechsler, Nicole Göckel, and Bernd Becker. Learning Heuristics for OBDD Minimization by Evolutionary Algorithms. In Parallel Problem Solving from Nature (PPSN’96), volume LNCS 1141, pages 730–739, Berlin, Germany, 1996. Springer-Verlag.

[262] N.M. Duarte, A. E. Ruano, C.M. Fonseca, and P.J. Fleming. Accelerating Multi-Objective Control System Design Using a Neuro-Genetic Approach. In 2000 Congress on Evolutionary Computation, volume 1, pages 392–397, Piscataway, New Jersey, July 2000. IEEE Service Center.

[263] Els I. Ducheyne, Robert R. De Wulf, and Bernard De Baets. Bi-objective genetic algorithm for forest management: a comparative study. In Proceedings of the 2001 Genetic and Evolutionary Computation Conference. Late-Breaking Papers, pages 63–66, San Francisco, California, July 2001.

[264] A. Duenas, K. Best, and N. Mort. Solving multiple criteria decision-making problem under uncertainty using protrade and @risk 4.0. In Advances in Fuzzy Systems and Evolutionary Computation. WSEAS Press, 2002.

[265] A. Duenas and N. Mort. A Genetic Algorithm for Multiobjective Optimisation using the interactive Sequential Multiobjective Problem Solving Method. In Proceedings of the International Conference on Artificial Intelligence (IC-AI’2001), pages 633–639, Las Vegas, Nevada, June 2001. CSREA Press.

[266] A. Duenas and N. Mort. A Multiobjective Optimiser with a Fuzzy Genetic Algorithm. In IASTED International Conference on Artificial Intelligence and Applications (AIA’2001), pages 191–197, Marbella, Spain, September 2001. ACTA Press.

[267] D. Dumitrescu, Crina Groşan, and Mihai Oltean. A New Evolutionary Approach for Multiobjective Optimization. Studia Universitatis Babeş–Bolyai, Informatica, XLV(1):51–68, 2000.

[268] D. Dumitrescu, Crina Groşan, and Mihai Oltean. Genetic Chromodynamics for Obtaining Continuous Representation of Pareto Regions. Studia Universitatis Babeş–Bolyai, Informatica, XLVI(1):15–30, 2001.

[269] Hugo Duncan, Gerard Leconte, and Peter Utiger. Using Genetic Algorithms in Industry – Art or Science? In Alwyn M. Barry, editor, GECCO 2002: Proceedings of the Bird of a Feather Workshops, Genetic and Evolutionary Computation Conference, pages 211–214, New York, July 2002. AAAI.

[270] Marc Ebner and Andreas Zell. Evolving a Task Specific Image Operator. In Riccardo Poli, Hans-Michael Voigt, Stefano Cagnoni, David Corne, George D. Smith, and Terence C. Fogarty, editors, Evolutionary Image Analysis, Signal Processing and Telecommunications, pages 74–89. Springer. Lecture Notes in Computer Science Volume 1596, Berlin, May 1999.

[271] R.M. Edwards and G.G. Cook. 3G Tri Band Probe Fed Printed Eccentric Spiral Antenna for Nomadic Wireless Devices Using Optimal Convergence for Pareto Ranked Genetic Algorithm. In Eleven International Conference on Antennas and Propagation, volume 2, pages 537–541. IEEE, 2001.

[272] Matthias Ehrgott. Approximation algorithms for combinatorial multicriteria optimization problems. International Transactions in Operational Research, 7:5–31, 2000.

[273] Matthias Ehrgott and Xavier Gandibleux. An Annotated Bibliography of Multi-objective Combinatorial Optimization. Technical Report 62/2000, Fachbereich Mathematik, Universitat Kaiserslautern, Kaiserslautern, Germany, 2000.

[274] Matthias Ehrgott and Xavier Gandibleux. A Survey and Annotated Bibliography of Multiobjective Combinatorial Optimization. OR Spektrum, 22:425–460, 2000.

[275] Matthias Ehrgott and Xavier Gandibleux. Multiobjective Combinatorial Optimization—Theory, Methodology, and Applications. In Matthias Ehrgott and Xavier Gandibleux, editors, Multiple Criteria Optimization: State of the Art Annotated Bibliographic Surveys, pages 369–444. Kluwer Academic Publishers, Boston, 2002.

[276] Matthias Ehrgott and Xavier Gandibleux, editors. Multiple Criteria Optimization: State of the Art Annotated Bibliographic Surveys. Kluwer Academic Publishers, Boston, 2002. ISBN 1-4020-7128-0.

[277] M. Eisenring, L. Thiele, and E. Zitzler. Conflicting Criteria in Embedded System Design. IEEE Design and Test, 17(2):51–59, April–June 2000.

[278] Anikó Ekárt and S.Z. Németh. Selection Based on the Pareto Nondomination Criterion for Controlling Code Growth in Genetic Programming. Genetic Programming and Evolvable Machines, 2(1):61–73, March 2001.

[279] Neil H. Eklund and Mark J. Embrechts. GA-Based Multi-Objective Optimization of Visible Spectra for Lamp Design. In Cihan H. Dagli, Anna L. Buczak, Joydeep Ghosh, Mark J. Embrechts, and Okan Ersoy, editors, Smart Engineering System Design: Neural Networks, Fuzzy Logic, Evolutionary Programming, Data Mining and Complex Systems, pages 451–456, New York, November 1999. ASME Press.

[280] Neil H. Eklund and Mark J. Embrechts. Determining the Color-Efficiency Pareto Optimal Surface for Filtered Light Sources. In Eckart Zitzler, Kalyanmoy Deb, Lothar Thiele, Carlos A. Coello Coello, and David Corne, editors, First International Conference on Evolutionary Multi-Criterion Optimization, pages 603–611. Springer-Verlag. Lecture Notes in Computer Science No. 1993, 2001.

[281] T. A. Ely, W. A. Crossley, and E. A. Williams. Satellite Constellation Design for Zonal Coverage using Genetic Algorithms. In 8th AAS/AIAA Space Flight Mechanics Meeting, Monterey, California, February 1998.

[282] C. Emmanouilidis, A. Hunter, and J. MacIntyre. A Multiobjective Evolutionary Setting for Feature Selection and a Commonality-Based Crossover Operator. In 2000 Congress on Evolutionary Computation, volume 1, pages 309–316, Piscataway, New Jersey, July 2000. IEEE Service Center.

[283] C. Emmanouilidis, A. Hunter, J. MacIntyre, and C. Cox. Multiple Criteria Genetic Algorithms for Feature Selection in Neurofuzzy Modeling. In 1999 International Joint Conference on Neural Networks, volume 6, pages 4387– 4392, Washington, D.C., July 1999.

[284] C. Emmanouilidis, A. Hunter, J. MacIntyre, and C. Cox. Selecting Features in Neurofuzzy Modelling by Multiobjective Genetic Algorithms. In 9th International Conference on Artificial Neural Networks, volume 2, pages 749–754, Edinburgh, UK, September 1999. IEEE.

[285] Christos Emmanouilidis and Andrew Hunter. A comparison of crossover operators in neural network feature selection with multiobjective evolutionary algorithms. In GECCO-2000 Workshop on Evolutionary Computation in the Development of Artificial Neural Networks, pages 58–60, Las Vegas, Nevada, July 2000.

[286] Mark Erickson, Alex Mayer, and Jeffrey Horn. The Niched Pareto Genetic Algorithm 2 Applied to the Design of Groundwater Remediation Systems. In Eckart Zitzler, Kalyanmoy Deb, Lothar Thiele, Carlos A. Coello Coello, and David Corne, editors, First International Conference on Evolutionary Multi-Criterion Optimization, pages 681–695. Springer-Verlag. Lecture Notes in Computer Science No. 1993, 2001.

[287] Mark Erickson, Alex Mayer, and Jeffrey Horn. Multi-objective optimal design of groundwater remediation systems: application of the niched Pareto genetic algorithm (NPGA). Advances in Water Resources, 25(1):51–65, January 2002.

[288] Mark Erickson, Alex S. Mayer, and Jeffrey Horn. Development of a Multi-Objective Optimization Framework for Groundwater Remediation Design Using the Niched-Pareto Genetic Algorithm. In EOS: Transactions of American Geophysical Union, pages F840–F843. Am. Geophys. Union, 1999.

[289] Henrik Esbensen and Ernest S. Kuh. Design space exploration using the genetic algorithm. In IEEE International Symposium on Circuits and Systems (ISCAS’96), pages 500–503, Piscataway, NJ, 1996. IEEE.

[290] Henrik Esbensen and Ernest S. Kuh. EXPLORER: An Interactive Floorplaner for Design Space Exploration. In Proceedings of the European Design Automation Conference, pages 356–361, 1996.

[291] S. Esquivel, S. Ferrero, and R. Gallard. Parameter settings and representations in pareto-based optimization for job shop scheduling. Cybernetics and Systems, 33(6):559–578, September 2002.

[292] S. Esquivel, S. Ferrero, R. Gallard, C. Salto, H. Alfonso, and M. Schütz. Enhanced evolutionary algorithms for single and multiobjective optimization in the job scheduling problem. Knowledge-Based Systems, 15(1–2):13–25, January 2002.

[293] Susana C. Esquivel, Héctor A. Leiva, and Raúl H. Gallard. Multiplicity in Genetic Algorithms to face Multicriteria Optimization. In 1999 Congress on Evolutionary Computation, pages 85–90, Washington, D.C., July 1999. IEEE Service Center.

[294] Richard M. Everson, Jonathan E. Fieldsend, and Sameer Singh. Full Elite Sets for Multi-Objective Optimisation. In I.C. Parmee, editor, Proceedings of the Fifth International Conference on Adaptive Computing Design and Manufacture (ACDM 2002), volume 5, pages 343–354, University of Exeter, Devon, UK, April 2002. Springer-Verlag.

[295] Ali Farhang-Mehr and Shapour Azarm. Multi-Objective Genetic Algorithms With Concepts from Statistical Thermodynamics. In Lee Spector, Erik D. Goodman, Annie Wu, William B. Langdon, Hans-Michael Voigt, Mitsuo Gen, Sandip Sen, Marco Dorigo, Shahram Pezeshk, Max. H. Garzon, and Edmund Burke, editors, Proceedings of the Genetic and Evolutionary Computation Conference (GECCO’2001), page 1075. Morgan Kaufmann Publishers, San Francisco, California, July 2001.

[296] Ali Farhang-Mehr and Shapour Azarm. Diversity assessment of pareto optimal solution sets: An entropy approach. In Congress on Evolutionary Computation (CEC’2002), volume 1, pages 723–728, Piscataway, New Jersey, May 2002. IEEE Service Center.

[297] M. Farina. A Neural Network Based Generalized Response Surface Multiobjective Evolutionary Algorithm. In Congress on Evolutionary Computation (CEC’2002), volume 1, pages 956–961, Piscataway, New Jersey, May 2002. IEEE Service Center.

[298] M. Farina and P. Amato. On the Optimal Solution Definition for Many-criteria Optimization Problems. In Proceedings of the NAFIPS-FLINT International Conference’2002, pages 233–238, Piscataway, New Jersey, June 2002. IEEE Service Center.

[299] M. Farina, A. Bramanti, and P. Di Barba. Combining Global and Local Search of Non-dominated Solutions in Inverse Electromagnetism. In K.C. Giannakoglou, D.T. Tsahalis, J. Periaux, K.D. Papailiou, and T. Fogarty, editors, Proceedings of the EUROGEN’2001 Conference, pages 196–201, Barcelona, Spain, March 2001. International Center for Numerical Methods in Engineering (CIMNE).

[300] Marco Farina. Cost-effective Evolutionary Strategies for Pareto Optimal Front Approximation in Multiobjective Shape Design Optimization of Electromagnetic Devices. PhD thesis, Department of Electrical Engineering. University of Pavia, Italy, 2001.

[301] Marco Farina. A minimal cost hybrid strategy for pareto optimal front approximation. Evolutionary Optimization, 3(1):41–52, 2001.

[302] Marco Farina, Alessandro Bramanti, and Paolo Di Barba. A GRS Method for Pareto-Optimal Front Identification in Electromagnetic Synthesis. IEE Proceedings—Science, Measurement and Technology, 2002. (In Press).

[303] Hala Fayad. Application of Neural Networks and Genetic Algorithms for Solving Conjunctive Water Use Problems. PhD thesis, Department of Biological and Irrigation Engineering. Utah State University, Logan, Utah, 2001.

[304] C.-M. Feng and J.-J. Lin. Using a genetic algorithm to generate alternative sketch maps for urban planning. Computers, Environment and Urban Systems, 23(2):91–108, March 1999.

[305] Chung-Wei Feng, Liang Liu, and Scott A. Burns. Using Genetic Algorithms to Solve Construction Time-Cost Trade-Off Problems. Journal of Computing in Civil Engineering, 10(3):184–189, 1999.

[306] Luis Fernandes, Isabel Figueiredo, Joaquin Júdice, Lino Costa, and Pedro Oliveira. Application of Genetic Algorithms to Plate Optimization. In S. R. Idelshon, E. O nate, and E. Dvorkin, editors, Computational Mechanics, New Trends and Aplications, Barcelona, Spain, 1998.

[307] Jonathan E. Fieldsend and Sameer Singh. A Multi-Objective Algorithm based upon Particle Swarm Optimisation, an Efficient Data Structure and Turbulence. In Proceedings of the 2002 U.K. Workshop on Computational Intelligence, pages 37–44, Birmingham, UK, September 2002.

[308] Jonathan E. Fieldsend and Sameer Singh. Pareto Multi-Objective Non-Linear Regression Modelling to Aid CAPM Analogous Forecasting. In Proceedings of the IEEE/INNS Joint International Conference on Neural Networks (ICNN’02).

World Congress on Computational Intelligence, volume 1, pages 388–393. IEEE, May 2002.

[309] John H. Fisch, Yun Li, P.C. Kjaer, J.J. Gribble, and T.J.E. Miller. Paretooptimal firing angles for switched reluctance motor control. In Proceedings of the 2nd IEE/IEEE International Conference on Genetic Algorithms in Engineering Systems: Innovations and Applications (GALESIA’97), pages 90–96, Glasgow, Scotland, September 1997. IEE.

[310] P.J. Fleming and R.C. Purshouse. Genetic algorithms in control systems engineering. Technical Report 789, Department of Automatic Control and Systems Engineering, University of Sheffield, Sheffield, UK, May 2001.

[311] Robert Flynn and Porter D. Sherman. Multicriteria Optimization of Aircraft Panels: Determining Viable Genetic Algorithm Configurations. International Journal of Intelligent Systems, 10:987–999, 1995.

[312] Carlos M. Fonseca and Peter J. Fleming. Genetic Algorithms for Multiobjective Optimization: Formulation, Discussion and Generalization. In Stephanie Forrest, editor, Proceedings of the Fifth International Conference on Genetic Algorithms, pages 416–423, San Mateo, California, 1993. University of Illinois at Urbana-Champaign, Morgan Kauffman Publishers.

[313] Carlos M. Fonseca and Peter J. Fleming. An Overview of Evolutionary Algorithms in Multiobjective Optimization. Technical report, Department of Automatic Control and Systems Engineering, University of Sheffield, Sheffield, U. K., 1994.

[314] Carlos M. Fonseca and Peter J. Fleming. Multiobjective Genetic Algorithms Made Easy: Selection, Sharing, and Mating Restriction. In Proceedings of the First International Conference on Genetic Algorithms in Engineering Systems: Innovations and Applications, pages 42–52, Sheffield, UK, September 1995. IEE.

[315] Carlos M. Fonseca and Peter J. Fleming. Multiobjective Optimization and Multiple Constraint Handling with Evolutionary Algorithms I: A Unified Formulation. Technical Report 564, University of Sheffield, Sheffield, UK, January 1995.

[316] Carlos M. Fonseca and Peter J. Fleming. Multiobjective Optimization and Multiple Constraint Handling with Evolutionary Algorithms II: Application Example. Technical Report 565, University of Sheffield, Sheffield, UK, January 1995.

[317] Carlos M. Fonseca and Peter J. Fleming. An Overview of Evolutionary Algorithms in Multiobjective Optimization. Evolutionary Computation, 3(1):1–16, Spring 1995.

[318] Carlos M. Fonseca and Peter J. Fleming. Nonlinear System Identification with Multiobjective Genetic Algorithms. In Proceedings of the 13th World Congress of the International Federation of Automatic Control, pages 187– 192, San Francisco, California, 1996. Pergamon Press.

[319] Carlos M. Fonseca and Peter J. Fleming. On the Performance Assessment and Comparison of Stochastic Multiobjective Optimizers. In Hans-Michael Voigt, Werner Ebeling, Ingo Rechenberg, and Hans-Paul Schwefel, editors, Parallel Problem Solving from Nature—PPSN IV, Lecture Notes in Computer Science, pages 584–593, Berlin, Germany, September 1996. Springer-Verlag.

[320] Carlos M. Fonseca and Peter J. Fleming. Multiobjective Optimization. In Thomas Bäck, David B. Fogel, and Zbigniew Michalewicz, editors, Handbook of Evolutionary Computation, volume 1, pages C4.5:1–C4.5:9. Institute of Physics Publishing and Oxford University Press, 1997.

[321] Carlos M. Fonseca and Peter J. Fleming. Multiobjective Optimization and Multiple Constraint Handling with Evolutionary Algorithms—Part I: A Unified Formulation. IEEE Transactions on Systems, Man, and Cybernetics, Part A: Systems and Humans, 28(1):26–37, 1998.

[322] Carlos M. Fonseca and Peter J. Fleming. Multiobjective Optimization and Multiple Constraint Handling with Evolutionary Algorithms—Part II: A Application Example. IEEE Transactions on Systems, Man, and Cybernetics, Part A: Systems and Humans, 28(1):38–47, 1998.

[323] M. P. Fourman. Compaction of Symbolic Layout using Genetic Algorithms. In Genetic Algorithms and their Applications: Proceedings of the First International Conference on Genetic Algorithms, pages 141–153. Lawrence Erlbaum, 1985.

[324] K. Fujita, N. Hirokawa, S. Akagi, S. Kitamura, and H. Yokohata. Multiobjective optimal design of automotive engine using genetic algorithm. In Proceedings of DETC’98 – ASME Design Engineering Technical Conferences, page 11, 1998.

[325] Tomonari Furukawa. Parameter identification with weightless regularization. International Journal for Numerical Methods in Engineering, 52:219–238, 2001.

[326] Tomonari Furukawa and Gamini Dissanayake. Parameter identification of autonomous vehicles using multi-objective optimization. Engineering Optimization, 34(4):369–395, 2002.

[327] Tomonari Furukawa, Shinobu Yoshimura, and Hiroshi Kawai. Human-like Optimization—A Novel Technique for Computational Design. In H.A. Mang, F.G. Rammerstorfer, and J. Eberhardsteiner, editors, Proceedings of the Fifth World Congress on Computational Mechanics (WCCM V), Vienna, Austria, July 2002. Vienna University of Technology.

[328] Louis Gacôgne. Research of Pareto Set by Genetic Algorithm, Application to Multicriteria Optimization of Fuzzy Controller. In 5th European Congress on Intelligent Techniques and Soft Computing EUFIT’97, pages 837–845, Aachen, Germany, September 1997.

[329] Louis Gacôgne. Multiple Objective Optimization of Fuzzy Rules for Obstacles Avoiding by an Evolution Algorithm with Adaptative Operators. In Proceedings of the Fifth International Mendel Conference on Soft Computing (Mendel’99), pages 236–242, Brno, Czech Republic, June 1999.

[330] C. Gagné, M. Gravel, and W. Price. Scheduling a single machine where setup times are sequence dependent using an ant-colony heuristic. In Abstract Proceedings of ANTS’2000, pages 157–160, Brussels, Belgium, September 2000.

[331] Caroline Gagné, Wilson L. Price, and Marc Gravel. Scheduling a Single Machine with Sequence Dependent Setup Time Using Ant Colony Optimization. Technical Report 2001–003, Faculté des Sciences de L’Administration, Université Laval, Québec, Canada, April 2001. Available

at http://www.fsa.ulaval.ca/rd.

[332] Thierry Galinho, Alain Cardon, and Jean-Philippe Vacher. Genetic Integration in a Multiagent System for Job-Shop Scheduling. In Helder Coelho, editor, Progress in Artificial Intelligence—IBERAMIA’98, pages 76–87, Lisbon, Portugal, October 1998. Springer-Verlag.

[333] Luca Maria Gambardella, Éric Taillard, and Giovanni Agazzi. MACS-VRPTW: A Multiple Ant Colony System for Vehicle Routing Problems with Time Windows. In David Corne, Marco Dorigo, and Fred Glover, editors, New Ideas in Optimization, pages 63–76. McGraw-Hill, 1999.

[334] Xavier Gandibleux and Arnaud Freville. Tabu Search Based Procedure for Solving the 0-1 Multi-Objective Knapsack Problem: The Two Objectives Case. Journal of Heuristics, 6(3):361–383, August 2000.

[335] Xavier Gandibleux, Nazik Mezdaoui, and Arnaud Fréville. A Tabu Search Procedure to Solve Combinatorial Optimisation Problems. In Rafael Caballero, Francisco Ruiz, and Ralph E. Steuer, editors, Advances in Multiple Objective and Goal Programming, volume 455 of Lecture Notes in Economics and Mathematical Systems, pages 291–300. Springer-Verlag, 1997.

[336] Xavier Gandibleux, Hiroyuki Morita, and Naoki Katoh. The Supported Solutions Used as a Genetic Information in a Population Heuristic. In Eckart Zitzler, Kalyanmoy Deb, Lothar Thiele, Carlos A. Coello Coello, and David Corne, editors, First International Conference on Evolutionary Multi-Criterion Optimization, pages 429–442. Springer-Verlag. Lecture Notes in Computer Science No. 1993, 2001.

[337] Sanjeev Garg and Santosh K. Gupta. Multiobjective optimization of a free radical bulk polymerization reactor using genetic algorithm. Macromolecular Theory and Simulations, 8(1):46–53, 1999.

[338] A. Gaspar-Cunha. Modelling and Optimisation of Single Screw Extrusion. PhD thesis, University of Minho, Guimarães, Portugal, 2000.

[339] M. Gen, K. Ida, and Y. Li. Solving bicriteria solid transportation problem with fuzzy numbers by genetic algorithm. International Journal of Computers and Industrial Engineering, 29:537–543, 1995.

[340] Mitsuo Gen and Runwei Cheng. Genetic Algorithms and Engineering Design. John Wiley and Sons, Inc., New York, 1997.

[341] Mitsuo Gen and Runwei Cheng. Genetic Algorithms and Engineering Optimization. Wiley Series in Engineering Design and Automation. John Wiley & Sons, New York, 2000.

[342] Mitsuo Gen, Kenichi Ida, and Jongryul Kim. A Spanning Tree-Based Genetic Algorithm for Bicriteria Topological Network Design. In Proceedings of the 5th IEEE Conference on Evolutionary Computation, pages 15–20, Piscataway, New Jersey, 1998. IEEE Press.

[343] Mitsuo Gen and Yin-Zhen Li. Solving Multi-Objective Transportation Problems by Spanning Tree-based Genetic Algorithm. In Ian Parmee, editor, The Integration of Evolutionary and Adaptive Computing Technologies with Product/System Design and Realisation, pages 95–108, Plymouth, United Kingdom, April 1998. Plymouth Engineering Design Centre, Springer-Verlag.

[344] John S. Gero and Sushil J. Louis. Improving Pareto Optimal Designs Using Genetic Algorithms. Microcomputers in Civil Engineering, 10(4):241–249, 1995.

[345] John S. Gero, Sushil J. Louis, and Sourav Kundu. Evolutionary learning of novel grammars for design improvement. Artificial Intelligence for Engineering Design, Analysis and Manufacturing, 8:83–94, 1994.

[346] V.J. Gillet, W. Khatib, P. Willett, P.J. Fleming, and D.V.S. Green. Combinatorial Library Design using a Multiobjective Genetic Algorithm. Journal of Chemical Information and Computer Sciences, 42(2):375–385, March-April 2002.

[347] V. Gineityte, B. Csukas, and S. Balogh. Combining genetic programming with generic simulation models in evolutionary synthesis. Computers in Industry, 36(3):181–197, June 1998.

[348] A. P. Giotis and K. C. Giannakoglou. Single- and Multi-Objective Airfoil Design Using Genetic Algorithms and Artificial Intelligence. In Kaisa Miettinen, Marko M. Mäkelä, Pekka Neittaanmäki, and Jacques Periaux, editors, Proceedings of EUROGEN’99, Jyväskyl, Finland, 30 May-6 June 1999. University of Jyváskylä.

[349] David E. Goldberg. Genetic Algorithms in Search, Optimization and Machine Learning. Addison-Wesley Publishing Company, Reading, Massachusetts, 1989.

[350] I. Golovkin, R. Mancini, S. Louis, Y. Ochi, K. Fujita, H. Nishimura, H. Shirga, N. Miyanaga, H. Azechi, R. Butzbach, I. Uschmann, E. Förster, J. Delettrez, J. Koch, R.W. Lee, and L. Klein. Spectroscopic Determination of Dynamic Plasma Gradients in Implosion Cores. Physical Review Letters, 88(4), January 2002.

[351] Igor E. Golovkin, Sushil J. Louis, and Roberto C. Mancini. Parallel Implementation of Niched Pareto Genetic Algorithm Code for X-ray Plasma Spectroscopy. In Late Breaking Papers at the 2000 Genetic and Evolutionary Computation Conference, pages 222–227, Las Vegas, Nevada, July 2000.

[352] Igor E. Golovkin, Sushil J. Louis, and Roberto C. Mancini. Parallel Implementation of Niched Pareto Genetic Algorithm Code for X-ray Plasma Spectroscopy. In Congress on Evolutionary Computation (CEC’2002), volume 2, pages 1820–1824, Piscataway, New Jersey, May 2002. IEEE Service Center.

[353] Igor E. Golovkin, Roberto C. Mancini, Sushil J. Louis, Richard W. Lee, and Lewis Klein. Multi-criteria Search and Optimization: an Application to Xray Plasma Spectroscopy. In 2000 Congress on Evolutionary Computation, volume 2, pages 1521–1527, Piscataway, New Jersey, July 2000. IEEE Service Center.

[354] Antonio F. Gómez-Skarmeta, Fernando Jiménez, and Jesús Ibáñez. Pareto– optimality in Fuzzy Modeling. In 6th European Congress on Intelligent Techniques and Soft Computing EUFIT’98, pages 694–700, Aachen, Germany, September 1998.

[355] Marc Gravel, Aaron Luntala Nsakanda, and Wilson Price. Efficient solutions to the cell-formation problem with multiple routings via a double-loop genetic algorithm. European Journal of Operational Research, 109:286–298, 1998.

[356] Marc Gravel, Wilson L. Price, and Caroline Gagné. Scheduling Continuous Casting of Aluminum Using a Multiple-Objective Ant Colony Optimization Metaheuristic. Technical Report 2001–004, Faculté des Sciences de L’Administration, Université Laval, Québec, Canada, April 2001. Available

at http://www.fsa.ulaval.ca/rd.

[357] Garrison W. Greenwood, Xiaobo Sharon Hu, and Joseph G. D’Ambrosio. Fitness Functions for Multiple Objective Optimization Problems: Combining Preferences with Pareto Rankings. In Richard K. Belew and Michael D. Vose, editors, Foundations of Genetic Algorithms 4, pages 437–455, San Mateo, California, 1997. Morgan Kaufmann.

[358] D. Greiner, G. Winter, and J.M. Emperador. Optimising frame structures by different strategies of genetic algorithms. Finite Elements in Analysis and Design, 37(5):381–402, May 2001.

[359] Pierre Grignon and G. M. Fadel. Bi-objective Optimization by Iterative Genetic Algorithms. In EURO-INFORMS conference, Barcelona, Spain, July 1997.

[360] Pierre Grignon and G. M. Fadel. Quality Criteria for Multi-objective optimization solutions obtained with a Genetic Algorithm. In 38Th AIAA/ASME/ASCE/AHS/ASC SDM conference, Orlando, Florida, April 1997. AIAA-97-1658.

[361] Pierre Grignon and Georges M. Fadel. Configuration design optimization method. In Proceedings of DETC’99 – ASME Design Engineering Technical Conferences, Las Vegas, Nevada, September 1999.

[362] Pierre Grignon and Georges M. Fadel. Multiobjective optimization by iterative genetic algorithm. In Proceedings of DETC’99 – ASME Design Engineering Technical Conferences, Las Vegas, Nevada, September 1999.

[363] Pierre Grignon, J. Wodziack, and G. M. Fadel. Bi-Objective optimization of components packing using a genetic algorithm. In NASA/AIAA/ISSMO Multidisciplinary Design and Optimization Conference, pages 352–362, Seattle, Washington, September 1996. AIAA-96-4022-CP.

[364] Pierre M. Grignon. Configuration Design. PhD thesis, Mechanical Engineering Department, Clemson University, Clemson, SC, May 1999.

[365] R. Groppetti and R. Muscia. On a Genetic Multiobjective Approach for the Integration and Optimization of Assembly Product Design and Process Planning. In P. Chedmail, J. C. Bocquet, and D. Dornfeld, editors, Integrated Design and Manufacturing in Mechanical Engineering, pages 61–70. Kluwer Academic Publishers, The Netherlands, 1997.

[366] Darko Grundler. Multiobjective Optimization of Heat Transfer Plant using Decision Table Controller and Genetic Algorithm. In 2000 Congress on Evolutionary Computation, volume 1, pages 517–521, San Diego, California, July 2000. IEEE Service Center.

[367] P.-Y. Gueugniaud, M. Bertin-Maghit, C. Hirschauer, C. Bouchard, B. Vilasco, P. Petit, M. Gen, K. Ida, J. Lee, and J. Kim. Fuzzy Nonlinear Goal Programming Using Genetic Algorithm. Computers & Industrial Engineering, 33(1):39–42, October 1997.

[368] H. V. Gupta, S. Sorroshian, and P. O. Yapo. Towards Improved Calibration of Hydrologic Models: Multiple and Non-Commensurable Measures of Information. Water Resources Research, 34(4):751–763, 1998.

[369] H. Altay Güvenir. A Genetic Algorithm for Multicriteria Inventory Classification. In D.W. Pearson, N.C. Steele, and R.F. Albrecht, editors, Artificial Neural Nets and Genetic Algorithms. Proceedings of the International Conference, pages 6–9, Wien, April 1995. Springer-Verlag.

[370] H. Altay Güvenir and E. Erel. Multicriteria inventory classification using a genetic algorithm. European Journal of Operational Research, 105(1):29–37, February 1998.

[371] Emmanuel Guy and Jean-Philippe Vacher. Application des algorithmes génétiques à l’ordonnancement d’atelier de type job-shop. Master’s thesis, Ecole Supérieure d’Ingénieurs en Génie Electrique, 1996. (In French).

[372] P. Haastrup and Â. Guimarães Pereira. Exploring the Use of Multi-Objective Genetic Algorithms for Reducing Traffic Generated Urban Air and Noise Pollution. In Proceedings of the 5th European Congress on Intelligent and Soft Computing, pages 819–825, Aachen, Germany, September 1997.

[373] Song Yop Hahn. Application of Vector Optimization Employing Modified Genetic Algorithm to Permanent Magnet Motor Design. In Proceedings of the IEEE Conference on Electromagnetic Field Computation, pages 0D1–7, Okayama, Japan, 1996. IEEE Press.

[374] P. Hajela and C. Y. Lin. Genetic search strategies in multicriterion optimal design. Structural Optimization, 4:99–107, 1992.

[375] P. Hajela and J. Yoo. GA Based Fuzzy Optimization for Non-Convex Pareto Surfaces. In 2001 Genetic and Evolutionary Computation Conference. Workshop Program, pages 85–90, San Francisco, California, July 2001.

[376] D. Halhal, G. A. Walters, D. Ouazar, and D. A. Savic. Multi-objective improvement of water distribution systems using a structure messy genetic algorithm approach. Journal of Water Resources Planning and Management ASCE, 123(3):137–146, 1997.

[377] Hatem Hamda, Olga Roudenko, and Marc Schoenauer. Multi-Objective Evolutionary Topological Optimum Design. In I.C. Parmee, editor, Proceedings of the Fifth International Conference on Adaptive Computing Design and Manufacture (ACDM 2002), volume 5, pages 121–132, University of Exeter, Devon, UK, April 2002. Springer-Verlag.

[378] Thomas Hanne. Concepts of a learning objected-oriented problem solver LOOPS. In Proceedings of the 12th International Conference on Multiple Criteria Decision Making, pages 330–339. Springer-Verlag, 1995.

[379] Thomas Hanne. On the convergence of multiobjective evolutionary algorithms. European Journal of Operational Research, 117(3):553–564, 1999.

[380] Thomas Hanne. Global Multiobjective Optimization Using Evolutionary Algorithms. Journal of Heuristics, 6(3):347–360, August 2000.

[381] Thomas Hanne. Global Multiobjective Optimization with Evolutionary Algorithms: Selection Mechanisms and Mutation Control. In Eckart Zitzler, Kalyanmoy Deb, Lothar Thiele, Carlos A. Coello Coello, and David Corne, editors, First International Conference on Evolutionary Multi-Criterion Optimization, pages 197–212. Springer-Verlag. Lecture Notes in Computer Science No. 1993, 2001.

[382] Michael Pilegaard Hansen. Generating a Diversity of Good Solutions to a Practical Combinatorial Problem using Vectorized Simulated Annealing. Technical report, Institute of Mathematical Modelling, Technical University of Denmark, August 1997. Working Paper.

[383] Michael Pilegaard Hansen. Tabu Search in Multiobjective Optimisation : MOTS. In Proceedings of the 13th International Conference on Multiple Criteria Decision Making (MCDM’97), Cape Town, South Africa, January 1997.

[384] Michael Pilegaard Hansen. Metaheuristics for multiple objective combinatorial optimization. PhD thesis, Institute of Mathematical Modelling, Technical University of Denmark, March 1998.

[385] Michael Pilegaard Hansen. Use of Substitute Scalarizing Functions to Guide a Local Search Based Heuristic: The Case of moTSP. Journal of Heuristics, 6(3):419–431, August 2000.

[386] Michael Pilegaard Hansen and Andrzej Jaszkiewicz. Evaluating the quality of approximations to the non-dominated set. Technical Report IMM-REP-1998-7, Technical University of Denmark, March 1998.

[387] M. Hapke, A. Jaszkiewicz, and R. Slowinski. Fuzzy multi-mode resourceconstrained project scheduling with multiple objectives. In J. Weglarz, editor, Recent Advances in Project Scheduling, chapter 16, pages 355–382. Kluwer Academic Publishers, 1998.

[388] Maciej Hapke, Andrzej Jaszkiewicz, and Roman Slowinski. Pareto Simulated Annealing for Fuzzy Multi-Objective Combinatorial Optimization. Journal of Heuristics, 6(3):329–345, August 2000.

[389] Stephen P. Harris and Emmanuel C. Ifeachor. Nonlinear FIR Filter Design by Genetic Algorithm. In 1st Online Conference on Soft Computing, August 1996.

[390] John W. Hartmann. Low-thrust Trajectory Optimization Using Stochastic Optimization Methods. Master’s thesis, Department of Aeronautical and Astronautical Engineering, University of Illinois at Urbana-Champaign, January 1999.

[391] John W. Hartmann, Victoria L. Coverstone-Carroll, and Steven N. Williams. Optimal Interplanetary Spacecraft Trajectories Via A Pareto Genetic Algorithm. In AAS/AIAA Space Flight Mechanics Meeting, Monterey, California, February 1998. Paper No. AAS-98-202.

[392] John W. Hartmann, Victoria L. Coverstone-Carroll, and Steven N. Williams. Optimal Interplanetary Spacecraft Trajectories via a Pareto Genetic Algorithm. The Journal of the Astronautical Sciences, 46(3):267–282, July–September 1998.

[393] A. Heredia-Langner, D.C. Montgomery, and W.M. Carlyle. Solving a multistage partial inspection problem using genetic algorithms. International Journal of Production Research, 40(8):1923–1940, 2002.

[394] Alberto Herreros, Enrique Baeyens, and José R. Perán. Design of Multiobjective Robust Controllers Using Genetic Algorithms. In Annie S. Wu, editor, Proceedings of the 1999 Genetic and Evolutionary Computation Conference. Workshop Program, pages 131–132, Orlando, Florida, July 1999.

[395] Alberto Herreros, Enrique Baeyens, and José R. Perán. MRCD (Multiobjective Robust Controller Design) Genetic Algorithm: Mechanics and Evaluation. In PPSN/SAB Workshop on Multiobjective Problem Solving from Nature (MPSN), Paris, France, September 2000.

[396] Alain Hertz, Brigitte Jaumard, C.C. Ribeiro, and W.P. Formosinho Filho. A multi-criteria tabu search approach to cell formation problems in group technology with multiple objectives. RAIRO/Operations Research, 28(3):303–328, 1994.

[397] Tomoyouki Higashihara and Masayasu Atsumi. Evolutionary Acquisition of Sensory—Action Network of Mobile Robot using Multiobjective Genetic Algorithm. IPSJ SIG-ICS, 98-ICS-111:1–6, 1998. In Japanese.

[398] M. R. Hilliard, G. E. Liepins, M. Palmer, and G. Rangarajen. The computer as a partner in algorithmic design: Automated discovery of parameters for a multiobjective scheduling heuristic. In R. Sharda, B. L. Golden, E. Wasil, O. Balci, and W. Stewart, editors, Impacts of Recent Computer Advances on Operations Research, pages 321–331. North-Holland Publishing Company, New York, 1989.

[399] Mark Hinchliffe, Mark Willis, and Ming Tham. Chemical Process Systems Modelling using Multi-Objective Genetic Programming. In John R. Koza, Wolfgang Banzhaf, Kumar Chellapilla, Kalyanmoy Deb, Marco Dorigo, David B. Fogel, Max H. Garzon, David E. Goldberg, Hitoshi Iba, and Rick L. Riolo, editors, Proceedings of the Third Annual Conference on Genetic Programming, pages 134–139, San Mateo, California, July 1998. University of Wisconsin at Madison, Morgan Kaufmann Publishers.

[400] Tomoyuki Hiroyasu, Mitsunori Miki, and Shinya Watanabe. Distributed Genetic Algorithms with a New Sharing Approach in Multiobjective Optimization Problems. In 1999 Congress on Evolutionary Computation, pages 69–76, Washington, D.C., July 1999. IEEE Service Center.

[401] Tomoyuki Hiroyasu, Mitsunori Miki, and Shinya Watanabe. Divided Range Genetic Algorithms in Multiobjective Optimization Problems. In Proceedings of International Workshop on Emergent Synthesis, pages 57–66, Kobe, Japan, December 1999.

[402] Tomoyuki Hiroyasu, Mitsunori Miki, and Shinya Watanabe. The New Model of Parallel Genetic Algorithm in Multi-Objective Optimization Problems— Divided Range Multi-Objective Genetic Algorithm—. In 2000 Congress on Evolutionary Computation, volume 1, pages 333–340, Piscataway, New Jersey, July 2000. IEEE Service Center.

[403] Shinn-Ying Ho and Xiao-I Chang. An Efficient Generalized Multiobjective Evolutionary Algorithm. In Wolfgang Banzhaf, Jason Daida, Agoston E. Eiben, Max H. Garzon, Vasant Honavar, Mark Jakiela, and Robert E. Smith, editors, GECCO–99: Proceedings of the Genetic and Evolutionary Computation Conference, volume 1, pages 871–878, Orlando, Florida, USA, 1999. Morgan Kaufmann Publishers.

[404] Jeffrey Horn. The Nature of Niching: Genetic Algorithms and the Evolution of Optimal, Cooperative Populations. PhD thesis, University of Illinois at Urbana Champaign, Urbana, Illinois, 1997.

[405] Jeffrey Horn and Nicholas Nafpliotis. Multiobjective Optimization using the Niched Pareto Genetic Algorithm. Technical Report IlliGAl Report 93005, University of Illinois at Urbana-Champaign, Urbana, Illinois, USA, 1993.

[406] Jeffrey Horn, Nicholas Nafpliotis, and David E. Goldberg. A Niched Pareto Genetic Algorithm for Multiobjective Optimization. In Proceedings of the First IEEE Conference on Evolutionary Computation, IEEE World Congress on Computational Intelligence, volume 1, pages 82–87, Piscataway, New Jersey, June 1994. IEEE Service Center.

[407] Chih-Ming Hsu and Chao-Ton Su. Multi-objective machine-component grouping in cellular manufacturing: a genetic algorithm. Production Planning and Control, 9(2):155–166, 1998.

[408] Jianjun Hu, Kisung Seo, Shaobo Li, Zhun Fan, Ronald C. Rosenberg, and Erik D. Goodman. Structure Fitness Sharing (SFS) for Evolutionary Design by Genetic Programming. In W.B. Langdon, E. Cantú-Paz, K. Mathias, R. Roy, D. Davis, R. Poli, K. Balakrishnan, V. Honavar, G. Rudolph, J. Wegener, L. Bull, M.A. Potter, A.C. Schultz, J.F. Miller, E. Burke, and N. Jonoska, editors, Proceedings of the Genetic and Evolutionary Computation Conference (GECCO’2002), pages 780–787, San Francisco, California, July 2002. Morgan Kaufmann Publishers.

[409] Xiabo Sharon Hu, Garrison Greenwood, and Joseph G. D’Ambrosio. An Evolutionary Approach to Hardware/Software Partitioning. In Hans-Michael Voigt, Werner Ebeling, Ingo Rechenberg, and Hans-Paul Schwefel, editors, Parallel Problem Solving from Nature—PPSN IV, pages 900–909. Springer-Verlag. Lecture Notes in Computer Science No. 1141, September 1996.

[410] Xiaohui Hu and Russell Eberhart. Multiobjective Optimization Using Dynamic Neighborhood Particle Swarm Optimization. In Congress on Evolutionary Computation (CEC’2002), volume 2, pages 1677–1681, Piscataway, New Jersey, May 2002. IEEE Service Center.

[411] Evan J. Hughes. Multi-Objective Probabilistic Selection Evolutionary Algorithm. Technical Report DAPS/EJH/56/2000, Department of Aerospace, Power, & Sensors, Cranfield University, RMCS, Shrivenham, UK, SN6 8LA, September 2000.

[412] Evan J. Hughes. Constraint Handling With Uncertain and Noisy Multi-Objective Evolution. In Proceedings of the Congress on Evolutionary Computation 2001 (CEC’2001), volume 2, pages 963–970, Piscataway, New Jersey, May 2001. IEEE Service Center.

[413] Evan J. Hughes. Evolutionary Multi-objective Ranking with Uncertainty and Noise. In Eckart Zitzler, Kalyanmoy Deb, Lothar Thiele, Carlos A. Coello Coello, and David Corne, editors, First International Conference on Evolutionary Multi-Criterion Optimization, pages 329–343. Springer-Verlag. Lecture Notes in Computer Science No. 1993, 2001.

[414] Evan J. Hughes. Multi-Objective Evolutionary Guidance for Swarms. In Congress on Evolutionary Computation (CEC’2002), volume 2, pages 1127– 1132, Piscataway, New Jersey, May 2002. IEEE Service Center.

[415] Evan James Hughes. Radar Cross Section Modelling Using Genetic Algorithms. PhD thesis, Department of Aerospace, Power, & Sensors, Cranfield University, Royal Military College of Science, Shrivenham, UK, May 1998.

[416] Philip Husbands. Distributed Coevolutionary Genetic Algorithms for multi-Criteria and Multi-Constraint Optimisation. In T. Fogarty, editor, Evolutionary Computing, volume 865 of Lecture Notes in Computer Science, pages 150– 165. Springer Verlag, 1994.

[417] K. W. Hutchinson, David S. Todd, and Pratyush Sen. An Evolutionary Multiple Objective Strategy for the Optimisation of Made-To-Order Products with special reference to the Conceptual Design of High Speed Mono Hull Roll-On/Roll-Off Passenger Ferries. In Fast Sea Transportation, FAST 98, Freemantle, Australia, November 1998.

[418] Chul Ju Hyun, Yeongho Kim, and Yeo Keun Kim. A Genetic Algorithm for Multiple Objective Sequencing Problems in Mixed Model Assembly Lines. Computers & Operations Research, 25(7/8):675–690, 1998.

[419] Ester Bernadó i Mansilla and Josep M. Garrell i Guiu. MOLeCS: Using Multiobjective Evolutionary Algorithms for Learning. In Eckart Zitzler, Kalyanmoy Deb, Lothar Thiele, Carlos A. Coello Coello, and David Corne, editors, First International Conference on Evolutionary Multi-Criterion Optimization, pages 696–710. Springer-Verlag. Lecture Notes in Computer Science No. 1993, 2001.

[420] Kenichi Ida, Mitsuo Gen, and Yin-Zhen Li. Solving Multiobjective Chanceconstrained Solid Transportation Problem by Evolutionary Computation. In 5th European Congress on Intelligent Techniques and Soft Computing EUFIT’97, pages 743–747, Aachen, Germany, September 1997.

[421] Ilkka Ikonen, William E. Biles, Anup Kumar, John C. Wissel, and Rammohan K. Ragade. Genetic Algorithm for Packing Three-Dimensional Non-Convex Objects Having Cavities and Holes. In Proceedings of the 7th International Conference on Genetic Algortithms, pages 591–598, East Lansing, Michigan, July 1997. Morgan Kaufmann Publishers.

[422] Hisao Ishibuchi and Tadahiko Murata. Multi-Objective Genetic Local Search Algorithm. In Toshio Fukuda and Takeshi Furuhashi, editors, Proceedings of the 1996 International Conference on Evolutionary Computation, pages 119– 124, Nagoya, Japan, 1996. IEEE.

[423] Hisao Ishibuchi and Tadahiko Murata. Minimizing the Fuzzy Rule and Maximizing Its Performance by a Multi-Objective Genetic Algorithm. In Proceedings of the 6th International Conference on Fuzzy Systems, pages 259–264, Barcelona, Spain, July 1997.

[424] Hisao Ishibuchi and Tadahiko Murata. Multi-Objective Genetic Local Search Algorithm and Its Application to Flowshop Scheduling. IEEE Transactions on Systems, Man and Cybernetics, 28(3):392–403, August 1998.

[425] Hisao Ishibuchi and Tadahiko Murata. Multi-Objective Genetic Local Search for Minimizing the Number of Fuzzy Rules for Pattern Classification Problems. In David B. Fogel, editor, Proceedings of the 1998 IEEE International Conference on Evolutionary Computation, pages 1100–1105, Piscataway, New Jersey, May 1998. IEEE.

[426] Hisao Ishibuchi, Tadahiko Murata, and Mitsuo Gen. Performance Evaluation of Fuzzy Rule-Based Classification Systems Obtained by Multi-Objective Genetic Algorithms. Computers & Industrial Engineering, 35(3–4):575–578, December 1998.

[427] Hisao Ishibuchi, Tadahiko Murata, and I. B. Tursken. Selecting Linguistic Classification Rules by Two-Objective Genetic Algorithms. In Proceedings of the 1995 IEEE International Conference on Systems, Man and Cybernetics, pages 1410–1415, Vancouver, Canada, October 1995. IEEE.

[428] Hisao Ishibuchi, Tadahiko Murata, and I. B. Tursken. Single-Objective and Two-Objective Genetic Algorithms for Selecting Linguistic Rules for Pattern Classification Problems. Fuzzy Sets and Systems, 89(2):135–150, July 1997.

[429] Hisao Ishibuchi and Tomoharu Nakashima. Evolution of Reference Sets in Nearest Neighbor Classification. In B. McKay, X. Yao, C. S. Newton, J.-H. Kim, and T. Furuhashi, editors, Simulated Evolution and Learning. Second Asia-Pacific Conference on Simulated Evolution and Learning, SEAL’98, Canberra, Australia, 1998. Lecture Notes in Computer Science 1585, Springer-Verlag.

[430] Hisao Ishibuchi and Tomoharu Nakashima. Linguistic Rule Extraction by Genetics-Based Machine Learning. In Darrell Whitley, David Goldberg, Erick Cantú-Paz, Lee Spector, Ian Parmee, and Hans-Georg Beyer, editors, Proceedings of the Genetic and Evolutionary Computation Conference (GECCO’2000), pages 195–202, San Francisco, California, 2000. Morgan Kaufmann.

[431] Hisao Ishibuchi and Tomoharu Nakashima. Multi-Objective Pattern and Feature Selection by a Genetic Algorithm. In Darrell Whitley, David Goldberg, Erick Cantú-Paz, Lee Spector, Ian Parmee, and Hans-Georg Beyer, editors, Proceedings of the Genetic and Evolutionary Computation Conference (GECCO’2000), pages 1069–1076, San Francisco, California, 2000. Morgan Kaufmann.

[432] Hisao Ishibuchi and Tomoharu Nakashima. Three-Objective Optimization in Linguistic Function Approximation. In Proceedings of the Congress on Evolutionary Computation 2001 (CEC’2001), volume 1, pages 340–347, Piscataway, New Jersey, May 2001. IEEE Service Center.

[433] Hisao Ishibuchi, Tomoharu Nakashima, and Tetsuya Kuroda. A Hybrid Fuzzy GBML Algorithm for Designing Compact Fuzzy Rule-Based Classification Systems. In Proceedings of 9th IEEE International Conference on Fuzzy Systems, pages 706–711, San Antonio, Texas, May 7–10 2000.

[434] Hisao Ishibuchi, Tomoharu Nakashima, and Tetsuya Kuroda. Minimizing the Number of Fuzzy Rules by Fuzzy Genetics-Based Machine Learning for Pattern Classification Problems. In Proceedings of the 8th Conference on Information Processing and Management of Uncertainty in Knowledge-Based Systems, pages 96–103, Madrid, Spain, July 3–7 2000.

[435] Hisao Ishibuchi and Takashi Yamamoto. Fuzzy Rule Selection by Data Mining Criteria and Genetic Algorithms. In W.B. Langdon, E. Cantú-Paz, K. Mathias, R. Roy, D. Davis, R. Poli, K. Balakrishnan, V. Honavar, G. Rudolph, J. Wegener, L. Bull, M.A. Potter, A.C. Schultz, J.F. Miller, E. Burke, and N. Jonoska, editors, Proceedings of the Genetic and Evolutionary Computation Conference (GECCO’2002), pages 399–406, San Francisco, California, July 2002. Morgan Kaufmann Publishers.

[436] Hisao Ishibuchi, Tadashi Yoshida, and Tadahiko Murata. Balance between Genetic Search and Local Search in Hybrid Evolutionary Multi-Criterion Optimization Algorithms. In W.B. Langdon, E. Cantú-Paz, K. Mathias, R. Roy, D. Davis, R. Poli, K. Balakrishnan, V. Honavar, G. Rudolph, J. Wegener, L. Bull, M.A. Potter, A.C. Schultz, J.F. Miller, E. Burke, and N. Jonoska, editors, Proceedings of the Genetic and Evolutionary Computation Conference (GECCO’2002), pages 1301–1308, San Francisco, California, July 2002. Morgan Kaufmann Publishers.

[437] Hisao Ishibuchi, Tadashi Yoshida, and Tadahiko Murata. Selection of Initial Solutions for Local Search in Multiobjective Genetic Local Search. In Congress on Evolutionary Computation (CEC’2002), volume 1, pages 950– 955, Piscataway, New Jersey, May 2002. IEEE Service Center.

[438] R. S. H. Istepanian and J. F. Whidborne. Multi-objective design of finite wordlength controller structures. In 1999 Congress on Evolutionary Computation, pages 61–68, Washington, D.C., July 1999. IEEE Service Center.

[439] K. Ito, S. Akagi, and M. Nishikawa. A multiobjective optimization approach to a design problem of heat insulation for thermal distribution piping network systems. Journal of Mechanisms, Transmissions and Automation in Design (Transactions of the ASME), 105:206–213, jun 1983.

[440] Y. Itoh, C. Liu, and A. Hammad. Optimizing Rehabilitation Plan of Concrete Bridge Decks Using Multi-Objective Genetic Algorithm. In Optimization in Industry, pages 115–120. American Society of Mechanical Engineers, 1997.

[441] W. Jakob, M. Gorges-Schleuter, and C. Blume. Application of Genetic Algorithms to task planning and learning. In R. Männer and B. Manderick, editors, Parallel Problem Solving from Nature, 2nd Workshop, Lecture Notes in Computer Science, pages 291–300, Amsterdam, 1992. North-Holland Publishing Company.

[442] Ahmad Jan, Masahito Yamamoto, and Azuma Ohuchi. Evolutionary Algorithms for Nurse Scheduling Problem. In 2000 Congress on Evolutionary Computation, volume 1, pages 196–203, Piscataway, New Jersey, July 2000. IEEE Service Center.

[443] Arne Jansson. Fluid Power System Design—A Simulation Approach. PhD thesis, Department of Mechanical Engineering, Linköping University, Linköping, Sweden, 1994.

[444] Arne Jansson, Petter Krus, and Jan-Ove Palmberg. Optimisation of fluid power systems with two alternative non-derivative methods. Technical Report LiTH-IDA-R-94-29, Department of Mechanical Engineering, Linköping University, S-581 83 Linköping, Sweden, 1994.

[445] C. H. Jarvis, N. Stuart, K. Kelsey, and R. H. A. Baker. Towards a Methodology for Selecting a “characteristic” Sample from an Existing Database: An Evolutionary Approach. In Third International Conference/Workshop on Integrating GIS and Environmental Modeling, Santa Fe, New Mexico, January 1996. National Center for Geographic Information and Analysis.

[446] Andrzej Jaszkiewicz. A metaheuristic approach to multiple objective nurse scheduling. Foundations of Computing and Decision Sciences, 22(3):169–184, 1997.

[447] Andrzej Jaszkiewicz. Genetic local search for multiple objective combinatorial optimization. Technical Report RA-014/98, Institute of Computing Science, Poznan University of Technology, 1998.

[448] Andrzej Jaszkiewicz. On the computational effectiveness of multiple objective metaheuristics. In Proceedings of the Fourth International Conference on Multi-Objective Programming and Goal Programming MOPGP’00. Theory & Applications, Berlin–Heidelberg, May 29–June 1 2000. Springer-Verlag.

[449] Andrzej Jaszkiewicz. On the performance of multiple objective genetic local search on the 0/1 knapsack problem. a comparative experiment. Technical Report RA-002/2000, Institute of Computing Science, Poznan University of Technology, Poznań, Poland, July 2000.

[450] Andrzej Jaszkiewicz. A comparative study of multiple-objective metaheuristics on the bi-objective set covering problem and the Pareto memetic algorithm. Technical Report RA-003/01, Institute of Computing Science, Poznan University of Technology, Poznan, Poland, 2001.

[451] Andrzej Jaszkiewicz. Multiple objective metaheuristic algorithms for combinatorial optimization. Poznan University of Technology, Poznan, Poland, 2001. Habilitation thesis.

[452] Andrzej Jaszkiewicz. On the performance of multiple-objective genetic local search on the 0/1 knapsack problem—a comparative experiment. IEEE Transactions on Evolutionary Computation, 6(4):402–412, August 2002.

[453] Andrzej Jaszkiewicz, Maciej Hapke, and Pawel Kominek. Performance of Multiple Objective Evolutionary Algorithms on a Distribution System Design Problem—Computational Experiment. In Eckart Zitzler, Kalyanmoy Deb, Lothar Thiele, Carlos A. Coello Coello, and David Corne, editors, First International Conference on Evolutionary Multi-Criterion Optimization, pages 241–255. Springer-Verlag. Lecture Notes in Computer Science No. 1993, 2001.

[454] Mikkel T. Jensen. Robust and Flexible Scheduling with Evolutionary Computation. PhD thesis, Department of Computer Science. University of Aarhus, Aarhus, Denmark, October 2001.

[455] L.E. Jeremiah, L.L. Gibson, M. Gen, K. Ida, J. Lee, and J. Kim. Fuzzy Nonlinear Goal Programming Using Genetic Algorithm. Computers and Industrial Engineering, 33(1):39–42, October 1997.

[456] F. Jiménez and J. M. Cadenas. An evolutionary program for the multiobjective solid transportation problem with fuzzy goals. Operations Research and Decisions, 2:5–20, 1995.

[457] Fernando Jiménez, Antonio F. Gómez-Skarmeta, Gracia Sánchez, and Kalyanmoy Deb. An Evolutionary Algorithm for Constrained Multi-objective Optimization. In Congress on Evolutionary Computation (CEC’2002), volume 2, pages 1133–1138, Piscataway, New Jersey, May 2002. IEEE Service Center.

[458] Fernando Jiménez, Gracia Sánchez, and Antonio Gómez-Skarmeta. A Pareto-Evolutionary Approach for Goal and Priority Based Multi-objective Optimization Problems. In Proceedings of the 6th Joint Conference on Information Sciences (JCIS 2002), 4th International Workshop on Frontiers in Evolutionary Algorithms, pages 590–593, 2002.

[459] Fernando Jiménez, Gracia Sánchez, Antonio F. Gómez-Skarmeta, H. Roubos, and R. Babuska. Fuzzy modeling with multi-objective neuro-evolutionary algorithms. In Proceedings of the 2002 IEEE International Conference on Systems, Man and Cybernetics (SMC’02). IEEE, 2002.

[460] Fernando Jiménez and José L. Verdegay. Interval multiobjective solid transportation problem via genetic algorithms (IPMU’96). In Proceedings of Information Processing and Management of Uncertainty in Knowledge-Based Systems, pages 787–792, Granada, Spain, 1996.

[461] Fernando Jiménez and José L. Verdegay. Constrained multiobjective optimization by evolutionary algorithms. In Proceedings of the International ICSC Symposium on Engineering of Intelligent Systems (EIS’98), pages 266–271, University of La Laguna, Tenerife, Spain, 1998.

[462] Fernando Jiménez, José L. Verdegay, and Antonio F. Gómez-Skarmeta. Evolutionary Techniques for Constrained Multiobjective Optimization Problems. In Annie S. Wu, editor, Proceedings of the 1999 Genetic and Evolutionary Computation Conference. Workshop Program, pages 115–116, Orlando, Florida, July 1999.

[463] Fernando Jiménez, Antonio F. Gómez-Skarmeta, Hans Roubos, and Robert Babuška. Accurate, transparent, and compact fuzzy models for function approximation and dynamic modeling through multi-objective evolutionary optimization. In Eckart Zitzler, Kalyanmoy Deb, Lothar Thiele, Carlos A. Coello Coello, and David Corne, editors, First International Conference on Evolutionary Multi-Criterion Optimization, pages 653–667. Springer-Verlag. Lecture Notes in Computer Science No. 1993, 2001.

[464] Yaochu Jin, Tatsuya Okabe, and Bernhard Sendhoff. Adapting Weighted Aggregation for Multiobjective Evolution Strategies. In Eckart Zitzler, Kalyanmoy Deb, Lothar Thiele, Carlos A. Coello Coello, and David Corne, editors, First International Conference on Evolutionary Multi-Criterion Optimization, pages 96–110. Springer-Verlag. Lecture Notes in Computer Science No. 1993, 2001.

[465] Yaochu Jin, Tatsuya Okabe, and Bernhard Sendhoff. Dynamic Weighted Aggregation for Evolutionary Multi-Objective Optimization: Why Does It Work and How? In Lee Spector, Erik D. Goodman, Annie Wu, W.B. Langdon, Hans-Michael Voigt, Mitsuo Gen, Sandip Sen, Marco Dorigo, Shahram Pezeshk, Max H. Garzon, and Edmund Burke, editors, Proceedings of the Genetic and Evolutionary Computation Conference (GECCO’2001), pages 1042– 1049, San Francisco, California, 2001. Morgan Kaufmann Publishers.

[466] Yaochu Jin, Markus Olhofer, and Bernhard Sendhoff. On Evolutionary Optimization with Approximate Fitness Functions. In Darrell Whitley, David Goldberg, Erick Cantú-Paz, Lee Spector, Ian Parmee, and Hans-Georg Beyer, editors, Proceedings of the Genetic and Evolutionary Computation Conference (GECCO’2000), pages 786–793, San Francisco, California, 2000. Morgan Kaufmann.

[467] Yaochu Jin and Bernhard Sendhoff. Incorporation of Fuzzy Preferences into Evolutionary Multiobjective Optimization. In W.B. Langdon, E. Cantú-Paz, K. Mathias, R. Roy, D. Davis, R. Poli, K. Balakrishnan, V. Honavar, G. Rudolph, J. Wegener, L. Bull, M.A. Potter, A.C. Schultz, J.F. Miller, E. Burke, and N. Jonoska, editors, Proceedings of the Genetic and Evolutionary Computation Conference (GECCO’2002), page 683, San Francisco, California, July 2002. Morgan Kaufmann Publishers.

[468] Yaochu Jin and Bernhard Sendhoff. Incorporation of Fuzzy Preferences into Evolutionary Multiobjective Optimization. In 4th Asia-Pacific Conference on Simulated Evolution and Learning (SEAL’2002), Singapore, November 2002.

[469] Yaochu Jin, Werner von Seelen, and Bernhard Sendhoff. On generating flexible, complete, consistent and compact(FC^{3}) fuzzy rule systems from data using evolution strategies. IEEE Transactions on Systems, Man, and Cybernetics, 29(4):829–845, 1999.

[470] Yaouchu Jin, Markus Olhofer, and Bernhard Sendhoff. Managing Approximate Models in Evolutionary Aerodynamic Design Optimization. In Proceedings of the Congress on Evolutionary Computation 2001 (CEC’2001), volume 1, pages 592–599, Piscataway, New Jersey, May 2001. IEEE Service Center.

[471] Brian R. Jones, William A. Crossley, and Anastasios S. Lyrintzis. Aerodynamic and Aeroacoustic Optimization of Airfoils via a Parallel Genetic Algorithm. In Proceedings of the 7th AIAA/USAF/NASA/ISSMO Symposium on Multidisciplinary Analysis and Optimization, AIAA-98-4811. AIAA, 1998.

[472] D. F. Jones, M. Tamiz, and S. K. Mirrazavi. Using Genetic Algorithms to Solve Difficult Goal Programs. In Proceedings of the Third International Conference on Multi-Objective Programming and Goal Programming: Theory and Applications (MOPGP’98), Quebec City, Canada, 1998.

[473] D.F. Jones, S.K. Mirrazavi, and M. Tamiz. Multi-objective metaheuristics: An overview of the current state-of-the-art. European Journal of Operational Research, 137(1):1–9, February 2002.

[474] Gareth Jones, Robert D. Brown, David E. Clark, Peter Willett, and Robert C. Glen. Searching Databases of Two-Dimensional and Three-Dimensional Chemical Structures using Genetic Algorithms. In Stephanie Forrest, editor, Proceedings of the Fifth International Conference on Genetic Algorithms, pages 597–602, San Mateo, California, 1993. Morgan Kaufmann.

[475] B. Anthony Kadrovach, Steven R. Michaud, Jesse B. Zydallis, Gary B. Lamont, Barry Secrest, and David Strong. Extending the Simple Genetic Algorithm into Multi-Objective Problems via Mendelian Pressure. In 2001 Genetic and Evolutionary Computation Conference. Workshop Program, pages 181–188, San Francisco, California, July 2001.

[476] B. Anthony Kadrovach, Jesse B. Zydallis, and Gary B. Lamont. Use of mendelian pressure in a multi-objective genetic algorithm. In Congress on Evolutionary Computation (CEC’2002), volume 1, pages 962–967, Piscataway, New Jersey, May 2002. IEEE Service Center.

[477] Tai Kang, G.Y. Cui, and Tapabrata Ray. Design Synthesis of Path Generating Compliant Mechanisms by Evolutionary Optimization of Topology and Shape. In ASME DETC 2000 Design Automation Conference, Baltimore, Maryland, 2000.

[478] Esra Köktener Karasakal and Murat Köksalan. A Simulated Annealing Approach to Bicriteria Scheduling Problems on a Single Machine. Journal of Heuristics, 6(3):311–327, August 2000.

[479] Kosuke Kato and Masatoshi Sakawa. Genetic Algorithms with Decomposition Procedures for Fuzzy Multiobjective 0-1 Programming Problems with Block Angular Structure. In Proceedings of the 1996 International Conference on Evolutionary Computation (ICEC’96), pages 706–709, 1996.

[480] Kosuke Kato and Masatoshi Sakawa. Interactive Decision Making for Multiobjective Block Angular 0-1 Programming Problems with Fuzzy Parameters Through Genetic Algorithms with Decomposition Procedures. In Proceedings of the Sixth IEEE Conference on Fuzzy Systems, pages 1645–1650, 1997.

[481] Kosuke Kato and Masatoshi Sakawa. An interactive fuzzy satisficing method for large scale multiobjective 0-1 programming problems with fuzzy parameters through genetic algorithms. European Journal of Operational Research, 107(3):590–598, June 1998.

[482] Kosuke Kato and Masatoshi Sakawa. Large scale fuzzy multiobjective 0-1 programs through genetic algorithms with decomposition procedures. In L.C. Jain and R.K. Jain, editors, Second International Conference on Knowledge-Based Intelligent Electronic Systems, pages 278–284, Adelaide, Australia, 1998. IEEE.

[483] Kosuke Kato, Masatoshi Sakawa, and Toshinori Ikegame. Interactive Decision Making for Multiobjective Block Angular 0-1 Programming Problems with Fuzzy Parameters Through Genetic Algorithms. Japanese Journal of Fuzzy Theory and Systems, 9(1):49–59, 1997.

[484] Kosuke Kato, Masatoshi Sakawa, and Toshinori Ikegame. An Interactive Fuzzy Criteria Method for Multiobjective 0-1 Programming Problems with Block Angular Structure Using Genetic Algorithms. Electronics and Communications in Japan, 81(8):10–17, 1998.

[485] T. Kawabe and T. Tagami. A New Genetic Algorithm using Pareto Partitioning Method for Robust Partial Model Matching PID Design with Two Degrees of Freedom. In Proceedings of the Third International ICSC (International Computer Science Conventions) Symposia on Intelligent Industrial Automation (IIA’99) and Soft Computing (SOCO’99), pages 562–567, Genova, 1999.

[486] S. Khajehpour. Optimal Conceptual Design of High-Rise Office Buldings. PhD thesis, Civil Engineering Department, University of Waterloo, Ontario, Canada, 2001.

[487] S. Khajehpour and D.E. Grierson. Conceptual Design using Adaptive Computing. In 2001 Genetic and Evolutionary Computation Conference. Workshop Program, pages 62–67, San Francisco, California, July 2001.

[488] Junaid A. Khan, Sadiq M. Sait, and Mahmood R. Minhas. Fuzzy Bialess Simulated Evolution for Multiobjective VLSI Placement. In Congress on Evolutionary Computation (CEC’2002), volume 2, pages 1642–1647, Piscataway, New Jersey, May 2002. IEEE Service Center.

[489] Nazan Khan, David E. Goldberg, and Martin Pelikan. Multi-Objective Bayesian Optimization Algorithm. In W.B. Langdon, E. Cantú-Paz, K. Mathias, R. Roy, D. Davis, R. Poli, K. Balakrishnan, V. Honavar, G. Rudolph, J. Wegener, L. Bull, M.A. Potter, A.C. Schultz, J.F. Miller, E. Burke, and N. Jonoska, editors, Proceedings of the Genetic and Evolutionary Computation Conference (GECCO’2002), page 684, San Francisco, California, July 2002. Morgan Kaufmann Publishers.

[490] Nazan Khan, David E. Goldberg, and Martin Pelikan. Multi-Objective Bayesian Optimization Algorithm. Technical Report 2002009, Illinois Genetic Algorithms Laboratory, University of Illinois at Urbana-Champaign, Urbana, Illinois, March 2002.

[491] Wael Khatib and Peter J. Fleming. An Introduction to Evolutionary Computing for Multidisciplinary Optimization. In Genetic Algorithms in Engineering Systems: Innovations and Applications, pages 7–12. IEE, 1997.

[492] E.F. Khor, K.C. Tan, and T.H. Lee. Multi-Objective Evolutionary Algorithm with Non-Stationary Search Space. In Proceedings of the Congress on Evolutionary Computation 2001 (CEC’2001), volume 1, pages 527–535, Piscataway, New Jersey, May 2001. IEEE Service Center.

[493] E.F. Khor, K.C. Tan, and T.H. Lee. Tabu-Based Exploratory Evolutionary Algorithm for Effective Multi-objective Optimization. In Eckart Zitzler, Kalyanmoy Deb, Lothar Thiele, Carlos A. Coello Coello, and David Corne, editors, First International Conference on Evolutionary Multi-Criterion Optimization, pages 344–358. Springer-Verlag. Lecture Notes in Computer Science No. 1993, 2001.

[494] E.F. Khor, K.C. Tan, M.L. Wang, and T.H. Lee. Evolutionary Algorithm with Dynamic Population Size for Multi-Objective Optimization. In 26th Annual Conference of the IEEE Industrial Electronics Society, volume 3, pages 1686– 1691. IEEE, 2000.

[495] Soon-Thiam Khu. Automatic Calibration of NAM Model with Multi-Objectives Consideration. Technical Report 1298-1, National University of Singapore/Danish Hydraulic Institute, December 1998.

[496] A. A. Khwaja, M. O. Rahman, and M.G. Wagner. Inverse Kinematics of Arbitrary Robotic Manipulators using Genetic Algorithms. In J. Lenarcic and M. L. Justy, editors, Advances in Robot Kinematics: Analysis and Control, pages 375–382. Kluwer Academic Publishers, 1998.

[497] J.-H. Kim and K.-C. Kim. Multicriteria Fuzzy Control Using Evolutionary Programming. Information Sciences, 103(1):71–86, December 1997.

[498] Min-Kyu Kim, Cheol-Gyun Lee, and Hyun-Kyo Jung. Multiobjective Optimal Design of Three-Phase Induction Motor Using Improved Evolution Strategy. IEEE Transactions on Magnetics, 34(5):2980–2983, September 1998.

[499] Y.-J. Kim and J. Ghaboussi. A New Genetic Algorithm Based Control Method Using State Space Reconstruction. In Proceedings of the Second World Conference on Struc. Control, pages 2007–2014, Kyoto, Japan, 1998.

[500] Y.-J. Kim and J. Ghaboussi. A New Method of Reduced Order Feedback Control Using Genetic Algorithms. Earthquake Engineering and Structural Dynamics, 28(2):193–212, 1999.

[501] YongSeong Kim. Feature Selection in Supervised and Unsupervised Learning via Evolutionary Search. PhD thesis, Graduate College, The University of Iowa, December 2001.

[502] YongSeong Kim, W. Nick Street, and Filippo Menczer. An Evolutionary Multi-Objective Local Selection Algorithm for Customer Targeting. In Proceedings of the Congress on Evolutionary Computation 2001 (CEC’2001), volume 2, pages 759–766, Piscataway, New Jersey, May 2001. IEEE Service Center.

[503] Y.S. Kim, W.N. Street, and F. Menczer. Feature Selection in Unsupervised Learning via Evolutionary Search. In Proceedings of the Sixth ACM SIGKDD International Conference on Knowledge Discovery and Data Mining, 2000.

[504] Michael Kirley. MEA: A metapopulation evolutionary algorithm for multiobjective optimisation problems. In Proceedings of the Congress on Evolutionary Computation 2001 (CEC’2001), volume 2, pages 949–956, Piscataway, New Jersey, May 2001. IEEE Service Center.

[505] Marek Kisiel-Dorohinicki, Grzegorz Dobrowolski, and Edward Nawarecki. Evolutionary multi-agent system in multiobjective optimisation. In M. Hamza, editor, Proceedings of the IASTED International Symposium on Applied Informatics (AI’2001), pages 360–365. IASTED/ACTA Press, 2001.

[506] Marek Kisiel-Dorohinicki and Krzysztof Socha. Crowding Factor in Evolutionary Multi-Agent System for Multiobjective Optimization. In Hamid R. Arabnia, editor, Proceedings of the International Conference on Artificial Intelligence (IC-AI’2001), pages 695–700, Las Vegas, Nevada, June 2001. CSREA Press.

[507] Hajime Kita, Yasuyuki Yabumoto, Naoki Mori, and Yoshikazu Nishikawa. Multi-Objective Optimization by Means of the Thermodynamical Genetic Algorithm. In Hans-Michael Voigt, Werner Ebeling, Ingo Rechenberg, and Hans-Paul Schwefel, editors, Parallel Problem Solving from Nature—PPSN IV, Lecture Notes in Computer Science, pages 504–512, Berlin, Germany, September 1996. Springer-Verlag.

[508] J. D. Knowles and D. W. Corne. Local Search, Multiobjective Optimization and the Pareto Archived Evolution Strategy. In B. et al. McKay, editor, Proceedings of Third Australia-Japan Joint Workshop on Intelligent and Evolutionary Systems, pages 209–216, Ashikaga, Japan, November 1999. Ashikaga Institute of Technology.

[509] Joshua Knowles and David Corne. M-PAES: A Memetic Algorithm for Multiobjective Optimization. In 2000 Congress on Evolutionary Computation, volume 1, pages 325–332, Piscataway, New Jersey, July 2000. IEEE Service Center.

[510] Joshua Knowles and David Corne. On Metrics for Comparing Nondominated Sets. In Congress on Evolutionary Computation (CEC’2002), volume 1, pages 711–716, Piscataway, New Jersey, May 2002. IEEE Service Center.

[511] Joshua D. Knowles. Local-Search and Hybrid Evolutionary Algorithms for Pareto Optimization. PhD thesis, The University of Reading, Department of Computer Science, Reading, UK, January 2002.

[512] Joshua D. Knowles and David W. Corne. Assessing the Performance of the Pareto Archived Evolution Strategy. In Annie S. Wu, editor, Proceedings of the 1999 Genetic and Evolutionary Computation Conference. Workshop Program, pages 123–124, Orlando, Florida, July 1999.

[513] Joshua D. Knowles and David W. Corne. The Pareto Archived Evolution Strategy: A New Baseline Algorithm for Multiobjective Optimisation. In 1999 Congress on Evolutionary Computation, pages 98–105, Washington, D.C., July 1999. IEEE Service Center.

[514] Joshua D. Knowles and David W. Corne. Approximating the Nondominated Front Using the Pareto Archived Evolution Strategy. Evolutionary Computation, 8(2):149–172, 2000.

[515] Joshua D. Knowles and David W. Corne. A Comparison of Diverse Approaches to Memetic Multiobjective Combinatorial Optimization. In Proceedings of the 2000 Genetic and Evolutionary Computation Conference Workshop Program, pages 103–108, Las Vegas, Nevada, July 2000.

[516] Joshua D. Knowles and David W. Corne. Benchmark Problem Generators and Results for the Multiobjective Degree-Constrained Minimum Spanning Tree Problem. In Lee Spector, Erik D. Goodman, Annie Wu, W.B. Langdon, Hans-Michael Voigt, Mitsuo Gen, Sandip Sen, Marco Dorigo, Shahram Pezeshk, Max H. Garzon, and Edmund Burke, editors, Proceedings of the Genetic and Evolutionary Computation Conference (GECCO’2001), pages 424–431, San Francisco, California, 2001. Morgan Kaufmann Publishers.

[517] Joshua D. Knowles and David W. Corne. A Comparative Assessment of Memetic, Evolutionary, and Constructive Algorithms for the Multiobjective d-MST Problem. In 2001 Genetic and Evolutionary Computation Conference. Workshop Program, pages 162–167, San Francisco, California, July 2001.

[518] Joshua D. Knowles and David W. Corne. A Comparison of Encodings and Algorithms for Multiobjective Minimum Spanning Tree Problems. In Proceedings of the Congress on Evolutionary Computation 2001 (CEC’2001), volume 1, pages 544–551, Piscataway, New Jersey, May 2001. IEEE Service Center.

[519] Joshua D. Knowles, David W. Corne, and Martin J. Oates. On the Assessment of Multiobjective Approaches to the Adaptive Distributed Database Management Problem. In Marc Schoenauer, Kalyanmoy Deb, Günter Rudolph, Xin Yao, Evelyne Lutton, Juan Julian Merelo, and Hans-Paul Schwefel, editors, Proceedings of the Sixth International Conference on Parallel Problem Solving from Nature (PPSN VI), pages 869–878, Berlin, September 2000. Springer.

[520] Joshua D. Knowles, David W. Corne, and Martin J. Oates. The Pareto-Envelope based Selection Algorithm for Multiobjective Optimization. In Proceedings of the Sixth International Conference on Parallel Problem Solving from Nature (PPSN VI), pages 839–848, Berlin, September 2000. Springer.

[521] Joshua D. Knowles, Martin J. Oates, and David W. Corne. Multiobjective Evolutionary Algorithms Applied to two Problems in Telecommunications. BT Technology Journal, 18(4):51–64, October 2000.

[522] Joshua D. Knowles, Richard A. Watson, and David W. Corne. Reducing Local Optima in Single-Objective Problems by Multi-objectivization. In Eckart Zitzler, Kalyanmoy Deb, Lothar Thiele, Carlos A. Coello Coello, and David Corne, editors, First International Conference on Evolutionary Multi-Criterion Optimization, pages 268–282. Springer-Verlag. Lecture Notes in Computer Science No. 1993, 2001.

[523] Shigenobu Kobayashi, Koji Yoshida, and Masayuki Asada. Generating a Set of Pareto Optimal Decision Trees by Genetic Algorithms. Journal of the Japanese Society for Artificial Intelligence, 11(5):725–732, September 1996.

[524] Thomas E. Koch and Andreas Zell. MOCS: Multi-Objective Clustering Selection Evolutionary Algorithm. In W.B. Langdon, E. Cantú-Paz, K. Mathias, R. Roy, D. Davis, R. Poli, K. Balakrishnan, V. Honavar, G. Rudolph, J. Wegener, L. Bull, M.A. Potter, A.C. Schultz, J.F. Miller, E. Burke, and N. Jonoska, editors, Proceedings of the Genetic and Evolutionary Computation Conference (GECCO’2002), pages 423–430, San Francisco, California, July 2002. Morgan Kaufmann Publishers.

[525] Ikeda Kokolo, Kita Hajime, and Kobayashi Shigenobu. Failure of Pareto-based MOEAs: Does Non-dominated Really Mean Near to Optimal? In Proceedings of the Congress on Evolutionary Computation 2001 (CEC’2001), volume 2, pages 957–962, Piscataway, New Jersey, May 2001. IEEE Service Center.

[526] Michael Kolonko and Stefan Voget. Multidimensional Optimization with Genetic Algorithms using Fuzzy Fitness Functions. In Tagungsband der Fuzzy-Neuro-Systeme, 1997.

[527] Mario Köppen and Stephan Rudlof. Multiobjective Optimization by Nessy Algorithm. In R. Roy, T. Furuhashi, and P.K. Chawdhry, editors, Advances in Soft Computing, pages 357–368, London, 1998. Springer.

[528] Urlich Korn and To Thanh Binh. The parallel evolution strategy toolbox. Automatisierungstechnik, 4:207–208, 1998.

[529] Matthias Krause and Volker Nissen. On Using Penalty Functions and Multicriteria Optimisation Techniques in Facility Layout. In J. Biethahn and Volker Nissen, editors, Evolutionary Algorithms in Management Applications, pages 153–166. Springer-Verlag, Berlin, 1995.

[530] Rajeev Kumar. On Generalisation of machine learning with neuralevolutionary computations. In Proceedings of the Third International Conference on Computational Intelligence and Multimedia Applications (IC-CIMA’99), pages 112–116, Los Alamitos, California, 1999. IEEE Computer Society Press.

[531] Rajeev Kumar. Codebook Design for Vector Quantisation using Multiobjective Genetic Algorithms. In PPSN/SAB Workshop on Multiobjective Problem Solving from Nature (MPSN), Paris, France, September 2000.

[532] Rajeev Kumar, N. Vinay Kumar, and I. J. Nagrath. Object oriented toolkit for multiobjective genetic optimisation. In 3rd International Conference on Computational Intelligence and Multimedia Applications, New Delhi, India, September 1999.

[533] Rajeev Kumar, Prajna P. Parida, and Mohit Gupta. Topological Design of Communication Networks using Multiobjective Genetic Optimization. In Congress on Evolutionary Computation (CEC’2002), volume 1, pages 425–430, Piscataway, New Jersey, May 2002. IEEE Service Center.

[534] Rajeev Kumar, S. Prasanth, and M.S. Sudarshan. Topological Design of Mesh Communication Networks using Multiobjecitve Genetic Optimisation. In PPSN/SAB Workshop on Multiobjective Problem Solving from Nature (MPSN), Paris, France, September 2000.

[535] Rajeev Kumar and Peter Rockett. Decomposition of High Dimensional Pattern Spaces for Hierarchical Classification. In Proceedings of the Workshop on Statistical Techniques in Pattern Recognition, Prague, Czech Republic, June 1997.

[536] Rajeev Kumar and Peter Rockett. Decomposition of High Dimensional Pattern Spaces for Hierarchical Classification. Kybernetika, 34(4):435–442, 1998.

[537] Rajeev Kumar and Peter Rockett. Multiobjective Genetic Algorithm Partitioning for Hierarchical Learning of High-Dimensional Pattern Spaces: A Learning-Follows-Decomposition Strategy. IEEE Transactions on Neural Networks, 9(5):822–830, 1998.

[538] Rajeev Kumar and Peter Rockett. Improved Sampling of the Pareto-Front in Multiobjective Genetic Optimizations by Steady-State Evolution: A Pareto Converging Genetic Algorithm. Evolutionary Computation, 10(3):283–314, Fall 2002.

[539] Rajeev Kumar and Peter I. Rockett. Assessing the Convergence of Rank-Based Multiobjective Genetic Algorithms. In Proceedings of the 2nd IEE/IEEE International Conference on Genetic Algorithms in Engineering Systems: Innovations and Applications (GALESIA’97), pages 19–23, Glasgow, Scotland, September 1997. IEE.

[540] Sujay V. Kumar and S. Ranji Ranjithan. Evaluation of the Constraint Method-Based Evolutionary Algorithm (CMEA) for a Tree-Objective Optimization Problem. In W.B. Langdon, E. Cantú-Paz, K. Mathias, R. Roy, D. Davis, R. Poli, K. Balakrishnan, V. Honavar, G. Rudolph, J. Wegener, L. Bull, M.A. Potter, A.C. Schultz, J.F. Miller, E. Burke, and N. Jonoska, editors, Proceedings of the Genetic and Evolutionary Computation Conference (GECCO’2002), pages 431–438, San Francisco, California, July 2002. Morgan Kaufmann Publishers.

[541] Sourav Kundu. A multicriteria genetic algorithm to solve optimization problems in structural engineering design. In B. Kumar, editor, Information Processing in Civil and Structural Engineering Design, pages 225–233, Glasgow, Scotland, August 1996. Civil-Comp Press Ltd.

[542] Sourav Kundu. A Note on Optimality vs. Stability—A Genetic Algorithm based Approach. In Proceedings of the Third World Congress of Structural and Multidisciplinary Optimization (WCSMO), Buffalo, New York, May 1999.

[543] Sourav Kundu, S. Kawata, and A. Watanabe. A multicriteria approach to control system design with genetic algorithm. In Proceedings of the International Federation of Automatic Control (IFAC’96)—13th World Congress, volume D, pages 315–320, Klidington, UK, 1996. Elsevier Science.

[544] Sourav Kundu and Seichi Kawata. AI in Control System Design Using a New Paradigm for Design Representation. In J. S. Gero and F. Sudweeks, editors, Artificial Intelligence in Design, pages 135–150. Kluwer Academic Publishers, The Netherlands, 1996.

[545] Sourav Kundu and Seichi Kawata. A GA-based state feedback design method using bicriterion performance index and tournament selection. In Proceedings of the Fifth International Conference on Intelligent Systems, pages 169–173, Reno, Nevada, 1996. International Society for Computers and Their Applications (ISCA).

[546] Sourav Kundu, Seichi Kawata, and A. Watanabe. Optimal control system design using a Pareto genetic algorithm. In Proceedings of the Joint System and Information ’95 Symposium, pages 53–59, Toyama, Japan, November 1995. SICE.

[547] Sourav Kundu and Seiichi Kawata. Evolutionary Multicriteria Optimization for Improved Design of Optimal Control Systems. In I.C. Parmee, editor, Proceedings of the Fifth International Conference on Adaptive Computing Design and Manufacture (ACDM 2002), volume 5, pages 207–218, University of Exeter, Devon, UK, April 2002. Springer-Verlag.

[548] Sourav Kundu and Andrzej Osyczka. The effect of genetic algorithm selection mechanisms on multicriteria optimization using the distance method. In Proceedings of the Fifth International Conference on Intelligent Systems, pages 164–168, Reno, Nevada, 1996. International Society for Computers and Their Applications (ISCA).

[549] Sourav Kundu and Andrzej Osyczka. Genetic multicriteria optimization of structural systems. In Proceedings of the 19th International Congress on Theoretical and Applied Mechanics (ICTAM 1996), page 272, Kyoto, Japan, August 1996. International Union of Theoretical and Applied Mechanics (IUTAM). Volumen of Abstracts.

[550] M.A. Kupinski and M.A. Anastasio. Multiobjective Genetic Optimization of Diagnostic Classifiers with Implications for Generating Receiver Operating Characteristic Curves. IEEE Transactions on Medical Imaging, 18(8):675– 685, August 1999.

[551] Setsuya Kurahashi and Takao Terano. A Genetic Algorithm with Tabu Search for Multimodal and Multiobjective Function Optimization. In Darrell Whitley, David Goldberg, Erick Cantú-Paz, Lee Spector, Ian Parmee, and Hans-Georg Beyer, editors, Proceedings of the Genetic and Evolutionary Computation Conference (GECCO’2000), pages 291–298, San Francisco, California, 2000. Morgan Kaufmann.

[552] A. Kurpati and S. Azarm. Immune Network Simulation with Multiobjective Genetic Algorithms for Multidisciplinary Design Optimization. Engineering Optimization, 33:245–260, 2000.

[553] A. Kurpati, S. Azarm, and J. Wu. Constraint handling improvements for multiobjective genetic algorithms. Structural and Multidisciplinary Optimization, 23(3):204–213, April 2002.

[554] Frank Kursawe. A Variant of Evolution Strategies for Vector Optimization. In H. P. Schwefel and R. Männer, editors, Parallel Problem Solving from Nature. 1st Workshop, PPSN I, volume 496 of Lecture Notes in Computer Science, pages 193–197, Berlin, Germany, oct 1991. Springer-Verlag.

[555] Frank Kursawe. Evolution strategies for vector optimization. In Preliminary Procedings of the Tenth International Conference on Multiple Criteria Decision Making, pages 187–193, Taipei, China, July 1992. National Chiao Tung University.

[556] Andrew Kusiak. Evolutionary Computation and Data Mining. In B. Gopalakrishnan and A. Gunasekaran, editors, Proceedings of the SPIE Conference on Intelligent Systems and Advances Manufacturing, pages 1–10, Boston, Massachussetts, 2000. SPIE.

[557] Michael Lahanas. Anatomy-based three-dimensional dose optimization in brachytherapy using multiobjective genetic algorithms. Medical Physics, 26(9):1904–1918, September 1999.

[558] Michael Lahanas, Natasa Milickovic, Dimos Baltas, and Nikolaos Zamboglou. Application of Multiobjective Evolutionary Algorithms for Dose Optimization Problems in Brachytherapy. In Eckart Zitzler, Kalyanmoy Deb, Lothar Thiele, Carlos A. Coello Coello, and David Corne, editors, First International Conference on Evolutionary Multi-Criterion Optimization, pages 574–587. Springer-Verlag. Lecture Notes in Computer Science No. 1993, 2001.

[559] W. B. Langdon. Data Structures and Genetic Programming. PhD thesis, University College, London, September 1996.

[560] W. B. Langdon. Scheduling Maintenance of Electrical Power Transmission Networks Using Genetic Programming. In The 1st Online Workshop on Soft Computing (WSC1). Research Group on ECOmp of the Society of Fuzzy Theory and Systems (SOFT), Nagoya University, Japan, August 1996.

[561] W. B. Langdon. Scheduling Planned Maintenance of Electrical Power Transmission Networks Using Genetic Algorithms. In Gennady K. Voronovsky and Serguey A. Sergeev, editors, Artificial Neural Networks and Genetic Algorithms in Power Engineering. OSNOVA, Ukraine, 1997. (in Russian).

[562] W. B. Langdon and P. C. Treleaven. Scheduling Maintenance of Electrical Power Transmission Networks Using Genetic Programming. In Kevin Warwick, Arthur Ekwue, and Raj Aggarwal, editors, Artificial Intelligence Techniques in Power Systems, chapter 10, pages 220–237. IEE, 1997.

[563] William B. Langdon. Data Structures and Genetic Programming. Research Note RN/95/70, University College London, Gower Street, London WC1E 6BT, UK, September 1995.

[564] William B. Langdon. Evolving data structures using genetic programming. In Larry Eshelman, editor, Proceedings of the Sixth International Conference on Genetic Algorithms (ICGA’95), pages 295–302, Pittsburgh, PA, July 1995. Morgan Kaufmann.

[565] William B. Langdon. Evolving data structures using genetic programming. Research Note RN/95/1, University College London, Gower Street, London WC1E 6BT, UK, January 1995.

[566] William B. Langdon. Pareto, Population Partitioning, Price and Genetic Programming. Research Note RN/95/29, University College London, Gower Street, London WC1E 6BT, UK, April 1995.

[567] William B. Langdon. Data structures and genetic programming. In Peter J. Angeline and Kenneth E. Kinnear, Jr., editors, Advances in Genetic Programming 2, chapter 20, pages 395–414. MIT Press, Cambridge, MA, USA, 1996.

[568] William B. Langdon. Scheduling Maintenance of Electrical Power Transmission Networks Using Genetic Programming. Research Note RN/96/49, University College London, Gower Street, London WC1E 6BT, UK, June 1996.

[569] William B. Langdon. Scheduling Maintenance of Electrical Power Transmission Networks Using Genetic Programming. In John R. Koza, editor, Late Breaking Papers at the GP-96 Conference, pages 107–116, Stanford, CA, USA, 28–31 July 1996. Stanford Bookstore.

[570] William B. Langdon. Using Data Structures within Genetic Programming. In John R. Koza, David E. Goldberg, David B. Fogel, and Rick L. Riolo, editors, Genetic Programming 1996: Proceedings of the First Annual Conference, pages 141–148, Stanford University, CA, USA, 28–31 July 1996. MIT Press.

[571] William B. Langdon. Using Data Structures within Genetic Programming. Research Note RN/96/1, University College London, Gower Street, London WC1E 6BT, UK, January 1996.

[572] William B. Langdon. Data Structures and Genetic Programming: Genetic Programming + Data Structures = Automatic Programming! Kluwer, Boston, April 1998.

[573] S.P. Larcombe, D.J. Prendergast, N.A. Thacker, and P.A. Ivey. Initial Development of a Genetic Algorithm to Automate System Implementation in a Novel Electronic Packaging Technology. In Proceedings of GALESIA’97, Glasgow, September 1997.

[574] S.P. Larcombe and N.A. Thacker. Using Genetic Algorithms to Automate System Implementation in a Novel Three-Dimensional Packaging Technology. In Proceedings of IEEE ICCD, 1996.

[575] Marco Laumanns, Günter Rudolph, and Hans-Paul Schwefel. A Spatial Predator-Prey Approach to Multi-Objective Optimization: A Preliminary Study. In A. E. Eiben, M. Schoenauer, and H.-P. Schwefel, editors, Parallel Problem Solving From Nature — PPSN V, pages 241–249, Amsterdam, Holland, 1998. Springer-Verlag.

[576] Marco Laumanns, Günter Rudolph, and Hans-Paul Schwefel. Approximating the Pareto Set: Concepts, Diversity Issues, and Performance Assessment. Technical Report CI-72/99, Dortmund: Department of Computer Science/LS11, University of Dortmund, Germany, March 1999. ISSN 1433-3325.

[577] Marco Laumanns, Günter Rudolph, and Hans-Paul Schwefel. Adaptive Mutation Control in Panmictic and Spatially Distributed Multi-Objective Evolutionary Algorithms. In PPSN/SAB Workshop on Multiobjective Problem Solving from Nature (MPSN), Paris, France, September 2000.

[578] Marco Laumanns, Günter Rudolph, and Hans-Paul Schwefel. Mutation Control and Convergence in Evolutionary Multi-Objective Optimization. In Proceedings of the 7th International Mendel Conference on Soft Computing (MENDEL 2001), Brno, Czech Republic, June 2001.

[579] Marco Laumanns, Lothar Thiele, Kalyanmoy Deb, and Eckart Zitzler. On the Convergence and Diversity-Preservation Properties of Multi-Objective Evolutionary Algorithms. Technical Report 108, Computer Engineering and Networks Laboratory (TIK), Swiss Federal Institute of Technology (ETH) Zurich, Gloriastrasse 35, CH-8092 Zurich, Switzerland, May 2001.

[580] Marco Laumanns, Lothar Thiele, Kalyanmoy Deb, and Eckart Zitzler. Combining convergence and diversity in evolutionary multi-objective optimization. Evolutionary Computation, 10(3):263–282, Fall 2002.

[581] Marco Laumanns, Lothar Thiele, Eckart Zitzler, and Kalyanmoy Deb. Archiving with Guaranteed Convergence and Diversity in Multi-Objective Optimization. In W.B. Langdon, E. Cantú-Paz, K. Mathias, R. Roy, D. Davis, R. Poli, K. Balakrishnan, V. Honavar, G. Rudolph, J. Wegener, L. Bull, M.A. Potter, A.C. Schultz, J.F. Miller, E. Burke, and N. Jonoska, editors, Proceedings of the Genetic and Evolutionary Computation Conference (GECCO’2002), pages 439–447, San Francisco, California, July 2002. Morgan Kaufmann Publishers.

[582] Marco Laumanns, Lothar Thiele, Eckart Zitzler, Emo Welzl, and Kalyanmoy Deb. Running Time Analysis of Multi-objective Evolutionary Algorithms on a Simple Discrete Optimization Problem. In Juan Julián Merelo Guervós, Panagiotis Adamidis, Hans-Georg Beyer, José-Luis Fernández-Villaca nas, and Hans-Paul Schwefel, editors, Parallel Problem Solving from Nature—PPSN VII, pages 44–53, Granada, Spain, September 2002. Springer-Verlag. Lecture Notes in Computer Science No. 2439.

[583] Marco Laumanns, Eckart Zitzler, and Lothar Thiele. A Unified Model for Multi-Objective Evolutionary Algorithms with Elitism. In 2000 Congress on Evolutionary Computation, volume 1, pages 46–53, Piscataway, New Jersey, July 2000. IEEE Service Center.

[584] Marco Laumanns, Eckart Zitzler, and Lothar Thiele. On the Effects of Archiving, Elitism, and Density Based Selection in Evolutionary Multi-objective Optimization. In Eckart Zitzler, Kalyanmoy Deb, Lothar Thiele, Carlos A. Coello Coello, and David Corne, editors, First International Conference on Evolutionary Multi-Criterion Optimization, pages 181–196. Springer-Verlag. Lecture Notes in Computer Science No. 1993, 2001.

[585] Marco Laumanns, Eckart Zitzler, and Lother Thiele. Multiple Criteria Decision Support by Evolutionary Computation. In L.M. Hilty and P.W. Gilgen, editors, Sustainability in the Information Society, 15th International Symposium Informatics for Environmental Protection, Zurich, October 2001. Verlag.

[586] Nando Laumanns, Marco Laumanns, and Harmut Kitterer. Evolutionary Multi-Objective Integer Programming for the Design of Adaptive Cruise Control Systems. In Tim Hendtlass and Moonis Ali, editors, Proceedings of the Fifteenth International Conference on Industrial & Engineering Applications of Artificial Intelligence & Expert Systems (IEA/AIE-2002), pages 200–210, Cairns, Australia, June 2002. Springer-Verlag. Lecture Notes in Artificial Intelligence Vol. 2358.

[587] Nando Laumanns, Marco Laumanns, and Dirk Neunzig. Multi-objective Design Space Exploration of Road Trains with Evolutionary Algorithms. In Eckart Zitzler, Kalyanmoy Deb, Lothar Thiele, Carlos A. Coello Coello, and David Corne, editors, First International Conference on Evolutionary Multi-Criterion Optimization, pages 612–623. Springer-Verlag. Lecture Notes in Computer Science No. 1993, 2001.

[588] Dongkon Lee. Multiobjective Design of a Marine Vehicle with Aid of Design Knowledge. International Journal for Numerical Methods in Engineering, 40:2665–2677, 1997.

[589] Jongsoo Lee and Prabhat Hajela. Parallel Genetic Algorithm Implementation in Multidisciplinary Rotor Blade Design. Journal of Aircraft, 33(5):962–969, September-October 1996.

[590] M. A. Lee and H. Esbensen. Constructing Fuzzy/Evolutionary Multiobjective Optimization Algorithms. In Proceedings of the International IEEE Conference on Fuzzy System (FUZZ-IEEE96), New Orleans, Louisiana, 1996.

[591] M. A. Lee and H. Esbensen. Multiobjective Optimization using Fuzzy/Evolutionary Algorithms. In Proceedings of the International Society for Computers and Their Applications (ISCA96), San Francisco, California, 1996.

[592] M. A. Lee and H. Esbensen. Set Quality Measures for Characterizing Multiobjective Optimization Algorithm Behavior. In Procedings of the North American Fuzzy Information Processing Society (NAFIPS96), Berkeley, California, 1996.

[593] M. A. Lee and H. Esbensen. Fuzzy/Multiobjective Genetic Systems for Intelligent Systems Design Tools and Components. In Witold Pedrycz, editor, Fuzzy Evolutionary Computation, pages 57–80. Kluwer Academic Publishers, Boston, Massachusetts, 1997.

[594] M. A. Lee and R. Hartani. A Multiobjective Evolutionary Algorithms Approach to Fuzzy Modelling. In Proceedings of the Second Annual Joint Conference on Information Sciences (JCIS’95), pages 460–463, Wrightsville Beach, North Carolina, 1995.

[595] Michael A. Lee, Henrik Esbensen, and Laurent Lemaitre. The Design of Hybrid Fuzzy/Evolutionary Multiobjective Optimization Algorithms. In Proceedings of the 1995 IEEE/Nagoya University World Wiseperson Workshop, pages 118–125, Nagoya, Japan, 1995.

[596] T.H. Lee, K.C. Tan, and E.F. Khor. Control System Design Unification and Automation—A Way Forward in CACSD via Evolutionary Computation. In Proceedings of the Congress on Evolutionary Computation 2001 (CEC’2001), volume 1, pages 152–161, Piscataway, New Jersey, May 2001. IEEE Service Center.

[597] Yann Lefablec. Optimisation par algorithmes génétiques parallèles et multiobjectifs. Thesis, 1995. (in French).

[598] Patrick Christopher Leger. Automated Synthesis and Optimization of Robot Configurations: An Evolutionary Approach. PhD thesis, The Robotics Institute, Carnegie Mellon University, Pittsburgh, Pennsylvania, December 1999.

[599] Héctor A. Leiva, Susana C. Esquivel, and Raúl H. Gallard. Multiplicity and Local Search in Evolutionary Algorithms to Build the Pareto Front. In Proceedings of the XX International Conference of the Chilean Computer Science Society, pages 7–13, Piscataway, New Jersey, 2000. IEEE Computer Society Press.

[600] Yiu-Wing Leung and Yuping Wang. Multiobjective Programming Using Uniform Design and Genetic Algorithm. IEEE Transactions on Systems, Man, and Cybernetics—Part C: Applications and Reviews, 30(3):293–304, August 2000.

[601] Geoff Leyland. Multi-Objective Optimisation Applied to Industrial Energy Problems. PhD thesis, École Polytechnique Fédérale de Lausanne, Lausanne, Switzerland, May 2002.

[602] J.C. Leyva and E. Fernández. A genetic algorithm for deriving final ranking from a fuzzy outranking relation. Foundations of Computing and Decision Sciences, 24(1):33–47, 1999.

[603] Xiaodong Li, Tianzi Jiang, and D. J. Evans. Medical Image Reconstruction Using a Multi-objective Genetic Local Search Algorithm. International Journal on Computer Mathematics, 74(301–314), 2000.

[604] Yinzhen Li, Mitsuo Gen, and Kenichi Ida. Evolutionary Computation for Multicriteria Solid Transportation Problem with Fuzzy Numbers. In Thomas Bäck, Zbigniew Michalewicz, and Hiroaki Kitano, editors, Proceedings of the Third IEEE Conference on Evolutionary Computation, pages 596–601, Piscataway, New Jersey, 1996. IEEE Service Center.

[605] Yinzhen Li, Kenichi Ida, and Mitsuo Gen. Improved Genetic Algorithm for Solving Multiobjective Solid Transportation Problem with Fuzzy Numbers. Computers in Industrial Engineering, 33(3-4):589–592, 1997.

[606] Simon J. Liang and John M. Lewis. Job Shop Scheduling Using Multiple Criteria. In Proceedings of the Joint Hungarian-British Mechatronic Conference, pages 77–82. Computational Mechanics, September 1994.

[607] G. E. Liepins, M. R. Hilliard, J. Richardson, and M. Palmer. Genetic algorithms application to set covering and travelling salesman problems. In D. E. Brown and C. C. White, editors, Operations research and Artificial Intelligence: The integration of problem-solving strategies, pages 29–57. Kluwer Academic, Norwell, Massachusetts, 1990.

[608] Dan Lin, Shouyang Wang, and Hong Yan. A multiobjective genetic algorithm for portfolio selection. Working Paper, Institute of Systems Science, Academy of Mathematics and Systems Science Chinese Academy of Sciences, Beijing, China, 2001.

[609] Shie-Yui Liong, Soon-Thiam Khu, and Weng Tat Chan. Derivation of Pareto Front with Accelerated Convergence Genetic Algortithm (ACGA). In V. Babovic and L. C. Larsen, editors, Proceedings of the Third Hydroinformatics Conference, 1998.

[610] Shie-Yui Liong, Soon-Thiam Khu, and Weng Tat Chan. Novel Application of Genetic Algorithm and Neural Network in Water Resources: Development of Pareto Front. In Eleventh Congress of the International Association for Hydraulic Research—Asia and Pacific Division, pages 185–194, Yogyakarta, Indonesia, 1998.

[611] Joanna Lis and A. E. Eiben. A Multi-Sexual Genetic Algorithm for Multiobjective Optimization. In Toshio Fukuda and Takeshi Furuhashi, editors, Proceedings of the 1996 International Conference on Evolutionary Computation, pages 59–64, Nagoya, Japan, 1996. IEEE.

[612] Chunlu Liu, Amin Hammad, and Yoshito Itoh. Multiobjective Optimization of Bridge Deck Rehabilitation Using a Genetic Algorithm. Journal of Computer-Aided Civil and Infrastructure Engineering, 12(3):431–443, November 1997.

[613] Juan Liu and Hitoshi Iba. Selecting Informative Genes Using a Multiobjective Evolutionary Algorithm. In Congress on Evolutionary Computation (CEC’2002), volume 1, pages 297–302, Piscataway, New Jersey, May 2002. IEEE Service Center.

[614] Tung-Kuan Liu, Tadashi Ishihara, and Hikaru Inooka. Multiobjective Control Systems Design by Genetic Algorithms. In Proceedings of the 34th Society of Instrument and Control Engineers Annual Conference, pages 1521–1526, 1995.

[615] Tung-Kuan Liu, Tadashi Ishihara, and Hikaru Inooka. Design of Discrete-Time Control Systems by Multiobjective Genetic Algorithms. In Toshio Fukuda and Takeshi Furuhashi, editors, Proceedings of the 1996 IEEE IECON 22nd International Conference on Industrial Electronics, Control, and Instrumentation, pages 1618–1623. IEEE, 1996.

[616] Xiaojian Liu, D. W. Begg, and R. J. Fishwick. Genetic approach to optimal topology/controller design of adaptive structures. International Journal for Numerical Methods in Engineering, 41:815–830, 1998.

[617] Chi-Chun Lo and Wei-Hsin Chang. A Multiobjective Hybrid Genetic Algorithm for Capacitated Multipoint Network Design Problem. IEEE Transactions on Systems, Man, and Cybernetics Part B: Cybernetics, 30(3):461–470, June 2000.

[618] Jason D. Lohn, William F. Kraus, and Gary L. Haith. Comparing a Coevolutionary Genetic Algorithm for Multiobjective Optimization. In Congress on Evolutionary Computation (CEC’2002), volume 2, pages 1157–1162, Piscataway, New Jersey, May 2002. IEEE Service Center.

[619] Alberto Herreros López. Diseño de Controladores Robustos Multiobjetivo por Medio de Algoritmos Genéticos. PhD thesis, Departamento de Ingeniería de Sistemas y Automática, Universidad de Valladolid, Valladolid, España, Septiembre 2000. (In Spanish).

[620] Daniel H. Loughlin and S. Ranjithan. The Neighborhood constraint method: A Genetic Algorithm-Based Multiobjective Optimization Technique. In Thomas Bäck, editor, Proceedings of the Seventh International Conference on Genetic Algorithms, pages 666–673, San Mateo, California, July 1997. Michigan State University, Morgan Kaufmann Publishers.

[621] Daniel H. Loughlin, S. Ranji Ranjithan, E. Downey Brill Jr., and John W. Baugh Jr. Genetic Algorithm Approaches for Addressing Unmodeled Objectives in Optimization Problems. Engineering Optimization, 33:549–569, 2001.

[622] Daniel Hopkins Loughlin. Genetic Algorithm-Based Optimization in the Development of Tropospheric Ozone Control Strategies: Least Cost, Multiobjective, Alternative Generation, and Chance-Constrained Applications (Air Quality Management). PhD thesis, North Carolina State University, February 1998.

[623] Sushil J. Louis. Genetic Algorithms as a Computational Tool for Design. PhD thesis, Department of Computer Science, Indiana University, aug 1993.

[624] Sushil J. Louis and Gregory J. E. Rawlins. Pareto Optimality, GA-easiness and Deception. In Stephanie Forrest, editor, Proceedings of the Fifth International Conference on Genetic Algorithms, pages 118–123, University of Illinois at Urbana-Champaign, 1993. Morgan Kaufmann Publishers.

[625] Haiming Lu and Gary G. Yen. Multiobjective Optimization Design via Genetic Algorithm. In Proceedings of the 2001 International Conference on Control Applications, pages 1190–1195. IEEE, 2001.

[626] Haiming Lu and Gary G. Yen. Dynamic Population Size in Multiobjective Evolutionary Algorithms. In Congress on Evolutionary Computation (CEC’2002), volume 2, pages 1648–1653, Piscataway, New Jersey, May 2002. IEEE Service Center.

[627] Haiming Lu and Gary G. Yen. Rank-Density Based Multiobjective Genetic Algorithm. In Congress on Evolutionary Computation (CEC’2002), volume 1, pages 944–949, Piscataway, New Jersey, May 2002. IEEE Service Center.

[628] Sean Luke and Liviu Panait. Lexicographic parsimony pressure. In W.B. Langdon, E. Cantú-Paz, K. Mathias, R. Roy, D. Davis, R. Poli, K. Balakrishnan, V. Honavar, G. Rudolph, J. Wegener, L. Bull, M.A. Potter, A.C. Schultz, J.F. Miller, E. Burke, and N. Jonoska, editors, Proceedings of the Genetic and Evolutionary Computation Conference (GECCO’2002), pages 829–836, San Francisco, California, July 2002. Morgan Kaufmann Publishers.

[629] Panta Lučić and Dušan Teodorović. Simulated annealing for the multiobjective aircrew rostering problem. Transportation Research Part A, 33:19– 45, 1999.

[630] Mohammed Hakim Mabed, Malek Rahoual, El-Ghazali Talbi, and Clarisse Dhaenens. Using Genetic Algorithms to schedule multicriteria Flow-shop. In PPSN/SAB Workshop on Multiobjective Problem Solving from Nature (MPSN), Paris, France, September 2000.

[631] Nateri K. Madavan. Multiobjective Optimization Using a Pareto Differential Evolution Approach. In Congress on Evolutionary Computation (CEC’2002), volume 2, pages 1145–1150, Piscataway, New Jersey, May 2002. IEEE Service Center.

[632] M. Mahfouf, M. F. Abbod, and D. A. Linkens. Multi-Objective Genetic Optimization of the Performance Index of Self-Organizing Fuzzy Logic Control Algorithm Using a Fuzzy Ranking Approach. In H. J. Zimmerman, editor, Proceedings of the Sixth European Congress on Intelligent Techniques and Soft Computing, pages 1799–1808, Aachen, 1998. Verlag Mainz.

[633] R. Mäkinen, P. Neittaanmäki, J. Periaux, M. Sefrioui, and J. Toivanen. Parallel genetic solution for multiobjective MDO. In A. Schiano, A. Ecer, J. Périaux, and N. Satofuka, editors, Parallel CFD’96 Conference, pages 352–359, Capri, 1996. Elsevier.

[634] R. Mäkinen, P. Neittaanmäki, J. Périaux, and J. Toivanen. A genetic Algorithm for Multiobjective Design Optimization in Aerodynamics and Electromagnetics. In K. D. Papailiou et al., editor, Computational Fluid Dynamics ’98, Proceedings of the ECCOMAS 98 Conference, volume 2, pages 418–422, Athens, Greece, September 1998. Wiley.

[635] Raino A.E. Mäkinen, Jacques Periaux, and Jari Toivanen. Multidisciplinary shape optimization in aerodynamics and electromagnetics using genetic algorithms. International Journal for Numerical Methods in Fluids, 30(2):149– 159, May 1999.

[636] K. F. Man, K. S. Tang, and S. Kwong. Genetic Algorithms. Concepts and Designs. Springer-Verlag, New York, second edition, 1999.

[637] S. Afshin Mansouri. Manufacturing Cell Design in a Multi-Criterion Environment. PhD thesis, Amirkabir University of Technology, Tehran, Iran, January 2001.

[638] S.A. Mansouri, S.M. Moattar-Husseini, and S.T. Newman. A review of the modern approaches to multi-criteria cell design. International Journal of Production Research, 38(5):1201–1218, 2000.

[639] S.A. Mansouri, S.M. Moattar-Husseini, and S.H. Zegordi. Multi-criterion Tackling Bottleneck Machines and Exceptional Parts in Cell Formation Using Genetic Algorithms. In I.C. Parmee, editor, Proceedings of the Fifth International Conference on Adaptive Computing Design and Manufacture (ACDM 2002), volume 5, pages 181–192, University of Exeter, Devon, UK, April 2002. Springer-Verlag.

[640] Jiangming Mao, Kotaro Hirasawa, Jinglu Hu, and Junichi Murata. Genetic symbiosis algorithm for multiobjective optimization problem. In Proceedings of the 9th IEEE International Workshop on Robot and Human Interactive Communication (RO-MAN 2000), pages 137–142. IEEE, 2000.

[641] Jiangming Mao, Kotaro Hirasawa, Jinglu Hu, and Junichi Murata. Genetic Symbiosis Algorithm for Multiobjective Optimization Problems. In Proceedings of the 2001 Genetic and Evolutionary Computation Conference. Late-Breaking Papers, pages 267–274, San Francisco, California, July 2001.

[642] Jiangming Mao, Kotaro Hirasawa, Jinglu Hu, and Junichi Murata. Genetic Symbiosis Algorithm for Multiobjective Optimization Problems. In Lee Spector, Erik D. Goodman, Annie Wu, W.B. Langdon, Hans-Michael Voigt, Mitsuo Gen, Sandip Sen, Marco Dorigo, Shahram Pezeshk, Max H. Garzon, and Edmund Burke, editors, Proceedings of the Genetic and Evolutionary Computation Conference (GECCO’2001), page 771, San Francisco, California, July 2001. Morgan Kaufmann Publishers.

[643] N. Marco, S. Lanteri, J.-A. Desideri, and J. Périaux. A Parallel Genetic Algorithm for Multi-Objective Optimization in Computational Fluid Dynamics. In Kaisa Miettinen, Marko M. Mäkelä, Pekka Neittaanmäki, and Jacques Périaux, editors, Evolutionary Algorithms in Engineering and Computer Science, chapter 22, pages 445–456. John Wiley & Sons, Ltd, Chichester, UK, 1999.

[644] T. Marcu, L. Ferariu, and P. M. Frank. Genetic Evolving of Dynamic Neural Networks with Application to Process Fault Diagnosis. In Procedings of the EUCA/IFAC/IEEE European Control Conference ECC’99, Karlsruhe, Germany, 1999. CD-ROM, F-1046,1.

[645] Teodor Marcu. A multiobjective evolutionary approach to pattern recognition for robust diagnosis of process faults. In R. J. Patton and J. Chen, editors, IFAC Symposium on Fault Detection, Supervision and Safety for Technical Processes: SAFEPROCESS’97, pages 1183–1188, Kington Upon Hull, United Kingdom, August 1997.

[646] Teodor Marcu and Paul M. Frank. Parallel Evolutionary Approach to System Identification for Process Fault Diagnosis. In Prasad S. Dhurjati and Sylvie Cauvin, editors, Procedings of the IFAC Workshop on ‘On-line Fault Detection and Supervision in the Chemical Process Industries’, pages 113–118, Solaize (Lyon), France, 1998.

[647] Simon Mardle, Sean Pascoe, and Mehrdad Tamiz. An investigation of genetic algorithms for the optimisation of multi-objective fisheries bioeconomic models. Technical Report 136, Centre for the Economics and Management of Aquatic Resources, University of Portsmouth, 1998.

[648] Simon Mardle, Sean Pascoe, and Mehrdad Tamiz. An Investigation of Genetic Algorithms for the Optimization of Multiobjective Fisheries Bioeconomic Models. In Proceedings of the Third International Conference on Multi-Objective Programming and Goal Programming: Theory and Applications (MOPGP’98), Quebec City, Canada, 1998.

[649] Simon Mardle, Sean Pascoe, and Mehrdad Tamiz. An investigation of genetic algorithms for the optimisation of multi-objective fisheries bioeconomic models. International Transactions of Operations Research, 7(1):33–49, 2000.

[650] Richard Marett and Mike Wright. A Comparison of Neighborhood Search Techniques for Multi-Objective Combinatorial Problems. Computers and Operations Research, 23(5):465–483, 1996.

[651] Carlos E. Mariano and Eduardo Morales. A Multiple Objective Ant-Q Algorithm for the Design of Water Distribution Irrigation Networks. Technical Report HC-9904, Instituto Mexicano de Tecnología del Agua, June 1999.

[652] Carlos E. Mariano and Eduardo Morales. A New Distributed Reinforcement Learning Algorithm for Multiple Objective Optimization Problems. Technical Report HC-200001, Instituto Mexicano de Tecnología del Agua, January 2000.

[653] Carlos E. Mariano and Eduardo F. Morales. Distributed Reinforcement Learning for Multiple Objective Optimization Problems. In 2000 Congress on Evolutionary Computation, volume 1, pages 188–195, Piscataway, New Jersey, July 2000. IEEE Service Center.

[654] Carlos E. Mariano and Eduardo F. Morales. MDQL: A Reinforcement Learning Approach for the Solution of Multiple Objective Optimization Problems. In PPSN/SAB Workshop on Multiobjective Problem Solving from Nature (MPSN), Paris, France, September 2000.

[655] Nick Marvin, Mark Bower, and Jonathan E. Rowe. An evolutionary approach to constructing prognostic models. Artificial Intelligence in Medicine, 15(2):155–165, February 1999.

[656] W. Mason, V. Coverstone-Carroll, and J. Hartmann. Optimal Earth Orbiting Satellite Constellations via a Pareto Genetic Algorithm. In 1998 AIAA/AAS Astrodynamics Specialist Conference and Exhibit, pages 169–177, Boston, Massachusetts, August 1998. Paper No. AIAA 98-4381.

[657] William J. Mason. Satellite Constellation Design Via Evolutionary Computation. Master’s thesis, Department of Aeronautical and Astronautical Engineering, University of Illinois at Urbana Champaign, December 2000. (In process).

[658] Silvère Massebeuf, Christian Fonteix, Laszlo N. Kiss, Ivan Marc, Fernand Pla, and Kazimierz Zaras. Multicriteria Optimization and Decision Engineering of an Extrusion Process Aided by a Diploid Genetic Algorithm. In 1999 Congress on Evolutionary Computation, pages 14–21, Washington, D.C., July 1999. IEEE Service Center.

[659] Keith B. Matthews, Susan Craw, Stewart Elder, Alan R. Sibbald, and Ian MacKenzie. Applying Genetic Algorithms to Multi-Objective Land Use Planning. In Darrell Whitley, David Goldberg, Erick Cantú-Paz, Lee Spector, Ian Parmee, and Hans-Georg Beyer, editors, Proceedings of the Genetic and Evolutionary Computation Conference (GECCO’2000), pages 613–620, San Francisco, California, 2000. Morgan Kaufmann.

[660] Patrick R. McMullen. An ant colony optimization approach to addessing a JIT sequencing problem with multiple objectives. Artificial Intelligence in Engineering, 15:309–317, 2001.

[661] F. Menczer, W.N. Street, and M. Degeratu. Evolving heterogeneous neural agents by local selection. In V. Honavar, M. Patel, and K. Balakrishnan, editors, Advances in the Evolutionary Synthesis of Neural Systems. MIT Press, Cambridge, MA, 2000.

[662] Filippo Menczer, Melania Degeratu, and W. Nick Street. Efficient and Scalable Pareto Optimization by Evolutionary Local Selection Algorithms. Evolutionary Computation, 8(2):223–247, Summer 2000.

[663] G. Meneghetti, V. Pediroda, and C. Poloni. Application of a Multi Objective Genetic Algorithm and a Neural Network to the Optimisation of Foundry Processes. In Kaisa Miettinen, Marko M. Mäkelä, Pekka Neittaanmäki, and Jacques Périaux, editors, Evolutionary Algorithms in Engineering and Computer Science, chapter 23, pages 457–470. John Wiley & Sons, Ltd, Chichester, UK, 1999.

[664] Herve Meunier, El-Ghazali Talbi, and Philippe Reininger. A Multiobjective Genetic Algorithm for Radio Network Optimization. In 2000 Congress on Evolutionary Computation, volume 1, pages 317–324, Piscataway, New Jersey, July 2000. IEEE Service Center.

[665] Efrén Mezura Montes. Uso de la Técnica Multiobjetivo NPGA para el Manejo de Restricciones en Algoritmos Genéticos. Master’s thesis, Maestría en Inteligencia Artificial, Universidad Veracruzana, Xalapa, Veracruz, México, August 2001. (In Spanish).

[666] X. R. Miao, P. M. Zhang, Z. H. Xu, and C. H. Chen. Multi-objective Genetic Algorithm in IHSACC Optimization Design. In Artificial Intelligence and Soft Computing (ASC’99), Honolulu, Hawaii, August 1999. ACTA Press.

[667] Zbigniew Michalewicz. Genetic Algorithms + Data Structures = Evolution Programs. Springer-Verlag, third edition, 1996.

[668] E. Michielssen, J. M. Sajer, and R. Mittra. Pareto-optimal design of broadband microwave absorbers using genetic algorithms. In Proceedings of the IEEE Antennas and Propagation Society International Symposium, volume 3, pages 1167–1170, Ann Arbor, Michigan, June 1993. IEEE.

[669] E. Michielssen and D. S. Weile. Electromagnetic System Design using Genetic Algorithms. In Genetic Algorithms and Evolution Strategies in Engineering and Computer Science, pages 267–288. John Wiley and Sons, England, 1995.

[670] Eric Michielssen, Jean-Michel Sajer, S. Ranjithan, and Raj Mittra. Design of Lightweight, Broad-Band Microwave Absorbers Using Genetic Algorithms. IEEE Transactions on Microwave Theory and Techniques, 41(6/7):1024–1031, 1993.

[671] Natasa Milickovic, Michael Lahanas, Dimos Baltas, and Nikolaos Zamboglou. Comparison of Evolutionary and Deterministic Multiobjective Algorithms for Dose Optimization in Branchytherapy. In Eckart Zitzler, Kalyanmoy Deb, Lothar Thiele, Carlos A. Coello Coello, and David Corne, editors, First International Conference on Evolutionary Multi-Criterion Optimization, pages 167–180. Springer-Verlag. Lecture Notes in Computer Science No. 1993, 2001.

[672] George G. Mitchell, Diarmuid O’Donoghue, and Adrian Trenaman. A New Operator for Efficient Evolutionary Solutions to the Travelling Salesman Problem. In IASTED. Applied Informatics, pages 98–103, Innsbruck, Austria, 2000.

[673] K. Mitra, Kalyanmoy Deb, and Santosh K. Gupta. Multiobjective Dynamic Optimization of an Industrial Nylon 6 Semibatch Reactor Using Genetic Algorithm. Journal of Applied Polymer Science, 69(1):69–87, 1998.

[674] Hajime Kita Mitsuhiro Shibuya and Shigenobu Kobayashi. Integration of multi-objective and interactive genetic algorithms and its application to animation design. In Proceedings of IEEE Systems, Man, and Cybernetics, volume III, pages 646–651, 1999.

[675] O. A. Mohammed and G. F. Üler. Genetic Algorithms for the Optimal Design of Electromagnetic Devices. In Conference on the Annual Review of Progress in Applied Computational Electromagnetics, volume 11, pages 386–393, 1995.

[676] A.K. Molyneaux, G.B. Leyland, and D.Favrat. A New, Clustering Evolutionary Multi-Objective Optimisation Technique. In Proceedings of the Third International Symposium on Adaptive Systems—Evolutionary Computation and Probabilistic Graphical Models, pages 41–47, Havana, Cuba, March 19–23 2001. Institute of Cybernetics, Mathematics and Physics.

[677] David Montana, Garrett Bidwell, Gordon Vidaver, and Jose Herrero. Scheduling and Route Selection for Military Land Moves Using Genetic Algorithms. In 1999 Congress on Evolutionary Computation, volume 2, pages 1118–1123, Washington, D.C., July 1999. IEEE Service Center.

[678] David Montana, Marshall Brinn, Sean Moore, and Garrett Bidwell. Genetic Algorithms for Complex, Real-Time Scheduling. In Proceedings of the 1998 IEEE International Conference on Systems, Man, and Cybernetics, pages 2213–2218, La Jolla, California, October 1998. IEEE.

[679] Chiung Moon, Yin-Zhen Lin, and Mitsuo Gen. Evolutionary Algorithm for Flexible Process Sequencing with Multiple Objectives. In David B. Fogel, editor, Proceedings of the 1998 International Conference on Evolutionary Computation, pages 27–32, Piscataway, New Jersey, 1998. IEEE.

[680] Jacqueline Moore and Richard Chapman. Application of Particle Swarm to Multiobjective Optimization. Department of Computer Science and Software Engineering, Auburn University. (Unpublished manuscript), 1999.

[681] Naoki Mori, Junji Yoshida, Hisashi Tamaki, Hajime Kita, and Yoshikazu Nishikawa. Thermodynamical Selection Rule for the Genetic Algorithm. In David B. Fogel, editor, Proceedings of the Second IEEE Conference on Evolutionary Computation, pages 188–192, Piscataway, New Jersey, 1995. IEEE Service Center.

[682] Sanaz Mostaghim, Jürgen Teich, and Ambrish Tyagi. Comparison of Data Structures for Storing Pareto-sets in MOEAs. In Congress on Evolutionary Computation (CEC’2002), volume 1, pages 843–848, Piscataway, New Jersey, May 2002. IEEE Service Center.

[683] Walid El Moudani, Carlos Alberto Nunes Cosenza, Marc de Coligny, and Félix Mora-Camino. A Bi-Criterion Approach for the Airlines Crew Rostering Problem. In Eckart Zitzler, Kalyanmoy Deb, Lothar Thiele, Carlos A. Coello Coello, and David Corne, editors, First International Conference on Evolutionary Multi-Criterion Optimization, pages 486–500. Springer-Verlag. Lecture Notes in Computer Science No. 1993, 2001.

[684] Silla Mullei and Peter Beling. Hybrid Evolutionary Algorithms for a Multiobjective Financial Problem. In Proceedings of the 1998 IEEE International Conference on Systems, Man, and Cybernetics, volume 4, pages 3925–3930. IEEE, October 1998.

[685] Sibylle D. Müller, Ivo F. Sbalzarini, Jens H. Walther, and Petros D. Koumoutsakos. Evolution Strategies for the Optimization of Microdevices. In Proceedings of the Congress on Evolutionary Computation 2001 (CEC’2001), volume 1, pages 302–309, Piscataway, New Jersey, May 2001. IEEE Service Center.

[686] Tadahiko Murata. Genetic Algortithms for Multi-Objective Optimization. PhD thesis, Osaka Prefecture University, Japan, 1997.

[687] Tadahiko Murata and Hisao Ishibuchi. MOGA: Multi-Objective Genetic Algorithms. In Proceedings of the 2nd IEEE International Conference on Evolutionary Computing, pages 289–294, Perth, Australia, November 1995.

[688] Tadahiko Murata and Hisao Ishibuchi. Application of Two-Objective Genetic Algorithm to Flowshop Scheduling Problems with Interval Processing Time. In Proceedings of EUFIT’96, pages 443–447, Aachen, Germany, September 1996.

[689] Tadahiko Murata and Hisao Ishibuchi. Performance of Multi-Objective Genetic Algorithms for Flowshop Scheduling Problems. In Proceedings of the 14th International Conference on Production Research, pages 498–501, Osaka, Japan, August 1997.

[690] Tadahiko Murata and Hisao Ishibuchi. Constructing Multi-Objective Genetic Local Search Algorithms for Multi-Objective Flowshop Scheduling Problems. In Proceedings of the 1998 Japan-USA Symposium on Flexible Automation, pages 1353–1356, Ohtsu, Japan, July 1998.

[691] Tadahiko Murata, Hisao Ishibuchi, and Mitsuo Gen. Random Weights in Multi-Objective Genetic Algorithms. In Proceedings of the 2nd International Conference on Engineering Design and Automation, Maui, Hawaii, August 1998. Only CD-ROM Proceedings available.

[692] Tadahiko Murata, Hisao Ishibuchi, and Mitsuo Gen. Specification of Local Search Directions in Genetic Local Search Algorithms for Multi-Objective Optimization Problems. In W. Banzhaf, J. Daida, A. E. Eiben, M. H. Garzon, V. Honavar, M. Jakiela, and R. E. Smith, editors, Proceedings of the Genetic and Evolutionary Computation Conference (GECCO’99), volume 1, pages 441–448, San Francisco, California, July 1999. Morgan Kaufmann.

[693] Tadahiko Murata, Hisao Ishibuchi, and Mitsuo Gen. Cellular Genetic Local Search for Multi-Objective Optimization. In Darrell Whitley, David Goldberg, Erick Cantú-Paz, Lee Spector, Ian Parmee, and Hans-Georg Beyer, editors, Proceedings of the Genetic and Evolutionary Computation Conference (GECCO’2000), pages 307–314, San Francisco, California, 2000. Morgan Kaufmann.

[694] Tadahiko Murata, Hisao Ishibuchi, and Mitsuo Gen. Specification of Genetic Search Directions in Cellular Multi-objective Genetic Algorithms. In Eckart Zitzler, Kalyanmoy Deb, Lothar Thiele, Carlos A. Coello Coello, and David Corne, editors, First International Conference on Evolutionary Multi-Criterion Optimization, pages 82–95. Springer-Verlag. Lecture Notes in Computer Science No. 1993, 2001.

[695] Tadahiko Murata, Hisao Ishibuchi, and H. Tanaka. Multi-Objective Genetic Algorithm and Its Application to Flowshop Scheduling. Computers and Industrial Engineering, 30(4):957–968, September 1996.

[696] Tadahiko Murata, Shihei Kawakami, Hiroyuki Nozawa, Mitsuo Gen, and Hisao Ishibuchi. Three-Objective Genetic Algorithms for Designing Compact Fuzzy Rule-Based Systems for Pattern Classification Problems. In Lee Spector, Erik D. Goodman, Annie Wu, W.B. Langdon, Hans-Michael Voigt, Mitsuo Gen, Sandip Sen, Marco Dorigo, Shahram Pezeshk, Max H. Garzon, and Edmund Burke, editors, Proceedings of the Genetic and Evolutionary Computation Conference (GECCO’2001), pages 485–492, San Francisco, California, 2001. Morgan Kaufmann Publishers.

[697] Tadahiko Murata, Hiroyuki Nozawa, Yasuhiro Tsujimura, Mitsuo Gen, and Hisao Ishibuchi. Effect of Local Search on the Performance of Cellular Multi-Objective Genetic Algorithms for Designing Fuzzy Rule-based Classification Systems. In Congress on Evolutionary Computation (CEC’2002), volume 1, pages 663–668, Piscataway, New Jersey, May 2002. IEEE Service Center.

[698] Karl Nachtigall and Stefan Voget. Minimizing Waiting Times in Integrated Fixed Interval Timetables by Upgrading Railway Tracks. European Journal of Operational Research, 103:610–627, 1997.

[699] Prasanth B. Nair and Andrew J. Keane. Coevolutionary architecture for distributed optimization of complex coupled systems. AIAA Journal, 40(7):1434– 1443, July 2002.

[700] Hirotaka Nakayama, Masao Arakawa, and Ye Boon Yu. Data Envelopment Analysis in Multicriteria Decision Making. In Matthias Ehrgott and Xavier Gandibleux, editors, Multiple Criteria Optimization: State of the Art Annotated Bibliographic Surveys, pages 333–368. Kluwer Academic Publishers, Boston, 2002.

[701] Dongkyung Nam, Yun Deuk Seo, Lae-Jeong Park, Cheol Hoon Park, and Bumsup Kim. Parameter Optimization of a Voltage Reference Circuit using EP. In David B. Fogel, editor, Proceedings of the 1998 International Conference on Evolutionary Computation, pages 245–266, Piscataway, New Jersey, 1998. IEEE.

[702] S. Narayanan and S. Azarm. On Improving Multiobjective Genetic Algorithms for Design Optimization. In Proceedings of the Third World Congress of Structural and Multidisciplinary Optimization (WCSMO), Buffalo, New York, May 1999.

[703] S. Narayanan and S. Azarm. On Improving Multiobjective Genetic Algorithms for Design Optimization. Structural Optimization, 18:146–155, 1999.

[704] Sanjay Narayanan and Shapour Azarm. A Multiobjective Interactive Sequential Hybrid Optimization Technique for Design Decision Making. Engineering Optimization, 32:485–500, 2000.

[705] Boris Naujoks, Werner Haase, Jörg Ziegenhirt, and Thomas Bäck. Multi objective airfoil design using single parent populations. In W.B. Langdon, E. Cantú-Paz, K. Mathias, R. Roy, D. Davis, R. Poli, K. Balakrishnan, V. Honavar, G. Rudolph, J. Wegener, L. Bull, M.A. Potter, A.C. Schultz, J.F. Miller, E. Burke, and N. Jonoska, editors, Proceedings of the Genetic and Evolutionary Computation Conference (GECCO’2002), pages 1156–1163, San Francisco, California, July 2002. Morgan Kaufmann Publishers.

[706] Jorge S. Navarro, José A. Moreno, and Néstor Carrasquero. Evolutionary Multi-Objective Optimization of Simulation Models. In Alberto A. Ochoa Rodríguez, Marta R. Soto Ortiz, and Roberto Santana Hermida, editors, Second International Symposium on Artificial Intelligence (Adaptive Systems), ISAS’99, pages 242–250, La Havana, Cuba, 1999. Editorial Academia.

[707] Martijn Neef, Dirk Thierens, and Henryk Arciszewski. A Case Study of a Multiobjective Recombinative Genetic Algorithm with Coevolutionary Sharing. In 1999 Congress on Evolutionary Computation, pages 796–803, Washington, D.C., July 1999. IEEE Service Center.

[708] Walter Cede no and V. Rao Vemuri. On the Use of Niching for Dynamic Landscapes. In William Porto, editor, Proceedings of the 1997 IEEE International Conference on Evolutionary Computation, pages 361–366, Piscataway, New Jersey, April 1997. IEEE Press.

[709] Jason Noble and Richard A. Watson. Pareto coevolution: Using performance against coevolved opponents in a game as dimensions for Pareto selection. In Lee Spector, Erik D. Goodman, Annie Wu, W.B. Langdon, Hans-Michael Voigt, Mitsuo Gen, Sandip Sen, Marco Dorigo, Shahram Pezeshk, Max H. Garzon, and Edmund Burke, editors, Proceedings of the Genetic and Evolutionary Computation Conference (GECCO’2001), pages 493–500, San Francisco, California, 2001. Morgan Kaufmann Publishers.

[710] Stephen R. Norris and William A. Crossley. Pareto-Optimal Controller Gains Generated by a Genetic Algorithm. In AIAA 36th Aerospace Sciences Meeting and Exhibit, Reno, Nevada, January 1998. AIAA Paper 98-0010.

[711] S. Obayashi, S. Takahashi, and I. Fejtek. Transonic Wing Design by Inverse Optimization using MOGA. In Sixth Annual Conference of the Computational Fluid Dynamics Society of Canada, Quebec, Canada, June 1998.

[712] S. Obayashi, S. Takahashi, and Y. Takeguchi. Niching and Elitist Models for MOGAs. In A. E. Eiben, M. Schoenauer, and H.-P. Schwefel, editors, Parallel Problem Solving From Nature — PPSN V, pages 260–269, Amsterdam, Holland, 1998. Springer-Verlag.

[713] S. Obayashi, T. Tsukahara, and T. Nakamura. Cascade Airfoil Design by Multiobjective Genetic Algorithms. In Second International Conference on Genetic Algorithms in Engineering Systems: Innovations and Applications, pages 24–29. IEEE Conference Publication No. 446, September 1997.

[714] S. Obayashi, Y. Yamaguchi, and T. Nakamura. Multiobjective Genetic Algorithm for Multidisciplinary Design of Transonic Wing Planform. Journal of Aircraft, 34(5):690–693, September-October 1997.

[715] Shigeru Obayashi. Aerodynamic Inverse Optimization with Genetic Algorithms. In Proceedings of the IEEE International Conference on Industrial Technology, pages 421–425, Tongji University, 96TH8151, December 1996. IEEE Press.

[716] Shigeru Obayashi. Pareto Genetic Algorithm for Aerodynamic Design using the Navier-Stokes Equations. In D. Quagliarella, J. Périaux, C. Poloni, and G. Winter, editors, Genetic Algorithms and Evolution Strategies in Engineering and Computer Science. Recent Advances and Industrial Applications, chapter 12, pages 245–266. John Wiley and Sons, West Sussex, England, 1997.

[717] Shigeru Obayashi. Multidisciplinary Design Optimization of Aircraft Wing Planform Based on Evolutionary Algorithms. In Proceedings of the 1998 IEEE International Conference on Systems, Man, and Cybernetics, La Jolla, California, October 1998. IEEE.

[718] Shigeru Obayashi. Pareto Solutions of Multipoint Design of Supersonic Wings using Evolutionary Algorithms. In I.C. Parmee, editor, Adaptive Computing in Design and Manufacture V, pages 3–15, London, 2002. Springer-Verlag.

[719] Shigeru Obayashi, Kazuhiro Nakahashi, Akira Oyama, and Nobuhisa Yoshino. Design Optimization of Supersonic Wings Using Evolutionary Algorithms. In Proceedings of the Fourth ECCOMAS Computational Fluid Dynamics Conference, Athens, Greece, September 1998.

[720] Shigeru Obayashi, Takanori Tsukahara, and Takashi Nakamura. Multiobjective Evolutionary Computation for Supersonic Wing-Shape Optimization. IEEE Transactions on Evolutionary Computation, 4(2):182–187, July 2000.

[721] Shigeru Obayashi, Takanori Tsukahara, and Takashi Nakamura. Multiobjective Genetic Algorithm applied to Aerodynamic Design of Cascade Airfoils. IEEE Transactions on Industrial Electronics, 47(1), February 2000.

[722] V. Oduguwa and R. Roy. Multi-objective optimisation of rolling rod product design using meta-modelling approach. In W.B. Langdon, E. Cantú-Paz, K. Mathias, R. Roy, D. Davis, R. Poli, K. Balakrishnan, V. Honavar, G. Rudolph, J. Wegener, L. Bull, M.A. Potter, A.C. Schultz, J.F. Miller, E. Burke, and N. Jonoska, editors, Proceedings of the Genetic and Evolutionary Computation Conference (GECCO’2002), pages 1164–1171, San Francisco, California, July 2002. Morgan Kaufmann Publishers.

[723] Jaewon Oh, Hyokyung Bahn, Chisu Wu, , and Kern Koh. Pareto-based Soft Real-Time Task Scheduling in Multiprocessor Systems. In 7th IEEE Asia-Pacific Software Engineering Conference (APSEC 2000), pages 24–29, Singapore, 2000. IEEE.

[724] Tatsuya Okabe, Yaochu Jin, and Bernhard Sendhoff. On the Dynamics of Evolutionary Multi-Objective Optimization. In W.B. Langdon, E. Cantú-Paz, K. Mathias, R. Roy, D. Davis, R. Poli, K. Balakrishnan, V. Honavar, G. Rudolph, J. Wegener, L. Bull, M.A. Potter, A.C. Schultz, J.F. Miller, E. Burke, and N. Jonoska, editors, Proceedings of the Genetic and Evolutionary Computation Conference (GECCO’2002), pages 247–255, San Francisco, California, July 2002. Morgan Kaufmann Publishers.

[725] Matthias Ortmann and Wolfgang Weber. Multi-Criterion Optimization of Robot Trajectories with Evolutionary Strategies. In Proceedings of the 2001 Genetic and Evolutionary Computation Conference. Late-Breaking Papers, pages 310–316, San Francisco, California, July 2001.

[726] Matthias Ortmann and Wolfgang Weber. Multi-Criterion Optimization of Robot Trajectories with Evolutionary Strategies. In Lee Spector, Erik D. Goodman, Annie Wu, W.B. Langdon, Hans-Michael Voigt, Mitsuo Gen, Sandip Sen, Marco Dorigo, Shahram Pezeshk, Max H. Garzon, and Edmund Burke, editors, Proceedings of the Genetic and Evolutionary Computation Conference (GECCO’2001), page 1453, San Francisco, California, 2001. Morgan Kaufmann Publishers.

[727] Keith A. Osman, Anthony M. Higginson, and John Moore. Improving the efficiency of vehicle water-pump designs using genetic algorithms. In C. Dagli, M. Akay, A. Buczak, O. Ersoy, and B. Fernandez, editors, Smart Engineering Systems: Proceedings of the Artificial Neural Networks in Engineering Conference (ANNIE ’98), volume 8, pages 291–296, New York, 1998. ASME, ASME Press.

[728] Andrzej Osyczka. Evolutionary Algorithms for Single and Multicriteria Design Optimization. Physica Verlag, Germany, 2002. ISBN 3-7908-1418-0.

[729] Andrzej Osyczka and Stanislaw Krenich. A New Constraint Tournament Selection Method for Multicriteria Optimization using Genetic Algorithm. In 2000 Congress on Evolutionary Computation, volume 1, pages 501–507, Piscataway, New Jersey, July 2000. IEEE Service Center.

[730] Andrzej Osyczka and Stanislaw Krenich. Evolutionary Algorithms for Multicriteria Optimization with Selecting a Representative Subset of Pareto Optimal Solutions. In Eckart Zitzler, Kalyanmoy Deb, Lothar Thiele, Carlos A. Coello Coello, and David Corne, editors, First International Conference on Evolutionary Multi-Criterion Optimization, pages 141–153. Springer-Verlag. Lecture Notes in Computer Science No. 1993, 2001.

[731] Andrzej Osyczka, Stanislaw Krenich, and K. Karaś. Optimum Design of Robot Grippers using Genetic Algorithms. In Proceedings of the Third World Congress of Structural and Multidisciplinary Optimization (WCSMO), Buffalo, New York, May 1999.

[732] Andrzej Osyczka and Sourav Kundu. A Genetic Algorithm-Based Multicriteria Optimization Method. In Proceedings of First World Congress of Structural and Multidisciplinary Optimization, pages 909–914, Goslar, Germany, May 1995. Elsevier Science.

[733] Andrzej Osyczka and Sourav Kundu. A new method to solve generalized multicriteria optimization problems using the simple genetic algorithm. Structural Optimization, 10:94–99, 1995.

[734] Andrzej Osyczka and Sourav Kundu. Using genetic algorithms to solve multicriteria nonlinear programming problems. In Proceedings of Mendel’95, the 1st International Mendel Conference on Genetic Algorithms, Brno, Czech Republic, September 1995.

[735] Andrzej Osyczka and Sourav Kundu. A Genetic Algorithm Approach to Multicriteria Network Optimization Problems. In Proceedings of the 20th International Conference on Computers and Industrial Engineering, pages 329–332, Kyongju, Korea, October 1996.

[736] Andrzej Osyczka and Sourav Kundu. A modified distance method for multicriteria optimization using genetic algorithms. Computers in Industrial Engineering, 30(4):871–882, 1996.

[737] Andrzej Osyczka and H. Tamura. Pareto set distribution method for multicriteria optimization using genetic algorithm. In Proceedings of the Second International Conference on Genetic Algorithms (Mendel’96), pages 97–102, Brno, Czech Republic, June 1996.

[738] A. Oyama, S. Obayashi, K. Nakahashi, and N. Hirose. Coding by Taguchi Method for Evolutionary Algorithms Applied to Aerodynamic Optimization. In Proceedings of the Fourth ECCOMAS Computational Fluid Dynamics Conference, pages 196–203, Athens, Greece, September 1998. John Wiley & Sons.

[739] Akira Oyama and Meng-Sing Liou. Multiobjective Optimization of Rocket Engine Pumps using Evolutionary Algorithm. In Proceedings of the 15th AIAA Computational Fluid Dynamics Conference, Paper A01-31074, Anaheim, California, June 2001.

[740] Akira Oyama and Meng-Sing Liou. Multiobjective Optimization of Rocket Engine Pumps using Evolutionary Algorithm. Journal of Propulsion and Power, 18(3):528–535, May-June 2002.

[741] Akira Oyama and Shigeru Obayashi. Multidisciplinary Wing Design Optimization Using Multiobjective Evolutionary Algorithm. In PPSN/SAB Workshop on Multiobjective Problem Solving from Nature (MPSN), Paris, France, September 2000.

[742] Timucin Ozugur, Anand Bellary, and Falguni Sarkar. Multiobjctive Hierarchical 2G/3G Mobility Management Optimization: Niched Pareto Genetic Algorithm. In Global Telecommunications Conference, volume 6, pages 3681– 3685. IEEE, 2001.

[743] Ben Paechter, R.C. Rankin, Andrew Cumming, and Terence C. Fogarty. Timetabling the Classes of an Entire University with an Evolutionary Algorithm. In A. E. Eiben, Thomas Bäck, Marc Schoenauer, and Hans-Paul Schwefel, editors, Parallel Problem Solving From Nature — PPSN V, Amsterdam, Holland, 1998. Springer-Verlag. Lecture Notes in Computer Science No. 1498.

[744] N. Palli, S. Azarm, P. McCluskey, and R. Sundararajan. An Interactive Multistage epsilon-Inequality Constraint Method for Multiple Objectives Decision Making. Journal of Mechanical Design, Transactions of the ASME, 120:678– 686, 1998.

[745] Kwang-Wook Park and Donald E. Grierson. Pareto-Optimal Conceptual Design of the Structural Layout of Buildings Using a Multicriteria Genetic Algorithm. Computer-Aided Civil and Infrastructure Engineering, 14(3):163–170, May 1999.

[746] Seong-Jin Park. A Data Allocation Methodology Using the Multiple Aspects Petri Net and the Pareto Genetic Algorithm in Distributed Databases. PhD thesis, Korea University, 1997.

[747] Geoffrey T. Parks. Multiobjective PWR Reload Core Optimization Using Genetic Algorithms. In Proceedings of the International Conference on Mathematics and Computations, Reactor Physics, and Environmental Analyses, pages 615–624, 1995.

[748] Geoffrey T. Parks. Multiobjective Pressurized Water Reactor Reload Core Design by Nondominated Genetic Algorithm Search. Nuclear Science and Engineering, 124(1):178–187, 1996.

[749] Geoffrey T. Parks. Multiobjective Pressurised Water Reactor Reload Core Design using a Genetic Algorithm. In George D. Smith, Nigel C. Steele, and Rudolf F. Albrecht, editors, Artificial Neural Nets and Genetic Algorithms, pages 53–57, Norwich, UK, 1997. Springer-Verlag.

[750] Geoffrey T. Parks and I. Miller. Selective Breeding in a Multiobjective Genetic Algorithm. In A. E. Eiben, M. Schoenauer, and H.-P. Schwefel, editors, Parallel Problem Solving From Nature — PPSN V, pages 250–259, Amsterdam, Holland, 1998. Springer-Verlag.

[751] G.T. Parks, J. Li, M.-E. Balazs, and I. Miller. An empirical investigation of elitism in multiobjective genetic algorithms. Foundations of Computing and Decision Sciences, 26(1):51–74, 2001.

[752] G.T. Parks and A. Suppapitnarm. Multiobjective optimization of PWR reload core designs using simulated annealing. In Mathematics & Computation, Reactor Physics and Environmental Analysis in Nuclear Applications, volume 2, pages 1435–1444, Madrid, Spain, 1999.

[753] I. C. Parmee and G. Purchase. The development of a directed genetic search technique for heavily constrained design spaces. In I. C. Parmee, editor, Adaptive Computing in Engineering Design and Control-’94, pages 97–102, Plymouth, UK, 1994. University of Plymouth, University of Plymouth.

[754] Ian C. Parmee, Dragan Cvetković, Andrew H. Watson, and Christopher R. Bonham. Multi–objective Satisfaction within an Interactive Evolutionary Design Environment. Evolutionary Computation, 8(2):197–222, 2000.

[755] Ian C. Parmee, Dragan Cvetković, Andrew H. Watson, and Christopher R. Bonham. Multiobjective Satisfaction within an Interactive Evolutionary Design Environment. Evolutionary Computation, 8(2):197–222, Summer 2000.

[756] Ian C. Parmee and Andrew H. Watson. Preliminary Airframe Design Using Co-Evolutionary Multiobjective Genetic Algorithms. In W. Banzhaf, J. Daida, A. E. Eiben, M. H. Garzon, V. Honavar, M. Jakiela, and R. E. Smith, editors, Proceedings of the Genetic and Evolutionary Computation Conference (GECCO’99), volume 2, pages 1657–1665, San Francisco, California, July 1999. Morgan Kaufmann.

[757] I.C. Parmee. Poor-Definition, Uncertainty, and Human Factors—Satisfying Multiple Objectives in Real-World Decision-Making Environments. In Eckart Zitzler, Kalyanmoy Deb, Lothar Thiele, Carlos A. Coello Coello, and David Corne, editors, First International Conference on Evolutionary Multi-Criterion Optimization, pages 67–81. Springer-Verlag. Lecture Notes in Computer Science No. 1993, 2001.

[758] R. Parsons and S.L. Canfield. Developing genetic programming techniques for the design of compliant mechanisms. Structural and Multidisciplinary Optimization, 24(1):78–86, August 2002.

[759] K.E. Parsopoulos and M.N. Vrahatis. Particle Swarm Optimization Method in Multiobjective Problems. In Proceedings of the 2002 ACM Symposium on Applied Computing (SAC’2002), pages 603–607, Madrid, Spain, 2002. ACM Press.

[760] R. Perez, K. Behdinan, and J. Chung. Airfoil Shape Optimization Using Genetic Algorithms. In Proceedings, 47th Annual Conference of the Canadian Aeronautic and Space Institute, Aircraft Design and Development Symposium, Ottawa, Canada, April 30 - May 3 2000.

[761] Ruben Perez and Kamran Behdinan. Effective Multi-Mission Aircraft Conceptual Design Optimization Using a Hybrid Multi-Objective Evolutionary Method. In Proceedings of the 9th AIAA/ISSMO Symposium on Multidisciplinary Analysis and Optimization, AIAA-2002-5464. American Institute of Aeronautics and Astronautics, 2002.

[762] Jacques Périaux, Mourad Sefrioui, and Bertrand Mantel. RCS multi-objective optimization of scattered waves by active control elements using GAs. In Proceedings of the Fourth International Conference on Control, Automation, Robotics and Vision (ICARCV’96), Singapore, 1996.

[763] Jacques Périaux, Mourad Sefrioui, and Bertrand Mantel. GA Multiple Objective Optimization Strategies for Electromagnetic Backscattering. In D. Quagliarella, J. Périaux, C. Poloni, and G. Winter, editors, Genetic Algorithms and Evolution Strategies in Engineering and Computer Science. Recent Advances and Industrial Applications, chapter 11, pages 225–243. John Wiley and Sons, West Sussex, England, 1997.

[764] Andrei Petrovski and John McCall. Multi-objective Optimisation of Cancer Chemotherapy Using Evolutionary Algorithms. In Eckart Zitzler, Kalyanmoy Deb, Lothar Thiele, Carlos A. Coello Coello, and David Corne, editors, First International Conference on Evolutionary Multi-Criterion Optimization, pages 531–545. Springer-Verlag. Lecture Notes in Computer Science No. 1993, 2001.

[765] Henri Pierreval and Marie-France Plaquin. An Evolutionary Approach of Multicriteria Manufacturing Cell Formation. International Transactions in Operational Research, 5(1):13–25, January 1998.

[766] Carlo Poloni. Hybrid GA for Multi-Objective Aerodynamic Shape Optimization. In G. Winter, J. Periaux, M. Galan, and P. Cuesta, editors, Genetic Algorithms in Engineering and Computer Science, pages 397–416. Wiley & Sons, Chichester, 1995.

[767] Carlo Poloni, M. Fearon, and D. Ng. Parallelisation of Genetic Algorithms for Aerodynamic Design Optimisation. In Ian C. Parmee and M. J. Denham, editors, Proceedings of the Second International Conference on Adaptive Computing in Engineering Design and Control, pages 59–64, Plymouth, UK, 1996. University of Plymouth.

[768] Carlo Poloni, Giovanni Mosetti, and Stefano Contessi. Multiobjective Optimization by GAs: Application to System and Component Design. In Computational Methods in Applied Sciences ’96: Invited Lectures and Special Technological Sessions of the Third ECCOMAS Computational Fluid Dynamics Conference and the Second ECCOMAS Conference on Numerical Methods in Engineering, pages 258–264, Chichester, 1996. Wiley.

[769] Carlo Poloni and Valentino Pediroda. GA coupled with computationally expensive simulations: tools to improve efficiency. In D. Quagliarella, J. Périaux, C. Poloni, and G. Winter, editors, Genetic Algorithms and Evolution Strategies in Engineering and Computer Science. Recent Advances and Industrial Applications, chapter 13, pages 267–288. John Wiley and Sons, West Sussex, England, 1997.

[770] Ângela Cristina Martinho Guimarães Pereira. Extending Environmental Impact Assessment Processes: Generation of Alternatives for Siting and Routing Infraestructural Facilities by Multi-Criteria Evaluation and Genetic Algorithms. PhD thesis, New University of Lisbon, Lisbon, Portugal, 1997.

[771] Ângela Guimarães Pereira. Generating Alternative Routes using Genetic Algorithms and Multi-Criteria Analysis Techniques. In Ray Wyatt and Hemayet Hossain, editors, Fourth International Conference on Computers in Urban Planning and Urban Management), pages 547–560, Melbourne, Australia, July 11–14 1995.

[772] Ângela Guimarães Pereira. Generating alternative routes by multicriteria evaluation and a genetic algorithms. Environment and Planning B: Planning and Design, 23:711–720, 1996.

[773] Ângela Guimarães Pereira, G. Munda, and M. Pariccini. Generating alternatives for siting retail and service facilities using genetic algorithms and multiple criteria devision techniques. Journal of Retailing and Consumer Services, 1(2):40–47, 1994.

[774] Ângela Guimarães Pereira, Robert J. Peckham, and M. Paula Antunus. GENET: A Method to Generate Alternatives for Facilities Siting using Genetic Algorithms. In Janjaap Harts, Henk F. L. Ottens, and Henk J. Scholten, editors, Fourth European Conference and Exhibition on Geographical Information Systems (EGIS’93), pages 973–981, Genoa, Italy, March 29–April 1 1993.

[775] W. Pullan. Optimising Multiple Aspects of Network Survivability. In Congress on Evolutionary Computation (CEC’2002), volume 1, pages 115–120, Piscataway, New Jersey, May 2002. IEEE Service Center.

[776] R.C. Purshouse and P.J. Fleming. The Multi-Objective Genetic Algorithm Applied to Benchmark Problems—An Analysis. Technical Report 796, Department of Automatic Control and Systems Engineering, University of Sheffield, Sheffield, UK, August 2001.

[777] R.C. Purshouse and P.J. Fleming. Elitism, Sharing, and Ranking Choices in Evolutionary Multi-Criterion Optimisation. Technical Report 815, Department of Automatic Control and Systems Engineering, University of Sheffield, Sheffield, UK, January 2002.

[778] Robin C. Purshouse and Peter J. Fleming. Why use Elitism and Sharing in a Multi-Objective Genetic Algorithm? In W.B. Langdon, E. Cantú-Paz, K. Mathias, R. Roy, D. Davis, R. Poli, K. Balakrishnan, V. Honavar, G. Rudolph, J. Wegener, L. Bull, M.A. Potter, A.C. Schultz, J.F. Miller, E. Burke, and N. Jonoska, editors, Proceedings of the Genetic and Evolutionary Computation Conference (GECCO’2002), pages 520–527, San Francisco, California, July 2002. Morgan Kaufmann Publishers.

[779] Min Qiu. Prioritizing and Scheduling Road Projects by Genetic Algorithm. Mathematics and Computers in Simulation, 43:569–574, 1997.

[780] Domenico Quagliarella and Alessandro Vicini. Coupling Genetic Algorithms and Gradient Based Optimization Techniques. In D. Quagliarella, J. Périaux, C. Poloni, and G. Winter, editors, Genetic Algorithms and Evolution Strategies in Engineering and Computer Science. Recent Advances and Industrial Applications, chapter 14, pages 289–309. John Wiley and Sons, West Sussex, England, 1997.

[781] Domenico Quagliarella and Alessandro Vicini. Sub-population Policies for a Parallel Multiobjective Genetic Algorithm with Applications to Wing Design. In 1998 IEEE International Conference On Systems, Man, And Cybernetics, volume 4, pages 3142–3147, San Diego, California, October 1998. Institute of Electrical and Electronic Engineers (IEEE).

[782] Domenico Quagliarella and Alessandro Vicini. Designing High-Lift Airfoils Using Genetic Algorithms. In Kaisa Miettinen, Marko M. Mäkelä, Pekka Neittaanmäki, and Jacques Periaux, editors, Proceedings of EUROGEN’99, Jyväskyl, Finland, 1999. University of Jyváskylä.

[783] I. J. Ramírez-Rosado, J. L. Bernal-Agustín, V. Miranda, and L. M. Barbosa-Proenca. Multiobjective Planning of Power Distribution Systems Using Evolutionary Algorithms. In 8th IASTED International Conference (Modelling, Identification and Control-MIC’99), Innsbruck, Austria, February 1999.

[784] Ignacio J. Ramírez Rosado and José L. Bernal Agustín. Reliability and Cost Optimization for Distribution Networks Expansion Using an Evolutionary Algorithm. IEEE Transactions on Power Systems, 16(1):111–118, February 2001.

[785] S. Ranji Ranjithan, S. Kishan Chetan, and Harish K. Dakshima. Constraint Method-Based Evolutionary Algorithm (CMEA) for Multiobjective Optimization. In Eckart Zitzler, Kalyanmoy Deb, Lothar Thiele, Carlos A. Coello Coello, and David Corne, editors, First International Conference on Evolutionary Multi-Criterion Optimization, pages 299–313. Springer-Verlag. Lecture Notes in Computer Science No. 1993, 2001.

[786] K. N. Rao and S. Ganesan. ASIC benchmarking using Niched Pareto Genetic Algorithm. In Proceedings of International ICSC Symposia on Soft Computing (SOCO 96), pages B32–38, United Kingdom, October 1996.

[787] S. S. Rao. Genetic Algorithmic Approach for Multiobjective Optimization of Structures. In ASME Annual Winter Meeting, Structures and Controls Optimization, volume AD-Vol. 38, pages 29–38, New Orleans, Louisiana, November 1993. ASME.

[788] G. Ravi, Santosh K. Gupta, and M.B. Ray. Multiobjective optimization of cyclone separators using genetic algorithm. Industrial and Engineering Chemistry Research, 39(11):4272–4286, November 2000.

[789] G. Ravi, Santosh K. Gupta, S. Viswanathan, and M.B. Ray. Optimization of Venturi Scrubbers Using Genetic Algorithm. Industrial and Engineering Chemistry Research, 41(12):2988–3002, June 2002.

[790] Tapabrata Ray. Constrained Robust Optimal Design using a Multiobjective Evolutionary Algorithm. In Congress on Evolutionary Computation (CEC’2002), volume 1, pages 419–424, Piscataway, New Jersey, May 2002. IEEE Service Center.

[791] Tapabrata Ray, Tai Kang, and Seow Kian Chye. An Evolutionary Algorithm for Constrained Optimization. In Darrell Whitley, David Goldberg, Erick Cantú-Paz, Lee Spector, Ian Parmee, and Hans-Georg Beyer, editors, Proceedings of the Genetic and Evolutionary Computation Conference (GECCO’2000), pages 771–777, San Francisco, California, 2000. Morgan Kaufmann.

[792] Tapabrata Ray, Tai Kang, and Seow Kian Chye. Multiobjective Design Optimization by an Evolutionary Algorithm. Engineering Optimization, 33(3):399– 424, 2001.

[793] Tapabrata Ray and K.M. Liew. A Swarm Metaphor for Multiobjective Design Optimization. Engineering Optimization, 34(2):141–153, March 2002.

[794] Tapabrata Ray and Kang Tai. An Evolutionary Algorithm with a Multilevel Pairing Strategy for Single and Multiobjective Optimization. Foundations of Computing and Decision Sciences, 26:75–98, 2001.

[795] Brian J. Reardon. Fuzzy Logic vs. Niched Pareto Multiobjective Genetic Algorithm Optimization: Part I. Shaffer’s F2 Problem. Technical Report LA-UR-97-3675, Los Alamos National Laboratory, Los Alamos, New Mexico, September 1997.

[796] Brian J. Reardon. Fuzzy Logic vs. Niched Pareto Multiobjective Genetic Algorithm Optimization: Part II. A Simplified Born-Mayer Problem. Technical Report LA-UR-97-3676, Los Alamos National Laboratory, Los Alamos, New Mexico, September 1997.

[797] Brian J. Reardon. Optimization of Densification Modeling Parameters of Beryllium Powder using a Fuzzy Logic Based Multiobjective Genetic Algorithm. Technical Report LA-UR-98-1036, Los Alamos National Laboratory, Los Alamos, New Mexico, March 1998.

[798] Brian J. Reardon. Optimization of Micromechanical Densification Modeling Parameters For Copper Powder using a Fuzzy Logic Based Multiobjective Genetic Algorithm. Technical Report LA-UR-98-0419, Los Alamos National Laboratory, Los Alamos, New Mexico, January 1998.

[799] Patrick M. Reed, Barbara S. Minsker, and David E. Goldberg. Designing a New Elitist Nondominated Sorted Genetic Algorithm for a Multiobjective Long Term Groundwater Monitoring Application. In Proceedings of the 2001 Genetic and Evolutionary Computation Conference. Late-Breaking Papers, pages 352–358, San Francisco, California, July 2001.

[800] Patrick M. Reed, Barbara S. Minsker, and David E. Goldberg. Designing a New Elitist Nondominated Sorted Genetic Algorithm for a Multiobjective Long Term Groundwater Monitoring Application. In Lee Spector, Erik D. Goodman, Annie Wu, W.B. Langdon, Hans-Michael Voigt, Mitsuo Gen, Sandip Sen, Marco Dorigo, Shahram Pezeshk, Max H. Garzon, and Edmund Burke, editors, Proceedings of the Genetic and Evolutionary Computation Conference (GECCO’2001), page 1454, San Francisco, California, 2001. Morgan Kaufmann Publishers.

[801] Patrick M. Reed, Barbara S. Minsker, and David E. Goldberg. A multiobjective approach to cost effective long-term groundwater monitoring using an elitist nondominated sorted genetic algorithm with historical data. Journal of Hydroinformatics, 3(2):71–89, April 2001.

[802] Patrick M. Reed, Barbara S. Minsker, and David E. Goldberg. The Practitioner’s Role in Competent Search and Optimization Using Genetic Algorithms. In Don Phelps and Gerald Sehlke, editors, Bridging the Gap: Meeting the World’s Water and Environmental Resources Challenges. Proceedings of the World Water and Environmental Resources Congress, Washington, DC, 2001. American Society of Civil Engineers. ISBN 0-7844-0569-7.

[803] Patrick M. Reed, Barbara S. Minsker, and David E. Goldberg. Why Optimize Long Term Groundwater Monitoring Design? A Multiobjective Case Study of Hill Air Force Base. In Don Phelps and Gerald Sehlke, editors, Bridging the Gap: Meeting the World’s Water and Environmental Resources Challenges. Proceedings of the World Water and Environmental Resources Congress, Washington, DC, 2001. American Society of Civil Engineers. ISBN 0-7844-0569-7.

[804] Patrick Michael Reed. Striking the Balance: Long-Term Groundwater Monitoring Design for Multiple Conflicting Objectives. PhD thesis, Graduate College of the University of Illinois at Urbana-Champaign, Urbana, Illinois, 2002.

[805] P.M. Reed and B.S. Minsker. Discovery & Negotiation using Multiobjective Genetic Algorithms: A Case Study in Groundwater Monitoring Design. In Proceedings of Hydroinformatics 2002, Cardiff, UK, 2002.

[806] B. Rekiek, P. de Lit, and A. Delchambre. Hybrid assembly line design and user’s preferences. International Journal of Production Research, 40(5):1095– 1111, March 2002.

[807] Brahim Rekiek. Multiple-Objectives Genetic Algorithm. In Annie S. Wu, editor, Proceedings of the 1999 Genetic and Evolutionary Computation Conference. Workshop Program, page 401, Orlando, Florida, July 1999.

[808] Brahim Rekiek. Assembly Line Design (multiple objective grouping genetic algorithm and the balancing of mixed-model hybrid assembly line). PhD thesis, Free University of Brussels, CAD/CAM Department, Brussels, Belgium, December 2000.

[809] Brahim Rekiek, Pierre de Lit, Fabrice Pellichero, Thomas L’Eglise, Patrick Fouda, Emanuel Falkenauer, and Alan Delchambre. A multiple objective grouping genetic algorithm for assembly line design. Journal of Intelligent Manufacturing, 12(5–6):467–485, 2001.

[810] Brahim Rekiek, Pierre De Lit, Fabrice Pellichero, Thomas L’Eglise, Emanuel Falkenauer, and Alain Delchambre. Dealing With User’s Preferences in Hybrid Assembly Lines Design. In Proceedings of the MCPL’2000 Conference, 2000.

[811] Brahim Rekiek, Fabrice Pellichero, Pierre De Lit, Emanuel Falkenauer, and Alain Delchambre. A Resource Planner for Hybrid Assembly Lines. In Proceedings of the 15th International Conference on CAD/CAM Robotics & Factories of the Future CAR & FOF’99, volume 1, pages MW6–18–MW6–23, August 1999.

[812] J. H. Reynolds and E. D. Ford. Multi-criteria assessment of ecological process models. Ecology, 80(5):538–553, may 1999.

[813] Heather Erin Rickard. Feature Selection for Self-Organizing Feature Map Neural Networks With Applications in Medical Image Segmentation. Master’s thesis, Department of Electrical Engineering, University of Louisville, December 2001.

[814] Brian J. Ritzel, J. Wayland Eheart, and S. Ranjithan. Using genetic algorithms to solve a multiple objective groundwater pollution containment problem. Water Resources Research, 30(5):1589–1603, may 1994.

[815] Katya Rodrígez-Vázquez and Peter J. Fleming. Functionality and Optimality in Circuit Design: A Genetic Programming Approach. In Proceedings of the Third International Symposium on Adaptive Systems—Evolutionary Computation and Probabilistic Graphical Models, pages 23–28, Havana, Cuba, March 19–23 2001. Institute of Cybernetics, Mathematics and Physics.

[816] Mayron Rodrigues de Almeida, Marco Aurélio Cavalcanti Pachecho, Sílvio Hamacher, and Marley B.R. Vellasco. The Energy Minimization Method: A Multiobjective Fitness Evaluation Technique and Its Application to the Production Scheduling in a Petroleum Refinery. In Proceedings of the Congress on Evolutionary Computation 2001 (CEC’2001), volume 1, pages 560–567, Piscataway, New Jersey, May 2001. IEEE Service Center.

[817] Katya Rodríguez-Vázquez. Multiobjective Evolutionary Algorithms in Non-Linear System Identification. PhD thesis, Department of Automatic Control and Systems Engineering, The University of Sheffield, Sheffield, UK, 1999.

[818] Katya Rodríguez-Vázquez. Identification of MIMO Non-Linear Systems Using Evolutionary Computation. In Late Breaking Papers at the 2000 Genetic and Evolutionary Computation Conference, pages 411–417, Las Vegas, Nevada, July 2000.

[819] Katya Rodríguez-Vázquez and Peter J. Fleming. A Genetic Programming/NARMAX Approach to Non-Linear System Identification. In Genetic Algorithms in Engineering Systems: Innovations and Applications (GALE-SIA’97), pages 409–414, 1997.

[820] Katya Rodríguez-Vázquez and Peter J. Fleming. Multiobjective Genetic Programming for a Gas Turbine Engine Model Identification. In UKACC International Conference on Control’98, volume 2, pages 1385–1390, 1998.

[821] Katya Rodríguez-Vázquez and Peter J. Fleming. Multiobjective Genetic Programming for Non-Linear System Identification. Electronics Letters, 34(9):930–931, 1998.

[822] Katya Rodríguez-Vázquez and Peter J. Fleming. Non-Linear System Identification: Use of Genetic Programming to Satisfy Multiple Objectives. In Ajtonyi and Czap, editors, INTCOM’98 Symposium on Intelligent Systems in Control and Measurement, pages 148–154, 1998.

[823] Katya Rodríguez-Vázquez and Peter J. Fleming. Genetic Programming for Dynamic Chaotic Systems Modelling. In 1999 Congress on Evolutionary Computation, pages 22–28, Washington, D.C., July 1999. IEEE Service Center.

[824] Katya Rodríguez-Vázquez, Carlos M. Fonseca, and Peter J. Fleming. Multiobjective Genetic Programming : A Nonlinear System Identification Application. In John R. Koza, editor, Late Breaking Papers at the Genetic Programming 1997 Conference, pages 207–212, Stanford University, California, July 1997. Stanford Bookstore.

[825] James L. Rogers. Optimum Actuator Placement with a Genetic Algorithm for Aircraft Control. In Cihan H. Dagli, Anna L. Buczak, Joydeep Ghosh, Mark J. Embrechts, and Okan Ersoy, editors, Smart Engineering System Design: Neural Networks, Fuzzy Logic, Evolutionary Programming, Data Mining, and Complex Systems (ANNIE’99), pages 355–360, New York, November 1999. ASME Press.

[826] James L. Rogers. A Parallel Approach to Optimum Actuator Selection With A Genetic Algorithm. In AIAA Paper No. 2000-4484, AIAA Guidance, Navigation, and Control Conference, Denver, Colorado, August 14–17 2000.

[827] Carlos Eduardo Mariano Romero. Aprendizaje por Refuerzo en Optimización Multiobjetivo. PhD thesis, Departamento de Ciencias Computacionales, Instituto Tecnológico y de Estudios Superiores de Monterrey, Cuernavaca, Morelos, México, Marzo 2001. (In Spanish).

[828] Carlos Eduardo Mariano Romero and Eduardo Morales Manzanares. MOAQ an Ant-Q Algorithm for Multiple Objective Optimization Problems. In W. Banzhaf, J. Daida, A. E. Eiben, M. H. Garzon, V. Honavar, M. Jakiela, and R. E. Smith, editors, Genetic and Evolutionary Computing Conference (GECCO 99), volume 1, pages 894–901, San Francisco, California, July 1999. Morgan Kaufmann.

[829] Carlos Eduardo Mariano Romero and Eduardo Morales Manzanares. A New Approach for the Solution of Multiple Objective Optimization Problems Based on Reinforcement Learning. In Osvaldo Cairo, L. Enrique Sucar, and Francisco J. Cantu, editors, MICAI’2000), pages 212–223, Acapulco, México, April 2000. Springer-Verlag.

[830] Edmund Ronald. When Selection Meets Seduction. In Larry J. Eshelman, editor, Proceedings of the Sixth International Conference on Genetic Algorithms, pages 167–173, San Mateo, California, July 1995. University of Pittsburgh, Morgan Kaufmann Publishers.

[831] R. S. Rosenberg. Simulation of genetic populations with biochemical properties. PhD thesis, University of Michigan, Ann Harbor, Michigan, 1967.

[832] Jon Rowe, Kevin Vinsen, and Nick Marvin. Parallel GAs for Multiobjective Functions. In Jarmo T. Alander, editor, Proceedings of the Second Nordic Workshop on Genetic Algorithms and Their Applications (2NWGA), pages 61– 70, Vaasa, Finland, August 1996. University of Vaasa.

[833] Rajkumar Roy, Ahsutosh Tiwari, Olivier Munaux, and Graham Jared. Real-life engineering design optimization: Features and techniques. In J. Martikainen and J. Tanskanen, editors, CDROM Proceedings of the 5th Online World Conference on Soft Computing in Industrial Applications (WSC5)—ISBN 951-22-5205-8, Finland, 2000. IEEE.

[834] Bonnie Rubenstein-Montano and Ross A. Malaga. A Weighted Sum Genetic Algorithm to Support Multiple-Party Multi-Objective Negotiations. IEEE Transactions on Evolutionary Computation, 6(4):366–377, August 2002.

[835] Günter Rudolph. Evolutionary Search for Minimal Elements in Partially Ordered Finite Sets. In V.W. Porto, N. Saravanan, D. Waagen, and A.E. Eiben, editors, Evolutionary Programming VII, Proceedings of the 7th Annual Conference on Evolutionary Programming, pages 345–353, Berlin, 1998. Springer.

[836] Günter Rudolph. On a Multi-Objective Evolutionary Algorithm and Its Convergence to the Pareto Set. In Proceedings of the 5th IEEE Conference on Evolutionary Computation, pages 511–516, Piscataway, New Jersey, 1998. IEEE Press.

[837] Günter Rudolph. Evolutionary Search under Partially Ordered Fitness Sets. In Proceedings of the International NAISO Congress on Information Science Innovations (ISI 2001), pages 818–822. ICSC Academic Press: Millet/Sliedrecht, 2001.

[838] Günter Rudolph. A Partial Order Approach to Noisy Fitness Functions. In Proceedings of the Congress on Evolutionary Computation 2001 (CEC’2001), volume 1, pages 318–325, Piscataway, New Jersey, May 2001. IEEE Service Center.

[839] Günter Rudolph. Some Theoretical Properties of Evolutionary Algorithms under Partially Ordered Fitness Values. In Cs. Fabian and I. Intorsureanu, editors, Proceedings of the Evolutionary Algorithms Workshop (EAW-2001), pages 9– 22, Bucharest, Romania, January 2001.

[840] Günter Rudolph and Alexandru Agapie. Convergence Properties of Some Multi-Objective Evolutionary Algorithms. In Proceedings of the 2000 Conference on Evolutionary Computation, volume 2, pages 1010–1016, Piscataway, New Jersey, July 2000. IEEE Press.

[841] Alex J. Ruiz-Torres, E. Emory Enscore, and Russell R. Barton. Simulated Annealing Heuristics for the Average Flow-Time and the Number of Tardy Jobs Bi-Criteria Identical Parallel Machine Problem. Computers and Industrial Engineering, 33(1–2):257–260, 1997.

[842] Enrique H. Ruspini and Igor S. Zwir. Automated Qualitative Description of Measurements. In Proceedings of the 16th IEEE Instrumentation and Measurement Technology Conference, Venice, Italy, 1999.

[843] Conor Ryan. Pygmies and Servants. In Jr. Kenneth E. Kinnear, editor, Advances in Genetic Programming, pages 243–263. The MIT Press, Cambridge, Massachussets, 1994.

[844] Conor Ryan. Racial Harmony and Function Optimization in Genetic Algorithms—The Races Genetic Algorithm. In John R. McDonnell, Robert G. Reynolds, and David B. Fogel, editors, Evolutionary Programming IV: Proceedings of the Fourth Annual Conference on Evolutionary Programming, pages 109–125, Cambridge, Massachusetts, 1995. MIT Press.

[845] Sadiq M. Sait, Mahmood R. Minhas, and Junaid A. Khan. Performance and Low Power Driven VLSI Standard Cell Placement using Tabu Search. In Congress on Evolutionary Computation (CEC’2002), volume 1, pages 372– 377, Piscataway, New Jersey, May 2002. IEEE Service Center.

[846] Sadiq M. Sait, Habib Youseff, and Hussain Ali. Fuzzy Simulated Evolution Algorithm for Multi-objective Optimization of VLSI Placement. In 1999 Congress on Evolutionary Computation, pages 91–97, Washington, D.C., July 1999. IEEE Service Center.

[847] Sadiq M. Sait, Habib Youssef, and Junaid A. Khan. Fuzzy Evolutionary Algorithm for VLSI Placement. In Lee Spector, Erik D. Goodman, Annie Wu, W.B. Langdon, Hans-Michael Voigt, Mitsuo Gen, Sandip Sen, Marco Dorigo, Shahram Pezeshk, Max H. Garzon, and Edmund Burke, editors, Proceedings of the Genetic and Evolutionary Computation Conference (GECCO’2001), pages 1056–1063, San Francisco, California, 2001. Morgan Kaufmann Publishers.

[848] Masatoshi Sakawa. Genetic Algorithms and Fuzzy Multiobjective Optimization. Kluwer Academic Publishers, Boston, 2002. ISBN 0-7923-7452-5.

[849] Masatoshi Sakawa, Masahiro Inuiguchi, Hideaki Sunada, and Kazuya Sawada. Fuzzy Multiobjective Combinatorial Optimization Through Revised Genetic Algorithms. Japanese Journal of Fuzzy Theory and Systems, 6(1):77–88, 1994.

[850] Masatoshi Sakawa, Kosuke Kato, and Toshihiro Shibano. An interactive fuzzy satisficing method for multiobjective multidimensional 0-1 knapsack problems through genetic algorithms. In Proceedings of the 1996 International Conference on Evolutionary Computation (ICEC’96), pages 243–246, 1996.

[851] Masatoshi Sakawa, Kosuke Kato, and Toshihiro Shibano. Fuzzy Programming For Multiobjective 0-1 Programming Problems Through Revised Genetic Algorithms. European Journal of Operational Research, 97(1):149–158, 1997.

[852] Masatoshi Sakawa and R. Kubota. Fuzzy programming for multiobjective job shop scheduling with fuzzy processing time and fuzzy duedate through genetic algorithms. European Journal of Operational Research, 120(2):393–407, 2000.

[853] Masatoshi Sakawa and Toshihiro Shibano. An interactive fuzzy satisficing method for multiobjective 0-1 programming problems with fuzzy numbers through genetic algorithms with double strings. European Journal of Operational Research, 107(3):564–574, June 1998.

[854] Masatoshi Sakawa, Toshihiro Shibano, and Kosuke Kato. An interactive fuzzy satisficing method for multiobjective integer programming problems through genetic algorithms. In L.C. Jain and R.K. Jain, editors, Second International Conference on Knowledge-Based Intelligent Electronic Systems, pages 94– 100, Adelaide, Australia, 1998. IEEE.

[855] Masatoshi Sakawa, Toshihiro Shibano, and Hidenobu Obata. An Interactive Fuzzy Method for Multiobjective 0-1 Programming Problems with Fuzzy Number Criteria Using Genetic Algorithms. Electronics and Communications in Japan (Part III: Fundamental Electronic Science), 81(8):64–72, 1998.

[856] R. Saker, H.A. Abbass, and S. Karim. An Evolutionary Algorithm for Constrained Multiobjective Optimization Problems. In The 5th Australasia-Japan Joint Workshop on Intelligent and Evolutionary Systems (AJWIS’2001), pages 113–122, Dunedin, New Zealand, November 2001.

[857] R. Saker, H.A. Abbass, and C. Newton. Solving Multiobjective Optimization Problems Using Evolutionary Algorithm. In The International Conference on Computational Intelligence for Modelling, Control and Automation (CIMCA’2001), pages 149–160, Las Vegas, Nevada, July 2001.

[858] F. Sibel Salman, Jayan Kalagnanam, and Sesh Murthy. Cooperative Strategies for Solving the Bicriteria Sparse Multiple Knapsack Problem. In 1999 Congress on Evolutionary Computation, pages 53–60, Washington, D.C., July 1999. IEEE Service Center.

[859] Fatma Sibel Salman, Jayan Kalagnanam, and Sesh Murthy. Heuristics for Solving the Bicriteria Sparse Multiple Knapsack Problem. Technical Report RC 21059, IBM T.J. Watson Research Center, 1997.

[860] Fatma Sibel Salman, Jayant Kalagnanam, and Sesh Murthy. Cooperative strategies for solving the bicriteria sparse multiple knapsack problem. In 1999 Congress on Evolutionary Computation, pages 53–60, Washington, D.C., July 1999. IEEE Service Center.

[861] L. Saludjian, J. L. Coulomb, and A. Izabelle. Genetic Algorithm and Taylor Development of the Finite Element Solution for Shape Optimization of Electromagnetic Devices. IEEE Transactions on Magnetics, 34(5):2841–2844, September 1998.

[862] Eric Sandgren. Multicriteria design optimization by goal programming. In Hojjat Adeli, editor, Advances in Design Optimization, chapter 23, pages 225– 265. Chapman & Hall, London, 1994.

[863] Amâncio Santos and António Dourado Pereira Correia. Constrained GA Applied to Production and Energy Mangement of a Pulp and Paper Mill. In Janice Carroll, Hisham Haddad, Dave Oppenheim, Barrett Bryant, and Gary B. Lamont, editors, Proceedings of the 1999 ACM Symposium on Applied Computing, pages 324–332, San Antonio, Texas, 1999. ACM.

[864] R. Sarker, K. Liang, and C. Newton. A Multiobjective Evolutionary Algorithm. In International Computer Science Convention Congress on Intelligent Systems and Applications (ISA’2000), volume 2, pages 125–131, Wollongong, Australia, 2000.

[865] R. Sarker, K. Liang, and C. Newton. A New Evolutionary Algorithm for Multiobjective Optimization. European Journal of Operational Research, 140(1):12–23, 2002.

[866] Ruhul Sarker, H. Abbass, and C. Newton. Solving Two Multi-objective Optimization Problems using Evolutionary Algorithm. In M. Mohammadian, R. Sarker, and X. Yao, editors, Computational Intelligence in Control. Idea Group Publishing, USA, 2002.

[867] Ruhul Sarker and Carlos A. Coello Coello. Assessment Methodologies for Multiobjective Evolutionary Algorithms. In Ruhul Sarker, Masoud Mohammadian, and Xin Yao, editors, Evolutionary Optimization, pages 177–195. Kluwer Academic Publishers, New York, February 2002. ISBN 0-7923-7654-4.

[868] Ruhul Sarker, Masoud Mohammadian, and Xin Yao. Evolutionary Optimization. Kluwer Academic Publishers, Boston, Massachusetts, February 2002. ISBN 0-7923-7654-4.

[869] Hiroshi Sato and Akira Namatame. Co-evolution in Social Interactions. In Proceedings of the Congress on Evolutionary Computation 2001 (CEC’2001), volume 2, pages 1109–1114, Piscataway, New Jersey, May 2001. IEEE Service Center.

[870] Dragan A. Savic, Godfrey A. Walters, and Martin Schwab. Multiobjective Genetic Algorithms for Pump Scheduling in Water Supply. In AISB International Workshop on Evolutionary Computing. Lecture Notes in Computer Science 1305, pages 227–236, Berlin, April 1997. Springer-Verlag.

[871] T. M. Sayers and J. M. Anderson. The multi-objective optimisation of a traffic control system. In Avi Ceder, editor, Proceedings of 14th International Symposium on Transportation and Traffic Theory, pages 153–176, Haifa, Israel, July 1999. Technion-Israel, Institute of Technology, Transportation Research Institute.

[872] Ivo F. Sbalzarini, Sibylle Müller, and Petros Koumoutsakos. Multiobjective Optimization using Evolutionary Algorithms. In Center for Turbulence Research. Proceedings of the 2000 Summer Program, pages 63–74. NASA Ames/Stanford University, 2000.

[873] Ivo F. Sbalzarini, Sibylle Müller, and Petros Koumoutsakos. Microchannel Optimization Using Multiobjective Evolution Strategies. In Eckart Zitzler, Kalyanmoy Deb, Lothar Thiele, Carlos A. Coello Coello, and David Corne, editors, First International Conference on Evolutionary Multi-Criterion Optimization, pages 516–530. Springer-Verlag. Lecture Notes in Computer Science No. 1993, 2001.

[874] J. David Schaffer. Multiple Objective Optimization with Vector Evaluated Genetic Algorithms. PhD thesis, Vanderbilt University, 1984.

[875] J. David Schaffer. Multiple Objective Optimization with Vector Evaluated Genetic Algorithms. In Genetic Algorithms and their Applications: Proceedings of the First International Conference on Genetic Algorithms, pages 93–100. Lawrence Erlbaum, 1985.

[876] J. David Schaffer and John J. Grefenstette. Multiobjective Learning via Genetic Algorithms. In Proceedings of the 9th International Joint Conference on Artificial Intelligence (IJCAI-85), pages 593–595, Los Angeles, California, 1985. AAAI.

[877] Michael Scheffler and Gerhard Tröster. A multi-objective test vs. cost optimization for electronic products. In Proceedings of the 26th IEEE/CPMT International Electronics Manufacturing Technology Symposium, pages 344– 351. IEEE, 2000.

[878] Frank Schmiedle, Nicole Drechsler, Draniel Große, and Rolf Drechsler. Priorities in Multi-Objective Optimization for Genetic Programming. In Lee Spector, Erik D. Goodman, Annie Wu, W.B. Langdon, Hans-Michael Voigt, Mitsuo Gen, Sandip Sen, Marco Dorigo, Shahram Pezeshk, Max H. Garzon, and Edmund Burke, editors, Proceedings of the Genetic and Evolutionary Computation Conference (GECCO’2001), pages 129–136, San Francisco, California, 2001. Morgan Kaufmann Publishers.

[879] Volker Schnecke and Oliver Vornberger. Hybrid Genetic Algorithms for Constrained Placement Problems. IEEE Transactions on Evolutionary Computation, 1(4):266–277, November 1997.

[880] Jason R. Schott. Fault Tolerant Design Using Single and Multicriteria Genetic Algorithm Optimization. Master’s thesis, Department of Aeronautics and Astronautics, Massachusetts Institute of Technology, Cambridge, Massachusetts, May 1995.

[881] P. Schroder, A. J. Chipperfield, P. J. Fleming, and N. Grum. Multi-Objective Optimization of Distributed Active Magnetic Bearing Controllers. In Genetic Algorithms in Engineering Systems: Innovations and Applications, pages 13– 18. IEE, September 1997.

[882] M. Schwab, D. A. Savic, and G. A. Walters. Multi-Objective Genetic Algorithm for Pump Scheduling in Water Supply Systems. Technical Report 96/02, Centre For Systems And Control Engineering, School of Engineering, University of Exeter, Exeter, United Kingdom, 1996.

[883] M. Sefrioui and J. Periaux. Nash Genetic Algorithms: examples and applications. In 2000 Congress on Evolutionary Computation, volume 1, pages 509–516, San Diego, California, July 2000. IEEE Service Center.

[884] Mourad Sefrioui and Jacques Périaux. A hierarchical Genetic Algorithm Using Multiple Models for Optimization. In Marc Schoenauer, Kalyanmoy Deb, Günter Rudolph, Xin Yao, Evelyne Lutton, Juan Julian Merelo, and Hans-Paul Schwefel, editors, Proceedings of the Parallel Problem Solving from Nature VI Conference, pages 879–888, Paris, France, 2000. Springer. Lecture Notes in Computer Science No. 1917.

[885] P. Sen and J. B. Yang. Multiple Criteria Decision Support in Engineering Design. Springer-Verlag, London, 1998.

[886] Paolo Serafini. Simulated Annealing for Multiple Objective Optimization Problems. In G.H. Tzeng, H.F. Wang, U.P. Wen, and P.L. Yu, editors, Proceedings of the Tenth International Conference on Multiple Criteria Decision Making: Expand and Enrich the Domains of Thinking and Application, volume 1, pages 283–294, Berlin, 1994. Springer-Verlag.

[887] Stefan Sette, Luc Boullart, and Lieva Van Langenhove. Optimizing a Production Process by a Neural Network/Genetic Algorithm Approach. Engineering Applications in Artificial Intelligence, 9(6):681–689, 1996.

[888] K. J. Shaw and P. J. Fleming. Initial Study of Practical Multi-Objective Genetic Algorithms for Scheduling the Production of Chilled Ready Meals. In Proceedings of Mendel’96, the 2nd International Mendel Conference on Genetic Algorithms, Brno, Czech Republic, September 1996.

[889] K. J. Shaw and P. J. Fleming. An Initial Study of Practical Multi-Objective Production Scheduling using Genetic Algorithms. In Proceedings of the International Conference on Control’96, University of Exeter, UK, September 1996.

[890] K. J. Shaw and P. J. Fleming. Including Real-Life Preferences in Genetic Algorithms to Improve Optimisation of Production Schedules. In Proceedings of the GALESIA’97, pages 239–244, Glasgow, Scotland, September 1997. IEE.

[891] K. J. Shaw and P. J. Fleming. An Overview of Multi-Objective Genetic Algorithms for Production Scheduling. In Proceedings of the DTI/ACTT Regional Seminar in Finite Capacity Scheduling in the Process Industries, Manchester, UK, January 1997.

[892] K. J. Shaw and P. J. Fleming. Use of Rules and Preferences for Schedule Builders in Genetic Algorithms Production Scheduling. In Corne and Shapiro, editors, Proceedings of the AISB’97 Workshop on Evolutionary Computation.

Lecture Notes in Computer Science No. 1305, Manchester University, 1997. Springer-Verlag.

[893] K. J. Shaw, C. M. Fonseca, and P. J. Fleming. A Simple Demonstration of a Quantitative Technique for Comparing Multiobjective Genetic Algorithm Performance. In Annie S. Wu, editor, Proceedings of the 1999 Genetic and Evolutionary Computation Conference. Workshop Program, pages 119–120, Orlando, Florida, July 1999.

[894] K. J. Shaw, P. L. Lee, H. P. Nott, and M. Thompson. Genetic Algorithms for Multiobjective Scheduling of Combined Batch/Continuous Process Plants. In 2000 Congress on Evolutionary Computation, volume 1, pages 293–300, Piscataway, New Jersey, July 2000. IEEE Service Center.

[895] K. J. Shaw, A. L. Nortcliffe, M. Thompson, J. Love, C. M. Fonseca, and P. J. Fleming. Assessing the Performance of Multiobjective Genetic Algorithms for Optimization of a Batch Process Scheduling Problem. In 1999 Congress on Evolutionary Computation, pages 37–45, Washington, D.C., July 1999. IEEE Service Center.

[896] Katharine Jane Shaw. Using Genetic Algorithms for Practical Multi-Objective Production Schedule Optimisation. PhD thesis, Department of Automatic Control and Systems Engineering, The University of Sheffield, Sheffield, UK, 1997.

[897] Porter Sherman. Ranking Techniques in Multicriteria Genetic Algorithm-Based Optimization. PhD thesis, Department of Computer and Information Science, Polytechnic University, Brooklyn, New York, 1995.

[898] Toshihiro Shibano and Masatoshi Sakawa. Interactive Decision Making for Fuzzy Multiobjective 0-1 Programs Through Genetic Algorithms with Double Strings. In Proceedings of the Sixth IEEE Conference on Fuzzy Systems, pages 1639–1644, 1997.

[899] Mun Bo Shim, Tomonari Furukawa, and Shinobu Yoshimura. Pareto-based continuous evolutionary algorithms for multi-objective optimization. Engineering Computations, 19(1):22–48, 2002.

[900] Yoshiaki Shimizu. Multi-Objective Optimization for Site Location Problems through Hybrid Genetic Algorithm with Neural Networks. Journal of Chemical Engineering of Japan, 32(1):51–58, 1999.

[901] Jaqueline S. Shoaf and James A. Foster. A Genetic Algorithm Solution to the Efficient Set Problem: A Technique for Portfolio Selection Based on the Markowitz Model. In Proceedings of the Decision Sciences Institute Annual Meeting, pages 571–573, Orlando, Florida, 1996.

[902] Eric V. Siegel and Alexander D. Chaffee. Genetically Optimizing the Speed of Programs evolved to Play Tetris. In Peter J. Angeline and Kenneth E. Kinnear, editors, Advances in Genetic Programming 2, pages 279–298. MIT Press, 1996.

[903] Dong-Joon Sim, Hyun-Kyo Jung, and Song-Yop Hahn. Multiobjective optimal design of interior permanent magnet synchronous motors considering improved core loss formula. In 1997 IEEE International Electric Machines and Drives Conference Record, pages MA1/2.1–MA1/2.3, Milwaukee, Wisconsin, May 1997.

[904] Dong-Joon Sim, Hyun-Kyo Jung, Song-Yop Hahn, and Jong-Soo Won. Application of vector optimization employing modified genetic algorithm to permanent magnet motor design. In Proceedings of Seventh Conference on Electromagnetic Field Computation - CEFEC, page 288, Okayama, Japan, March 1996.

[905] Krzysztof Socha and Marek Kisiel-Dorohinicki. Agent-based Evolutionary Multiobjective Optimisation. In Congress on Evolutionary Computation (CEC’2002), volume 1, pages 109–114, Piscataway, New Jersey, May 2002. IEEE Service Center.

[906] Grant A.E. Soremekun. Genetic Algorithms for Composite Laminate Design and Optimization. Master’s thesis, Department of Mechanical Engineering, Virginia Polytechnic Institute, Blacksburgh, Virginia, February 5 1997.

[907] S. Sorooshian, L. A. Bastidas, and H. V. Gupta. Application of Multi-Objective Optimization Algorithms for Hydrologic Model Identification and Parameterization. In Proceedings of Second International Conference on Multiple Objective Decision Support Systems for Land, Water, and Environmental Management, Brisbane, Australia, August 1999.

[908] Jagabandhu Sridhar and Chandrasekharan Rajendran. Scheduling in Flowshop and Cellular Manufacturing Systems with Multiple Objectives – A Genetic Algorithmic Approach. Production Planning & Control, 7(4):374–382, 1996.

[909] Kishan Chetan Srigiriraju. Noninferior Surface Tracing Evolutionary Algorithm (NSTEA) for Multi Objective Optimization. Master’s thesis, North Carolina State University, Raleigh, North Carolina, August 2000.

[910] N. Srinivas and Kalyanmoy Deb. Multiobjective optimization using nondominated sorting in genetic algorithms. Technical report, Department of Mechanical Engineering, Indian Institute of Technology, Kanpur, India, 1993.

[911] N. Srinivas and Kalyanmoy Deb. Multiobjective Optimization Using Nondominated Sorting in Genetic Algorithms. Evolutionary Computation, 2(3):221– 248, Fall 1994.

[912] N. Srinivas and Kalyanmoy Deb. Comparative study of vector evaluated GA and NSGA applied to multiobjective optimization. In P. K. Roy and S. D. Mehta, editors, Proceedings of the Symposium on Genetic Algorithms, pages 83–90, 1995.

[913] Paul M. Stanfield, Russell E. King, and Thom J. Hodgson. Multi-objective stochastic scheduling of job ready times. Annals of Operations Research, 70:221–239, 1997.

[914] Timothy J. Stanley and Trevor Mudge. A Parallel Genetic Algorithm for Multiobjective Microprocessor Design. In Larry J. Eshelman, editor, Proceedings of the Sixth International Conference on Genetic Algorithms, pages 597–604, San Mateo, California, July 1995. University of Pittsburgh, Morgan Kaufmann Publishers.

[915] G.E. Stravroulakis and H. Antes. Flaw identification in elastomechanics: BEM simulation with local and genetic optimization. Structural Optimization, 16(2/3):162–175, 1998.

[916] R. Sundararajan, S. Azarm, P. McCluskey, and N. Palli. A Stress Model for Multiobjective Design Optimization of a Power Electronic Module. Mechanics of Structures and Machines, 27(2):163–183, 1999.

[917] A. Suppapitnarm, G.T. Parks, K. Shea, and P.J. Clarkson. A Multiobjective Optimisation Approach for the Conceptual Design of Frame Structures. In I.C. Parmee, editor, Adaptive Computing in Design and Manufacture V, pages 109–120, London, 2002. Springer-Verlag.

[918] A. Suppapitnarm, K.A. Seffen, G.T. Parks, and P.J. Clarkson. A simulated annealing algorithm for multiobjective optimization. Engineering Optimization, 33(1):59–85, 2000.

[919] A. Suppapitnarm, K.A. Seffen, G.T. Parks, and A.M. Connor. Multiobjective optimisation of bicycle frames using simulated annealing. In Proceedings of the First ASMO/ISSMO Conference on Engineering Design Optimization, volume 1, pages 357–364, Ilkley, West Yorkshire, 1999.

[920] A. Suppapitnarm, K.A. Seffen, G.T. Parks, and J.-S. Liu. Design by multiobjective optimisation using simulated annealing. In Proceedings of the 12th International Conference in Engineering Design (ICED’99), volume 3, pages 1395–1400, Munich, Germany, 1999.

[921] Patrick D. Surry and Nicholas J. Radcliffe. The COMOGA Method: Constrained Optimisation by Multiobjective Genetic Algorithms. Control and Cybernetics, 26(3):391–412, 1997.

[922] Patrick D. Surry, Nicholas J. Radcliffe, and Ian D. Boyd. A Multi-Objective Approach to Constrained Optimisation of Gas Supply Networks : The CO-MOGA Method. In Terence C. Fogarty, editor, Evolutionary Computing. AISB Workshop. Selected Papers, Lecture Notes in Computer Science, pages 166– 180, Sheffield, U.K., 1995. Springer-Verlag.

[923] Gilbert Syswerda and Jeff Palmucci. The Application of Genetic Algorithms to Resource Scheduling. In Richard K. Belew and Lashon B. Booker, editors, Proceedings of the Fourth International Conference on Genetic Algorithms, pages 502–508, San Mateo, California, 1991. Morgan Kaufmann.

[924] Ricardo Szmit and Amnon Barak. Evolution Strategies for a Parallel Multi-Objective Genetic Algorithm. In Darrell Whitley, David Goldberg, Erick Cantú-Paz, Lee Spector, Ian Parmee, and Hans-Georg Beyer, editors, Proceedings of the Genetic and Evolutionary Computation Conference (GECCO’2000), pages 227–234, San Francisco, California, 2000. Morgan Kaufmann.

[925] Kanta Tachibana and Takeshi Furuhashi. A Structure Identification Method of Submodels for Hierarchical Fuzzy Modelling Using the Multiple Objective Genetic Algorithm. International Journal of Intelligent Systems, 17(5):495– 513, May 2002.

[926] T. Tagami and T. Kawabe. Genetic Algorithm with a Pareto Partitioning Method for Multi-objective Flowshop Scheduling. In Proceedings of the 1998 International Symposium of Nonlinear Theory and its Applications (NOLTA’98), pages 1069–1072, Crans-Montana, 1998.

[927] T. Tagami and T. Kawabe. Genetic Algorithms using Pareto Partitioning Method for Multiobjective Optimization Problems. Transactions of The Institute of Systems, Control and Information Engineers, 11(11):600–607, 1998. (In Japanese).

[928] T. Tagami and T. Kawabe. Genetic Algorithm based on a Pareto Neighborhood Search for Multiobjective Optimization. In Proceedings of the 1999 International Symposium of Nonlinear Theory and its Applications (NOLTA’99), pages 331–334, Hawaii, 1999.

[929] Kiyoharu Tagawa, Noboru Wakabayashi, Hiromasa Haneda, and Katsumi Inoue. An Imanishism-Based Genetic Algorithm for Sampling Various Pareto-Optimal Solutions: An Application to the Multiobjective Resource Division Problem. Electrical Engineering in Japan, Volume 139(2):23–35, April 2002.

[930] Kiyoharu Tagawa, Noboru Wakabayashi, Kenji Kanesige, and Hiromasa Haneda. A New Genetic Algorithm based on Anti-Darwinism for Multi-Objective Part-Tool Grouping Problem. In Proceedings of the 2000 IEEE International Symposium on Industrial Electronics (ISIE’2000), volume 2, pages 782–787. IEEE, 2000.

[931] Y. Takada, M. Yamamura, and S. Kobayashi. An Approach to Portfolio Selection Problems Using Multi-Objective Genetic Algorithms. In Proceedings of the 23rd Symposium on Intelligent Systems, pages 103–108, 1996.

[932] Ricardo H.C. Takahashi, Reinaldo M. Palhares, Daniel A. Cutra, and Leila P.S. Gonçalves. Synthesis and characterization of pareto-optimal solutions for the mixed H_{2}/H_{∞} control problem. In Proceedings of the 40th IEEE International Conference on Decision and Control, volume 4, pages 3997–4002. IEEE, 2001.

[933] S. Takahashi, S. Obayashi, and K. Nakahashi. Transonic Shock-Free Wing Design with Multiobjective Genetic Algorithms. In Proceedings of the International Conference on Fluid Engineering, volume 1, pages 425–429, Tokyo, Japan, July 1997.

[934] S. Takahashi, S. Obayashi, and K. Nakahashi. Inverse Optimization of Transonic Wing Shape for Mid-Size Regional Aircraft. AIAA Paper 98-0601, January 1998.

[935] El-Ghazali Talbi, Malek Rahoual, Mohamed Hakim Mabed, and Clarisse Dhaenens. A Hybrid Evolutionary Approach for Multicriteria Optimization Problems: Application to the Flow Shop. In Eckart Zitzler, Kalyanmoy Deb, Lothar Thiele, Carlos A. Coello Coello, and David Corne, editors, First International Conference on Evolutionary Multi-Criterion Optimization, pages 416–428. Springer-Verlag. Lecture Notes in Computer Science No. 1993, 2001.

[936] Hisashi Tamaki, Hajime Kita, and Shigenobu Kobayashi. Multi-Objective Optimization by Genetic Algorithms : A Review. In Toshio Fukuda and Takeshi Furuhashi, editors, Proceedings of the 1996 International Conference on Evolutionary Computation (ICEC’96), pages 517–522, Nagoya, Japan, 1996. IEEE.

[937] Hisashi Tamaki, M. Mori, and M. Araki. Generation of a Set of Pareto-Optimal Solutions by Genetic Algorithms. Transactions of the Society of Instrument and Control Engineers, 31(8):1185–1192, 1995.

[938] Hisashi Tamaki, M. Mori, M. Araki, Y. Mishima, and H. Ogai. Multi-Criteria Optimization by Genetic Algorithms : A Case of Scheduling in Hot Rolling Process. In Proceedings of the 3rd Conference of the Association of Asian-Pacific Operational Research Societies within IFORS (APORS’94), pages 374– 381. World Scientific, 1995.

[939] Hisashi Tamaki, T. Mukai, K. Kawakami, and M. Araki. Genetic Algorithm Approach to Multi-Objective Scheduling Problems with Regular and Non-Regular Objective Functions. In Toshio Fukuda and Takeshi Furuhashi, editors, Proceedings of the International Conferences on Advances in Production Management Systems (APMS’96), pages 553–556, 1996.

[940] Hisashi Tamaki and Etsuo Nishino. A Genetic Algorithm Approach to Multi-Objective Scheduling Problems with Regular and Non-Regular Objective Functions. In Proceedings of the 8th IFAC/IFORS/IMACS/IFIP Symposium on Large Scale Systems : Theory and Applications (LSS’98), pages 289–294, 1998.

[941] Hisashi Tamaki, Etsuo Nishino, and Shigeo Abe. A Genetic Algorithm Approach to Multi-Objective Scheduling Problems with Earliness and Tardiness Penalties. In 1999 Congress on Evolutionary Computation, pages 46–52, Washington, D.C., July 1999. IEEE Service Center.

[942] K. C. Tan, T. H. Lee, D. Khoo, and E. F. Khor. MOEA Toolbox for Computer-Aided Multi-Objective Optimization. In 2000 Congress on Evolutionary Computation, volume 1, pages 38–45, Piscataway, New Jersey, July 2000. IEEE Service Center.

[943] K. C. Tan, T. H. Lee, and E. F. Khor. Control system design automation with robust tracking thumbprint performance using a multi-objective evolutionary algorithm. In IEEE Int. Conf. Control Appl. and Sys. Design, pages 498–503, Hawaii, 1999.

[944] K. C. Tan, T. H. Lee, and E. F. Khor. Evolutionary Algorithms with Goal and Priority Information for Multi-Objective Optimization. In 1999 Congress on Evolutionary Computation, pages 106–113, Washington, D.C., July 1999. IEEE Service Center.

[945] K. C. Tan, T. H. Lee, E. F. Khor, and K. Ou. Control system design unification and automation using an incremented multi-objective evolutionary algorithm. In Proceedings of the 19th IASTED International Conference on Modeling, Identification and Control, Innsbruck, Austria, 2000.

[946] K. C. Tan, T. H. Lee, E. F. Khor, and R. Sathikannan. Incremented multiobjective evolutionary design automation of robust tracking thumbprint performances in QFT. In Proceedings of the International Conference on Evolutionary Computation for Computer, Communication, Control and Power, pages 137–142, Chennai, India, 2000.

[947] Kay Chen Tan and Yun Li. Multi-Objective Genetic Algorithm Based Time and Frequency Domain Design Unification of Linear Control Systems. Technical Report CSC-97007, Department of Electronics and Electrical Engineering, University of Glasglow, Glasglow, Scotland, 1997.

[948] Kay Chen Tan and Yun Li. Multi-Objective Genetic Algorithm Based Time and Frequency Domain Design Unification of Linear Control Systems. In Proceedings of the IFAC/IEEE International Symposium on Artificial Intelligence in Real-Time Control, pages 61–66, Kuala Lumpur, Malaysia, September 1997.

[949] Kay Chen Tan and Yun Li. Multi-objective genetic algorithm based time and frequency domain design unification of linear control systems. In IFAC International Symposium on Artificial Intelligence and Real-Time Control, pages 61–66, Kuala Lumpa, Malaysia, September 1997.

[950] K.C. Tan, E.F. Khor, C. M. Heng, and T.H. Lee. Exploratory Multi-Objective Evolutionary Algorithm: Performance Study and Comparisons. In Lee Spector, Erik D. Goodman, Annie Wu, W.B. Langdon, Hans-Michael Voigt, Mitsuo Gen, Sandip Sen, Marco Dorigo, Shahram Pezeshk, Max H. Garzon, and Edmund Burke, editors, Proceedings of the Genetic and Evolutionary Computation Conference (GECCO’2001), pages 647–654, San Francisco, California, 2001. Morgan Kaufmann Publishers.

[951] K.C. Tan, T.H. Lee, and E.F. Khor. Evolutionary Algorithms for Multi-Objective Optimization: Performance Assessments and Comparisons. In Proceedings of the Congress on Evolutionary Computation 2001 (CEC’2001), volume 2, pages 979–986, Piscataway, New Jersey, May 2001. IEEE Service Center.

[952] K.C. Tan, T.H. Lee, and E.F. Khor. Evolutionary Algorithms with Dynamic Population Size and Local Exploration for Multiobjective Optimization. IEEE Transactions on Evolutionary Computation, 5(6):565–588, December 2001.

[953] K.C. Tan, T.H. Lee, and E.F. Khor. Incrementing Multi-Objective Evolutionary Algorithms: Performance Studies and Comparisons. In Eckart Zitzler, Kalyanmoy Deb, Lothar Thiele, Carlos A. Coello Coello, and David Corne, editors, First International Conference on Evolutionary Multi-Criterion Optimization, pages 111–125. Springer-Verlag. Lecture Notes in Computer Science No. 1993, 2001.

[954] K.C. Tan, Tong H. Lee, D. Khoo, and E.F. Khor. A Multiobjective Evolutionary Algorithm Toolbox for Computer-Aided Multiobjective Optimization. IEEE Transactions on Systems, Man, and Cybernetics—Part B: Cybernetics, 31(4):537–556, August 2001.

[955] K.C. Tan, K. Sengupta, T.H. Lee, and R. Sthikannan. Autonomous Registration of Disparate Spatial Data via an Evolutionary Algorithm Toolbox. In Congress on Evolutionary Computation (CEC’2002), volume 1, pages 31–36, Piscataway, New Jersey, May 2002. IEEE Service Center.

[956] Masahiro Tanaka and Tetsuzo Tanino. Global optimization by the genetic algorithm in a multiobjective decision support system. In Proceedings of the 10th International Conference on Multiple Criteria Decision Making, volume 2, pages 261–270, 1992.

[957] Masahiro Tanaka, Hikaru Watanabe, Yasuyuki Furukawa, and Tetsuzo Tanino. GA-Based Decision Support System for Multicriteria Optimization. In Proceedings of the International Conference on Systems, Man, and Cybernetics, volume 2, pages 1556–1561, Piscataway, NJ, 1995. IEEE.

[958] K. S. Tang, K. F. Man, and D. W. Gu. Structured Genetic Algorithm for Robust h^{∞} Control Systems Design. IEEE Transactions on Industrial Electronics, 43(5):575–582, October 1996.

[959] K. S. Tang, K. F. Man, and K. T. Ko. Wireless LAN Design using Hierarchical Genetic Algorithm. In Thomas Bäck, editor, Proceedings of the Seventh International Conference on Genetic Algorithms, pages 629–635, San Mateo, California, July 1997. Michigan State University, Morgan Kaufmann Publishers.

[960] Kit-Sang Tang, King-Tim Ko, and Erick W.M. Wong. Optimal File Placement in VOD System Using Genetic Algorithm. IEEE Transactions on Industrial Electronics, 48(5):891–897, October 2001.

[961] Kit-Sang Tang, Kim-Fung Man, and S. Kwong. Wireless Communication Network Design in IC Factory. IEEE Transactions on Industrial Electronics, 48(2):452–459, April 2001.

[962] K.S. Tang, Kim-Fung Man, Guanrong Chen, and Sam Kwong. An Optimal Fuzzy PID Controller. IEEE Transactions on Industrial Electronics, 48(4):757–765, August 2001.

[963] Tetsuzo Tanino, Masahiro Tanaka, and Chihirp Hojo. An interactive multicriteria decision making method by using a genetic algorithm. In Proceedings of 2nd International Conference on Systems Science and Systens Engineering, pages 381–386, 1993.

[964] J. Teghem. Multiobjective Combinatorial Optimization. In PPSN/SAB Workshop on Multiobjective Problem Solving from Nature (MPSN), Paris, France, September 2000.

[965] Jürgen Teich. Pareto-Front Exploration with Uncertain Objectives. In Eckart Zitzler, Kalyanmoy Deb, Lothar Thiele, Carlos A. Coello Coello, and David Corne, editors, First International Conference on Evolutionary Multi-Criterion Optimization, pages 314–328. Springer-Verlag. Lecture Notes in Computer Science No. 1993, 2001.

[966] Jürgen Teich, Eckart Zitzler, and Shuvra S. Bhattacharyya. 3D Exploration of Software schedules for DSP Algorithms. In 7th International Workshop on Hardware/Software Codesign (CODES’99), pages 168–172, May 1999.

[967] Dirk Thierens and Peter A.N. Bosman. Multi-Objective Mixture-based Iterated Density Estimation Evolutionary Algorithms. In Lee Spector, Erik D. Goodman, Annie Wu, W.B. Langdon, Hans-Michael Voigt, Mitsuo Gen, Sandip Sen, Marco Dorigo, Shahram Pezeshk, Max H. Garzon, and Edmund Burke, editors, Proceedings of the Genetic and Evolutionary Computation Conference (GECCO’2001), pages 663–670, San Francisco, California, 2001. Morgan Kaufmann Publishers.

[968] Dirk Thierens and Peter A.N. Bosman. Multi-Objective Optimization with Iterated Density Estimation Evolutionary Algorithms using Mixture Models. In Proceedings of the Third International Symposium on Adaptive Systems— Evolutionary Computation and Probabilistic Graphical Models, pages 129– 136, Havana, Cuba, March 19–23 2001. Institute of Cybernetics, Mathematics and Physics.

[969] Mark W. Thomas. A Pareto Frontier for Full Stern Submarines via Genetic Algorithm. PhD thesis, Ocean Engineering Department, Massachusetts Institute of Technology, Cambridge, MA, june 1998.

[970] H. A. Thompson, P. J. Fleming, and A. J. Chipperfield. Multi-Objective Optimisation of Systems Architectures for Distributed Aero-Engine Control Systems. In The 43rd ASME Gas Turbine and Aeroengine Technical Congress, volume 98-GT-045, Stockolm, Sweden, June 1998. ASME Press.

[971] Robert Thomson and Tughrul Arslan. An Evolutionary Algorithm for the Multi-Objective Optimisation of VLSI Primitive Operator Filters. In Congress on Evolutionary Computation (CEC’2002), volume 1, pages 37–42, Piscataway, New Jersey, May 2002. IEEE Service Center.

[972] F. L. Tito, G. N. Taranto, and D. M. Falcão. Integrated Tuning of Generator Excitation Systems by a Multiobjective Genetic Algorithm. In Proceedings of the International Conference on Intelligent System Application to Power Systems (ISAP), Rio de Janeiro, Brazil, April 1999.

[973] Ahsutosh Tiwari, Rajkumar Roy, Graham Jared, and Olivier Munaux. Challenges in Real-Life Engineering Design Optimization: An Analysis. In 2001 Genetic and Evolutionary Computation Conference. Workshop Program, pages 289–294, San Francisco, California, July 2001.

[974] Ashutosh Tiwari and Rajkumar Roy. Application of Generalised Regression GA for Designing a Turbine Blade Cooling System. In Alwyn M. Barry, editor, GECCO 2002: Proceedings of the Bird of a Feather Workshops, Genetic and Evolutionary Computation Conference, pages 108–113, New York, July 2002. AAAI.

[975] Ashutosh Tiwari and Rajkumar Roy. Variable Response Interaction and Multiobjective Optimisation. In W.B. Langdon, E. Cantú-Paz, K. Mathias, R. Roy, D. Davis, R. Poli, K. Balakrishnan, V. Honavar, G. Rudolph, J. Wegener, L. Bull, M.A. Potter, A.C. Schultz, J.F. Miller, E. Burke, and N. Jonoska, editors, Proceedings of the Genetic and Evolutionary Computation Conference (GECCO’2002), pages 602–609, San Francisco, California, July 2002. Morgan Kaufmann Publishers.

[976] Ashutosh Tiwari, Rajkumar Roy, Graham Jared, and Olivier Munaux. Interaction and Multi-Objective Optimisation. In Lee Spector, Erik D. Goodman, Annie Wu, W.B. Langdon, Hans-Michael Voigt, Mitsuo Gen, Sandip Sen, Marco Dorigo, Shahram Pezeshk, Max H. Garzon, and Edmund Burke, editors, Proceedings of the Genetic and Evolutionary Computation Conference (GECCO’2001), pages 671–678, San Francisco, California, 2001. Morgan Kaufmann Publishers.

[977] David S. Todd. Multiple Criteria Genetic Algorithms in Engineering Design and Operation. PhD thesis, University of Newcastle, Newcastle-upon-Tyne, UK, October 1997.

[978] David S. Todd, J. Scott, and Pratyush Sen. A Genetic Algorithm Based Approach to Systems Scheduling. In IFAC Symposium on Large Scale Systems, Patras, Greece, July 1998.

[979] David S. Todd, J. Scott, and Pratyush Sen. Genetic Algorithms Applied to Ship Design and Manufacturing Processes. In PRADS 98, The Hague, September 1998.

[980] David S. Todd and Pratyush Sen. A Multiple Criteria Genetic Algorithm for Containership Loading. In Thomas Bäck, editor, Proceedings of the Seventh International Conference on Genetic Algorithms, pages 674–681, San Mateo, California, July 1997. Michigan State University, Morgan Kaufmann Publishers.

[981] David S. Todd and Pratyush Sen. Multiple Criteria Scheduling using Genetic Algorithms in a Shipyard Environment. In Proceedings of the 9th International Conference on Computer Applications in Shipbuilding, Yokohama, Japan, October 1997.

[982] David S. Todd and Pratyush Sen. Distributed Task Scheduling and Allocation Using Genetic Algorithms. In 24th International Conference on Computers and Industrial Engineering, Brunel University, UK, September 1998.

[983] David S. Todd and Pratyush Sen. Tackling Complex Job Shop Problems Using Operation Based Scheduling. In Ian Parmee, editor, The Integration of Evolutionary and Adaptive Computing Technologies with Product/System Design and Realisation, pages 45–58, Plymouth, United Kingdom, April 1998. Plymouth Engineering Design Centre, Springer-Verlag.

[984] David S. Todd and Pratyush Sen. Directed Multiple Objective Search of Design Spaces Using Genetic Algorithms and Neural Networks. In W. Banzhaf, J. Daida, A. E. Eiben, M. H. Garzon, V. Honavar, M. Jakiela, and R. E. Smith, editors, Proceedings of the Genetic and Evolutionary Computation Conference (GECCO’99), volume 2, pages 1738–1743, San Francisco, California, July 1999. Morgan Kaufmann.

[985] Andrea Toffolo and Ernesto Benini. A New Pareto-like Evaluation Method for Finding Multiple Global Optima in Evolutionary Algorithms. In Late Breaking Papers at the 2000 Genetic and Evolutionary Computation Conference, pages 405–410, Las Vegas, Nevada, July 2000.

[986] V. G. Toshinsky, H. Sekimoto, and G. I. Toshinsky. A method to improve multiobjective genetic algorithm optimization of a self-fuel-providing LMFBR by niche induction among nondominated solutions. Annals of Nuclear Energy, 27(5):397–410, March 2000.

[987] A. Trebi-Ollennu and B. A. White. Multiobjective Fuzzy Genetic Algorithm Optimisation Approach to Nonlinear Control System Design. In UKACC International Conference on Control, volume 1, pages 479–484, San Francisco, California, September 1996.

[988] A. Trebi-Ollennu and B. A. White. Multiobjective Fuzzy Genetic Algorithm Optimization Approach to Nonlinear Control System Design. IEE Proceedings, Control Theory and Applications, 144(2):137–142, March 1997.

[989] Effie Tsoi, Kit Po Wong, and Chun Che Fung. Hybrid GA/SA Algorithms for Evaluating Trade-off Between Economic Cost and Environmental Impact in Generation Dispatch. In David B. Fogel, editor, Proceedings of teh Second IEEE Conference on Evolutionary Computation (ICEC’95), pages 132–137, Piscataway, New Jersey, 1995. IEEE Press.

[990] D. Tuyttens, J. Teghem, Ph. Fortemps, and K. Van Nieuwenhuyze. Performance of the MOSA Method for the Bicriteria Assignment Problem. Journal of Heuristics, 6(3):295–310, August 2000.

[991] Gwo.Hshiung Tzeng and Jen-Swei Kuo. Fuzzy Multiobjective Double Samplig Plans with Genetic Algorithms Based on Bayesian Model. In Weiling Chiang and Jonathan Lee, editors, Proceedings of the International Joint Conference of CFSA/IFIS/SOFT95 on Fuzzy Theory and Applications, pages 59–64, Singapore, 1995. World Scientific.

[992] E.L. Ulungu, J. Teghem, and Ph. Fortemps. Heuristics for multi-objective combinatorial optimization by simulated annealing. In J. Gu, G. Chen, Q. Wei, and S. Wang, editors, Multiple Criteria Decision Making: Theory and Applications. Proceedings of the 6th National Conference on Multiple Criteria Decision Making, pages 228–238, Windsor, UK, 1995. Sci-Tech.

[993] E.L. Ulungu, J. Teghem, Ph. Fortemps, and D. Tuyttens. MOSA Method: A Tool for Solving Multiobjective Combinatorial Optimization Problems. Journal of Multi-Criteria Decision Analysis, 8(4):221–236, 1999.

[994] E.L. Ulungu, J. Teghem, and Ch. Ost. Efficiency of interactive multi-objective simulated annealing through a case study. Journal of the Operational Research Society, 49:1044–1050, 1998.

[995] Jean-Philippe Vacher, Alain Cardon, and Thierry Galinho. Genetic Algorithms on Multi-Agents Systems. In Proceedings of EUROGEN’97, Trieste, Italie, December 1997.

[996] Jean-Philippe Vacher, Alain Cardon, and Thierry Galinho. Algorithmes Génétiques Multi-Objectif en Ordonnancement de Production de type Job-Shop. In FRANCORO II, page 58, Sousse, Tunisie, Avril 1998. (In French).

[997] Jean-Philippe Vacher, Alain Cardon, and Thierry Galinho. Genetic Algorithm using Multi-Objective in a Multi-Agent System. In The Second International Symposium On Intelligent Manaufacturing Systems IMS’98, pages 273–282, Sakarya, Turquie, August 1998.

[998] Jean-Philippe Vacher, Alain Cardon, and Thierry Galinho. Heuristics Granularity in a Multi-Agents Systems (MAS): Application to the Production Management. In Proceedings of ????, page 8, Guildford, UK, March 1998.

[999] Jean-Philippe Vacher, Alain Cardon, and Thierry Galinho. Information System for Management : Random generators for Job-Shop Scheduling Problems. In The Seventh International Conference Information Systems Development ISD’98, Bled, Slovenia, September 1998.

[1000] Jean-Philippe Vacher, Alain Cardon, and Thierry Galinho. Information Systems for Management in Job-Shop Scheduling Problems using a Multi-Objective Genetic Algorithm. In NIMES’98, Complex Systems, Intelligent Systems & Interfaces, pages 97–100, May 1998.

[1001] Jean-Philippe Vacher, Alain Cardon, and Thierry Galinho. Macrophagic Agents in Multi-agent Systems to Resolve Job-Shop Scheduling Problem. In The 32^{th} Hawaii International Conference on System Sciences HICSS-32, Maui, Hawaii, January 1999.

[1002] Jean-Philippe Vacher, Alain Cardon, Franck Lesage, and Thierry Galinho. Genetic Algorithms in a Multi-Agent System. In Proceedings IEEE International Joint Symposia on Intelligence and Systems, pages 17–26, Rockville, MD, USA, May 1998.

[1003] Jean-Philippe Vacher and Thierry Galinho. A Multi-Agent System Using Multi-Objective Genetic Algorithm: A Solution with Macrophagic Agents. In Proceedings of the 4^{th} International Conference on Information Systems, Analysis and Synthesis, ISAS’98, World Multiconference on Systems, Cybernetics and Informatics (SCI’98), pages 326–332, Orlando, Florida, July 1998.

[1004] Jean-Philippe Vacher, Thierry Galinho, and Zoubir Mammeri. Une application des algorithmes génétiques à l’ordonnancement d’atelier. In Ed. Hermes, editor, Proceeding de la conférence MOSIM’97, pages 43–50, Mont-Saint-Aignan, Juin 1997. (in French).

[1005] Christine L. Valenzuela. A Simple Evolutionary Algorithm for Multi-Objective Optimization (SEAMO). In Congress on Evolutionary Computation (CEC’2002), volume 1, pages 717–722, Piscataway, New Jersey, May 2002. IEEE Service Center.

[1006] Manuel Valenzuela-Rendón and Eduardo Uresti-Charre. A Non-Generational Genetic Algorithm for Multiobjective Optimization. In Thomas Bäck, editor, Proceedings of the Seventh International Conference on Genetic Algorithms, pages 658–665, San Mateo, California, July 1997. Michigan State University, Morgan Kaufmann Publishers.

[1007] David A. Van Veldhuizen, Jesse B. Zydallis, and Gary B. Lamont. Issues in Parallelizing Multiobjective Evolutionary Algorithms for Real World Applications. In Proceedings of the 17th ACM Symposium on Applied Computing, pages 595–602, Madrid, Spain, 2002. ACM Press.

[1008] Ganesh Vedarajan, Louis Chi Chan, and David E. Goldberg. Investment Portfolio Optimization using Genetic Algorithms. In John R. Koza, editor, Late Breaking Papers at the Genetic Programming 1997 Conference, pages 255– 263, Stanford University, California, July 1997. Stanford Bookstore.

[1009] David A. Van Veldhuizen. Multiobjective Evolutionary Algorithms: Classifications, Analyses, and New Innovations. PhD thesis, Department of Electrical and Computer Engineering. Graduate School of Engineering. Air Force Institute of Technology, Wright-Patterson AFB, Ohio, May 1999.

[1010] David A. Van Veldhuizen and Gary B. Lamont. Evolutionary Computation and Convergence to a Pareto Front. In John R. Koza, editor, Late Breaking Papers at the Genetic Programming 1998 Conference, pages 221–228, Stanford University, California, July 1998. Stanford University Bookstore.

[1011] David A. Van Veldhuizen and Gary B. Lamont. Multiobjective Evolutionary Algorithm Research: A History and Analysis. Technical Report TR-98-03, Department of Electrical and Computer Engineering, Graduate School of Engineering, Air Force Institute of Technology, Wright-Patterson AFB, Ohio, 1998.

[1012] David A. Van Veldhuizen and Gary B. Lamont. Genetic Algorithms, Building Blocks, and Multiobjective Optimization. In Annie S. Wu, editor, Proceedings of the 1999 Genetic and Evolutionary Computation Conference. Workshop Program, pages 125–126, Orlando, Florida, July 1999.

[1013] David A. Van Veldhuizen and Gary B. Lamont. MOEA Test Suite Generation, Design & Use. In Annie S. Wu, editor, Proceedings of the 1999 Genetic and Evolutionary Computation Conference. Workshop Program, pages 113–114, Orlando, Florida, July 1999.

[1014] David A. Van Veldhuizen and Gary B. Lamont. Multiobjective Evolutionary Algorithm Test Suites. In Janice Carroll, Hisham Haddad, Dave Oppenheim, Barrett Bryant, and Gary B. Lamont, editors, Proceedings of the 1999 ACM Symposium on Applied Computing, pages 351–357, San Antonio, Texas, 1999. ACM.

[1015] David A. Van Veldhuizen and Gary B. Lamont. Multiobjective Evolutionary Algorithms: Analyzing the State-of-the-Art. Evolutionary Computation, 8(2):125–147, 2000.

[1016] David A. Van Veldhuizen and Gary B. Lamont. Multiobjective Optimization with Messy Genetic Algorithms. In Proceedings of the 2000 ACM Symposium on Applied Computing, pages 470–476, Villa Olmo, Como, Italy, 2000. ACM.

[1017] David A. Van Veldhuizen and Gary B. Lamont. On Measuring Multiobjective Evolutionary Algorithm Performance. In 2000 Congress on Evolutionary Computation, volume 1, pages 204–211, Piscataway, New Jersey, July 2000. IEEE Service Center.

[1018] David A. Van Veldhuizen, Brian S. Sandlin, , Robert M. Marmelstein, and Gary B. Lamont. Finding Improved Wire-Antenna Geometries with Genetic Algorithms. In David B. Fogel, editor, Proceedings of the 1998 International Conference on Evolutionary Computation, pages 102–107, Piscataway, New Jersey, 1998. IEEE.

[1019] V. Rao Vemuri and Walter Cedeño. A New Genetic Algorithm for Multi Objective Optimization in Water Resource Management. In Proceedings of the Second IEEE International Conference on Evolutionary Computation, pages 495–500, Piscataway, New Jersey, 1995. IEEE Press.

[1020] V. Rao Vemuri and Walter Cedeño. A New Genetic Algorithm for Multi-Objective Optimization in Water Resource Management. In 1996 Knowledgebased Computer Systems, Bombay, India, December 1996. KBIS Proceedings.

[1021] V. Venugopal and T. T. Narendran. A Genetic Algorithm Approach to the Machine-Component Grouping Problem with Multiple Objectives. Computers and Industrial Engineering, 22(4):469–480, 1992.

[1022] Alessandro Vicini and Domenico Quagliarella. Inverse and Direct Airfoil Design Using a Multiobjective Genetic Algorithm. AIAA Journal, 35(9):1499– 1505, September 1997.

[1023] Alessandro Vicini and Domenico Quagliarella. Multipoint transonic airfoil design by means of a multiobjective genetic algorithm. In 35th AIAA Aerospace Sciences Meeting and Exhibit, Reno, Nevada, January 1997. American Institute of Aeronautics and Astronautics (AIAA). AIAA Paper 97-0082.

[1024] Alessandro Vicini and Domenico Quagliarella. Airfoil and Wing Design Through Hybrid Optimization Strategies. In 16th Applied Aerodynamics Conference, Albuquerque, New Mexico, June 1998. American Institute of Aeronautics and Astronautics (AIAA). AIAA Paper 98-2729.

[1025] Rémy Viennet, Christian Fontiex, and Ivan Marc. New Multicriteria Optimization Method Based on the Use of a Diploid Genetic Algorithm: Example of an Industrial Problem. In J. M. Alliot, E. Lutton, E. Ronald, M. Schoenauer, and D. Snyers, editors, Proceedings of Artificial Evolution (European Conference, selected papers), pages 120–127, Brest, France, September 1995. Springer-Verlag.

[1026] Rémy Viennet, Christian Fontiex, and Ivan Marc. Multicriteria Optimization Using a Genetic Algorithm for Determining a Pareto Set. Journal of Systems Science, 27(2):255–260, 1996.

[1027] S. Voget. Multiobjective optimization with genetic algorithm and fuzzy control. In Proceedings of the Fourth European Conference on Intelligent Techniques and Soft Computing (EUFIT’96), pages 391–394, Aachen, Germany, 1996.

[1028] Stefan Voget and Michael Kolonko. Multidimensional Optimization with a Fuzzy Genetic Algorithm. Journal of Heuristics, 4(3):221–244, September 1998.

[1029] David R. Wallace, Mark J. Jakiela, and W.C. Flowers. Design Search Under Probabilistic Specifications Using Genetic Algorithms. Computer-Aided Design, 28(5):405–421, 1996.

[1030] Jiang Feng Wang and Jacques Periaux. Multi-Point Optimization using GAs and Nash/Stackelberg Games for High Lift Multi-airfoil Design in Aerodynamics. In Proceedings of the Congress on Evolutionary Computation 2001 (CEC’2001), volume 1, pages 552–559, Piscataway, New Jersey, May 2001. IEEE Service Center.

[1031] S. Watanabe, T. Hiroyasu, and M. Miki. Parallel Evolutionary Multi-Criterion Optimization for Mobile Telecommunication Networks Optimization. In K.C. Giannakoglou, D.T. Tsahalis, J. Periaux, K.D. Papailiou, and T. Fogarty, editors, Evolutionary Methods for Design, Optimization and Control with Applications to Industrial Problems. Proceedings of the EUROGEN’2001. Athens. Greece, September 19-21, pages 167–172, Baracelona, Spain, 2001. International Center for Numerical Methods in Engineering(CIMNE).

[1032] Shinya Watanabe, Tomoyuki Hiroyasu, and Mitsunori Miki. LCGA: Local Cultivation Genetic Algorithm for Multi-Objective Optimization Problems. In W.B. Langdon, E. Cantú-Paz, K. Mathias, R. Roy, D. Davis, R. Poli, K. Balakrishnan, V. Honavar, G. Rudolph, J. Wegener, L. Bull, M.A. Potter, A.C. Schultz, J.F. Miller, E. Burke, and N. Jonoska, editors, Proceedings of the Genetic and Evolutionary Computation Conference (GECCO’2002), page 702, San Francisco, California, July 2002. Morgan Kaufmann Publishers.

[1033] Shinya Watanabe, Tomoyuki Hiroyasu, and Mitsunori Miki. NCGA: Neighborhood Cultivation Genetic Algorithm for Multi-Objective Optimization Problems. In Erick Cantú-Paz, editor, 2002 Genetic and Evolutionary Computation Conference. Late-Breaking Papers, pages 458–465, New York, July 2002.

[1034] Richard A. Watson. Problem Decomposition and Multi-Objective Optimization. In PPSN/SAB Workshop on Multiobjective Problem Solving from Nature (MPSN), Paris, France, September 2000.

[1035] Andrew Webb. Evolutionary Techniques Applied to Wide Area Network Planning. PhD thesis, School of Engineering, Circuits And Systems Research Group, University of Wales Cardiff, UK, July 2000.

[1036] D. S. Weile and E. Michielssen. Integer coded Pareto genetic algorithm design of constrained antenna arrays. Electronics Letters, 32(19):1744–1745, September 1996.

[1037] D. S. Weile and E. Michielssen. Integer coded Pareto genetic algorithm design of constrained antenna arrays. Technical Report CCEM-13-96, Electrical and Computer Engineering Department, Center for Computational Electromagnetics, University of Illinois at Urbana-Champaign, November 1996.

[1038] D. S. Weile and E. Michielssen. Multiobjective optimization of electromagnetic devices using Pareto genetic algorithms. In Proceedings of the 1996 Antenna Applications Symposium, pages 1–18, Amherst, Massachusetts, 1996.

[1039] D. S. Weile, E. Michielssen, and D. E. Goldberg. Genetic algorithm design of pareto optimal broad-band microwave absorbers. Technical Report CCEM-4-96, Electrical and Computer Engineering Department, Center for Computational Electromagnetics, University of Illinois at Urbana-Champaign, May 1996.

[1040] D. S. Weile, E. Michielssen, and D. E. Goldberg. Genetic algorithm design of Pareto optimal broadband microwave absorbers. IEEE Transactions on Electromagnetic Compatibility, 38(3):518–525, August 1996.

[1041] D. S. Weile, E. Michielssen, and D. E. Goldberg. Multiobjective synthesis of electromagnetic devices using nondominated sorting genetic algorithms. In 1996 IEEE Antennas and Propagation Society International Symposium Digest, volume 1, pages 592–595, Baltimore, Maryland, July 1996.

[1042] J. F. Whidborne, D.-W. Gu, and I. Postlethwaite. Algorithms for the Method of Inequalities — A Comparative Study. In Procedings of the 1995 American Control Conference, pages 3393–3397, Seattle, Washington, 1995.

[1043] P. B. Wienke, C. Lucasius, and G. Kateman. Multicriteria target optimization of analytical procedures using a genetic algorithm. Analytica Chimica Acta, 265(2):211–225, 1992.

[1044] P. B. Wilson and M. D. Macleod. Low implementation cost IIR digital filter design using genetic algorithms. In IEE/IEEE Workshop on Natural Algorithms in Signal Processing, pages 4/1–4/8, Chelmsford, U.K., 1993.

[1045] Jonathan Wright and Heather Loosemore. An Infeasibility Objective for Use in Constrained Pareto Optimization. In Eckart Zitzler, Kalyanmoy Deb, Lothar Thiele, Carlos A. Coello Coello, and David Corne, editors, First International Conference on Evolutionary Multi-Criterion Optimization, pages 256– 268. Springer-Verlag. Lecture Notes in Computer Science No. 1993, 2001.

[1046] Jin Wu. Quality Assisted Multiobjective and Multidisciplinary Genetic Algorithms. PhD thesis, Department of Mechanical Engineering, University of Maryland at College Park, College Park, Maryland, 2001.

[1047] Jin Wu and Shapour Azarm. Metrics for Quality Assessment of a Multiobjective Design Optimization Solution Set. Transactions of the ASME, Journal of Mechanical Design, 123:18–25, 2001.

[1048] Jin Wu and Shapour Azarm. On a New Constraint Handling Technique for Multi-Objective Genetic Algorithms. In Lee Spector, Erik D. Goodman, Annie Wu, W.B. Langdon, Hans-Michael Voigt, Mitsuo Gen, Sandip Sen, Marco Dorigo, Shahram Pezeshk, Max H. Garzon, and Edmund Burke, editors, Proceedings of the Genetic and Evolutionary Computation Conference (GECCO’2001), pages 741–748, San Francisco, California, 2001. Morgan Kaufmann Publishers.

[1049] Ningchuan Xiao, David A. Bennet, and Marc P. Armstrong. Using evolutionary algorithms to generate alternatives for multiobjective site-search problems. Environment and Planning A, 34(4):639–656, April 2002.

[1050] Jian-Xin Xu, Sanjib Kumar Panda, and Qing Zheng. Multiobjective Optimization of Current Waveforms for Switched Reluctance Motors by Genetic Algorithm. In Congress on Evolutionary Computation (CEC’2002), volume 2, pages 1860–1865, Piscataway, New Jersey, May 2002. IEEE Service Center.

[1051] Kazuo Yamasaki. Dynamic Pareto Optimum GA against the changing environments. In 2001 Genetic and Evolutionary Computation Conference. Workshop Program, pages 47–50, San Francisco, California, July 2001.

[1052] Xiaofeng Yang and Mitsuo Gen. Evolution program for bicriteria transportation problem. In M. Gen and T. Kobayashi, editors, Proceedings of the 16th International Conference on Computers and Industrial Engineering, pages 451– 454, Ashikaga, Japan, 1994. Pergamon Press.

[1053] Zichen Yang and Bo Meng. A multi-objective genetic algorithms method for generating pareto solutions in bilateral negotiations. In Proceedings of the 4th World Congress on Intelligent Control and Automation, volume 2, pages 1985– 1989, 2002.

[1054] Leehter Yao, William A. Sethares, and Daniel C. Kammer. Sensor Placement for On-Orbit Modal Identification via a Genetic Algorithm. AIAA Journal, 31(10):1922–1928, October 1993.

[1055] P. O. Yapo, H. V. Gupta, and S. Sorooshian. Multi-Objective Global Optimization for Hydrologic Models. Journal of Hydrology, 204:83–97, 1998.

[1056] Patrice Ogou Yapo. A multiobjective global optimization algorithm with application to the calibration of hydrologic models. PhD thesis, Department of Systems and Industrial Engineering, The University of Arizona, Tucson, Arizona, 1996.

[1057] Chao-Hsien Yeh and John W. Labadie. Multiobjective Watershed-Level Planning of Storm-Water Detention Systems. Journal of Water Resources Planning and Management, 123(6):336–343, November/December 1997.

[1058] Gary G. Yen and Haiming Lu. Hierarchical Rank Density Genetic Algorithm for Radial-Basis Function Neural Network Design. In Congress on Evolutionary Computation (CEC’2002), volume 1, pages 25–30, Piscataway, New Jersey, May 2002. IEEE Service Center.

[1059] Kai Yen and Lajos Hanzo. Genetic Algorithm Based Antenna Diversity Assisted Multiuser Detection for Synchronous CDMA Systems. In IEEE Vehicular Technology Conference, volume 3, pages 1794–1798. IEEE, 2001.

[1060] L. Yinzhen, K. Ida, and M. Gen. Improved Genetic Algorithm for Solving Multiobjective Solid Transportation Problem with Fuzzy Numbers. Computers and Industrial Engineering, 33(3):589–592, December 1997.

[1061] J. Yoo and P. Hajela. Immune network simulations in multicriterion design. Structural Optimization, 18:85–94, 1999.

[1062] Seung-Ryul Yoo. Determination of Operational Frequencies on Express Bus Service using Dynamic Niche Sharing Pareto GA. PhD thesis, Graduate School of Korea University, Department of Industrial Engineering, Korea, 1997. (in Korean).

[1063] Koji Yoshida, Masayuki Yamamura, and Shigenobu Kobayashi. Generating Pareto Optimal Decision Trees by GAs. In Proceedings of the 4th International Conference on Soft Computing (IIZUKA’96), pages 854–859, 1996.

[1064] Kazuyuki Yoshimura and Ryohei Nakano. Genetic Algorithms for Information Operator Scheduling. In David B. Fogel, editor, Proceedings of the 1998 International Conference on Evolutionary Computation, pages 277–282, Piscataway, New Jersey, 1998. IEEE.

[1065] Susumu Yoshizawa, Tohru Kawabe, and Sadaaki Miyamoto. Robust Control Configured Design Method for Systems with Multi-objective Specifications. In Joint 9th IFSA World Congress and 20th NAFIPS International Conference, volume 3, pages 1746–1751. IEEE, 2001.

[1066] Abdulnasser Younes, Hamada Ghenniwa, and Shawki Areibi. An adaptive genetic algorithm for multi objective flexible manufacturing systems. In W.B. Langdon, E. Cantú-Paz, K. Mathias, R. Roy, D. Davis, R. Poli, K. Balakrishnan, V. Honavar, G. Rudolph, J. Wegener, L. Bull, M.A. Potter, A.C. Schultz, J.F. Miller, E. Burke, and N. Jonoska, editors, Proceedings of the Genetic and Evolutionary Computation Conference (GECCO’2002), pages 1241– 1248, San Francisco, California, July 2002. Morgan Kaufmann Publishers.

[1067] Habib Youssef, Sadiq M. Sait, and Salman A. Khan. Fuzzy Simulated Evolution Algorithm for Topology Design on Campus Networks. In 2000 Congress on Evolutionary Computation, volume 1, pages 180–187, Piscataway, New Jersey, July 2000. IEEE Service Center.

[1068] Habib Youssef, Sadiq M. Sait, and Salman A. Khan. Fuzzy Evolutionary Hybrid Metaheuristic for Network Topology Design. In Eckart Zitzler, Kalyanmoy Deb, Lothar Thiele, Carlos A. Coello Coello, and David Corne, editors, First International Conference on Evolutionary Multi-Criterion Optimization, pages 400–415. Springer-Verlag. Lecture Notes in Computer Science No. 1993, 2001.

[1069] Y. Yu, JB. Y. Zhang, and M. C. Schell. Multi-objective Stochastic Reasoning and Genetic Algorithm Optimization: Applications in Radiotherapy Treatment Planning. In D. W. Pearson, editor, Proceedings of the Second International ICSC Symposium on Soft Computing, pages 231–237, Nîmes, France, September 1997. ICSC Academic Press.

[1070] Yan Yu. Multi-objective decision theory for computational optimization in radiation therapy. Medical Physics, 24:1445–1454, 1997.

[1071] R. S. Zebulum, M. A. Pacheco, and M. Vellasco. A multi-objective optimisation methodology applied to the synthesis of low-power operational amplifiers. In Ivan Jorge Cheuri and Carlos Alberto dos Reis Filho, editors, Proceedings of the XIII International Conference in Microelectronics and Packaging, volume 1, pages 264–271, Curitiba, Brazil, August 1998.

[1072] R. S. Zebulum, M. A. Pacheco, and M. Vellasco. Synthesis of CMOS operational amplifiers through Genetic Algorithms. In Proceedings of the Brazilian Symposium on Integrated Circuits, SBCCI’98, pages 125–128, Rio de Janeiro, Brazil, September 1998. IEEE.

[1073] Ricardo Salem Zebulum, Marco Aurélio Pacheco, and Marley Vellasco. Artificial Evolution of Active Filters: A Case Study. In Proceedings of the First NASA/DoD Workshop on Evolvable Hardware, pages 66–75, Los Alamitos, California, July 1999. IEEE Computer Society.

[1074] Byoung-Tak Zhang and Heinz Mühlenbein. Adaptive Fitness Functions for Dynamic Growing/Pruning of Program Trees. In Peter J. Angeline and Jr. Kenneth E. Kinnear, editors, Advances in Genetic Programming 2, pages 241– 256. MIT Press, 1996.

[1075] Yong Zhao, Carlos A. Brizuela, and Nobuo Sannomiya. Application of the Partial Enumeration Selection Method in Genetic Algorithms to Solving a Multi-Objective Flowshop Problem. In 2001 IEEE International Conference on Systems, Man, and Cybernetics, volume 4, pages 2365–2370, 2001.

[1076] Gengui Zhou and Mitsuo Gen. Evolutionary Computation on Multicriteria Production Process Planning Problem. In Proceedings of the 1997 IEEE International Conference on Evolutionary Computation, pages 419–424, Piscataway, New Jersey, April 1997. IEEE Press.

[1077] Gengui Zhou and Mitsuo Gen. Genetic Algorithm Approach on Multi-Criteria Minimum Spanning Tree Problem. European Journal of Operational Research, 114(1), April 1999.

[1078] Bo Zhu, Ho Seong Lee, Lin Guo, and Masayoshi Tomizuka. Robust Tuning of Fixed-Structure Controller for Disk Drives using Statistical Model and Multi-Objective Genetic Algorithms. In Proceedings of the 2001 American Control Conference, volume 4, pages 2773–2778. IEEE, 2001.

[1079] Zhong-Yao Zhu and Kwong-Sak Leung. Asynchronous Self-Adjustable Island Genetic Algorithm for Multi-Objective Optimization Problems. In Congress on Evolutionary Computation (CEC’2002), volume 1, pages 837–842, Piscataway, New Jersey, May 2002. IEEE Service Center.

[1080] Zhong-Yao Zhu and Kwong-Sak Leung. An Enhanced Annealing Genetic Algorithm for Multi-Objective Optimization Problems. In W.B. Langdon, E. Cantú-Paz, K. Mathias, R. Roy, D. Davis, R. Poli, K. Balakrishnan, V. Honavar, G. Rudolph, J. Wegener, L. Bull, M.A. Potter, A.C. Schultz, J.F. Miller, E. Burke, and N. Jonoska, editors, Proceedings of the Genetic and Evolutionary Computation Conference (GECCO’2002), pages 658–665, San Francisco, California, July 2002. Morgan Kaufmann Publishers.

[1081] Eckart Zitzler. Evolutionary Algorithms for Multiobjective Optimization: Methods and Applications. PhD thesis, Swiss Federal Institute of Technology (ETH), Zurich, Switzerland, November 1999.

[1082] Eckart Zitzler, Kalyanmoy Deb, and Lothar Thiele. Comparison of Multiobjective Evolutionary Algorithms: Empirical Results. Technical Report 70, Computer Engineering and Networks Laboratory (TIK), Swiss Federal Institute of Technology (ETH) Zurich, Gloriastrasse 35, CH-8092 Zurich, Switzerland, December 1999.

[1083] Eckart Zitzler, Kalyanmoy Deb, and Lothar Thiele. Comparison of Multiobjective Evolutionary Algorithms on Test Functions of Different Difficulty. In Annie S. Wu, editor, Proceedings of the 1999 Genetic and Evolutionary Computation Conference. Workshop Program, pages 121–122, Orlando, Florida, July 1999.

[1084] Eckart Zitzler, Kalyanmoy Deb, and Lothar Thiele. Comparison of Multiobjective Evolutionary Algorithms: Empirical Results. Evolutionary Computation, 8(2):173–195, Summer 2000.

[1085] Eckart Zitzler, Marco Laumanns, and Lothar Thiele. SPEA2: Improving the Strength Pareto Evolutionary Algorithm. Technical Report 103, Computer Engineering and Networks Laboratory (TIK), Swiss Federal Institute of Technology (ETH) Zurich, Gloriastrasse 35, CH-8092 Zurich, Switzerland, May 2001.

[1086] Eckart Zitzler, Marco Laumanns, and Lothar Thiele. SPEA2: Improving the Strength Pareto Evolutionary Algorithm. In K. Giannakoglou, D. Tsahalis, J. Periaux, P. Papailou, and T. Fogarty, editors, EUROGEN 2001. Evolutionary Methods for Design, Optimization and Control with Applications to Industrial Problems, Athens, Greece, September 2001.

[1087] Eckart Zitzler, Marco Laumanns, Lothar Thiele, Carlos M. Fonseca, and Viviane Grunert da Fonseca. Why Quality Assessment of Multiobjective Optimizers Is Difficult. In W.B. Langdon, E. Cantú-Paz, K. Mathias, R. Roy, D. Davis, R. Poli, K. Balakrishnan, V. Honavar, G. Rudolph, J. Wegener, L. Bull, M.A. Potter, A.C. Schultz, J.F. Miller, E. Burke, and N. Jonoska, editors, Proceedings of the Genetic and Evolutionary Computation Conference (GECCO’2002), pages 666–673, San Francisco, California, July 2002. Morgan Kaufmann Publishers.

[1088] Eckart Zitzler, Jürgen Teich, and Shuvra S. Bhattacharyya. Evolutionary Algorithm Based Exploration of Software Schedules for Digital Signal Processors. In W. Banzhaf, J. Daida, A. E. Eiben, M. H. Garzon, V. Honavar, M. Jakiela, and R. E. Smith, editors, Proceedings of the Genetic and Evolutionary Computation Conference (GECCO’99), volume 2, pages 1762–1769, San Francisco, California, July 1999. Morgan Kaufmann.

[1089] Eckart Zitzler, Jürgen Teich, and Shuvra S. Bhattacharyya. Multidimensional Exploration of Software Implementations for DSP Algorithms. Journal of VLSI Signal Processing, 24(1):83–98, February 2000.

[1090] Eckart Zitzler and Lothar Thiele. An Evolutionary Algorithm for Multiobjective Optimization: The Strength Pareto Approach. Technical Report 43, Computer Engineering and Communication Networks Lab (TIK), Swiss Federal Institute of Technology (ETH), Zurich, Switzerland, May 1998.

[1091] Eckart Zitzler and Lothar Thiele. Multiobjective Optimization Using Evolutionary Algorithms—A Comparative Study. In A. E. Eiben, editor, Parallel Problem Solving from Nature V, pages 292–301, Amsterdam, September 1998. Springer-Verlag.

[1092] Eckart Zitzler and Lothar Thiele. Multiobjective Evolutionary Algorithms: A Comparative Case Study and the Strength Pareto Approach. IEEE Transactions on Evolutionary Computation, 3(4):257–271, November 1999.

[1093] Igor S. Zwir and Enrique H. Ruspini. Qualitative Object Description: Initial Reports of the Exploration of the Frontier. In Procedings of the Joint EUROFUSE—SIC’99 International Conference, Budapest, Hungary, 1999.

[1094] Jesse B. Zydallis and Gary B. Lamont. Solving of Discrete Multiobjective Problems Using an Evolutionary Algorithm with a Repair Mechanism. In Proceedings of the IEEE 2001 Midwest Symposium on Circuits and Systems, volume 1, pages 470–473. IEEE, 2001.

[1095] Jesse B. Zydallis, Gary B. Lamont, and David A. Van Veldhuizen. Messy Genetic Algorithm Based Multi-Objective Optimization: A Comparative Statistical Analysis. In PPSN/SAB Workshop on Multiobjective Problem Solving from Nature (MPSN), Paris, France, September 2000.

[1096] Jesse B. Zydallis, Todd A. Sriver, and Gary B. Lamont. Multiobjective Evolutionary Algorithm Approach for Solving Integer Based Optimization Problems. In W.B. Langdon, E. Cantú-Paz, K. Mathias, R. Roy, D. Davis, R. Poli, K. Balakrishnan, V. Honavar, G. Rudolph, J. Wegener, L. Bull, M.A. Potter, A.C. Schultz, J.F. Miller, E. Burke, and N. Jonoska, editors, Proceedings of the Genetic and Evolutionary Computation Conference (GECCO’2002), page 1276, San Francisco, California, July 2002. Morgan Kaufmann Publishers.

[1097] Jesse B. Zydallis, David A. Van Veldhuizen, and Gary B. Lamont. A Statistical Comparison of Multiobjective Evolutionary Algorithms Including the MOMGA–II. In Eckart Zitzler, Kalyanmoy Deb, Lothar Thiele, Carlos A. Coello Coello, and David Corne, editors, First International Conference on Evolutionary Multi-Criterion Optimization, pages 226–240. Springer-Verlag. Lecture Notes in Computer Science No. 1993, 2001.